### v2

# A Context for Manifold Calculus

Abstract.
We develop Weiss’s manifold calculus in the setting of $\infty$-categories, where we allow the target $\infty$-category to be any $\infty$-category with small limits. We will establish the connection between polynomial functors, Kan extensions, and Weiss sheaves, and will classify homogeneous functors. We will also generalize Weiss and Boavida de Brito’s theorem to functors taking values in arbitrary $\infty$-categories with small limits.
2020 Mathematics Subject Classification:
55U35, 57R40, 18F50
1. Introduction
Let $M$ be a smooth $n$-manifold and let $\operatorname{Open}\left(M\right)$ be the poset of open sets of $M$. The idea of Weiss’s manifold calculus is to study space-valued presheaves $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ on $M$ by constructing a tower of presheaves
$F\to\cdots\to T_{k}F\to\cdots\to T_{1}F\to T_{0}F$
such that each $T_{k}F$ is the best “polynomial functor” of degree $\leq k$ which approximates $F$. Borrowing an analogy from calculus, the tower is called the Taylor tower of $F$ and the functors $T_{k}F$ the polynomial approximations of $F$.
The Taylor tower is determined by how we define what it means for a functor to be polynomial, so we must make this explicit. There is a heuristic in homotopy theory that homology behaves like linear (or affine) functions. A homology theory is more or less characterized by the excision axiom, so we make the following definition:
Definition 1.1.
[Wei99, Definition 2.2] We say that $F$ is polynomial of degree $\leq k$, or $k$-excisive, if for each open set $U\subset M$ and pairwise disjoint closed subsets $A_{0},\dots,A_{k}$ of $U$, the map
$FU\to\underset{\emptyset\subsetneq S\subset\{0,\dots,k\}}{\operatorname{holim}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
is a homotopy equivalence.
For example, $1$-excisivity is equivalent to the condition that, for every open set $U\subset M$ and every pair of disjoint closed sets $A_{0},A_{1}\subset U$, the square
${{F(U)}}$${{F(U\setminus A_{1})}}$${{F(U\setminus A_{0})}}$${{F(U\setminus(A_{0}\cup A_{1}))}}$
is homotopy cartesian. This corresponds to the conventional usage of the term “excisivity,” hence justifying our terminology.
We also want to ensure that $F$ interacts well with the smooth structure on $M$, and also that $F$ is “continuous.” We thus make the following definition:
Definition 1.2.
[…]
  • Let $U\subset V$ be an inclusion of open sets of $M$ which is a smooth isotopy equivalence. Then the map $FV\to FU$ is a homotopy equivalence.
  • Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of open subsets of $M$. Then the map $F(\bigcup_{i\geq 0}U_{i})\to\operatorname{holim}_{i}F\left(U_{i}\right)$ is a homotopy equivalence.
With these definitions, we can now define polynomial approximations.
Definition 1.3.
Let $F,G:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be functors that carry smooth isotopy equivalences to homotopy equivalences. A natural transformation $\alpha:F\to G$ is said to exhibit $G$ as a $k$th polynomial approximation of $F$ if $G$ is polynomial of degree $\leq k$ and $\alpha$ is homotopically initial among the natural transformations to good polynomial functors of degree $\leq k$. In this case, we will write $G=T_{k}F$.
At this point, several questions come up naturally:
Question 1.4.
When does a good presheaf on $M$ admit polynomial approximations? If it does, how do we construct the approximations?
Question 1.5.
Can we classify polynomial functors?
Suppose now that $F$ takes values in the category of pointed spaces and that we have constructed its polynomial approximations. In good cases, we can expect that the analysis of $F$ reduces, up to extension problem, to that of the fibers $L_{k}F=\mathrm{hofib}\left(T_{k}F\to T_{k-1}F\right)$. The functor $L_{k}F$ has the special property that it is homogeneous: It is polynomial of degree $\leq k$, and its $\left(k-1\right)$th polynomial approximation vanishes. This leads to the following question:
Question 1.6.
Can we classify homogeneous functors?
And of course, we want to know:
Question 1.7.
Is manifold calculus actually useful?
Numerous studies suggest that the answer to Question 1.7 is an unequivocal yes: The idea of manifold calculus is a very, very good idea. Rather than making a long list of applications, we refer the reader to Munson’s survey [Mun10]. As for the remaining questions, Weiss gave complete answers in [Wei99], which we now review as two separate theorems (Theorems 1.8 and 1.9).
The first theorem answers Questions 1.4 and 1.5. We will let $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)$ denote the subset of $\operatorname{Open}\left(M\right)$ consisting of the elements that are diffeomorphic to $\mathbb{R}^{n}\times S$, where $S$ is a finite set of cardinality at most $k$.
Theorem 1.8.
[Wei99, Lemma 3.8, Theorem 4.1, Theorem 5.1] Let $G:\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be a functor which carries isotopy equivalences to homotopy equivalences. The homotopy right Kan extension of $G$ along the inclusion $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$ is good and $k$-excisive. Moreover, every good polynomial functor of degree $\leq k$ arises in this way.
In particular, every good space-valued presheaf $F$ on $M$ admits polynomial approximations in all degrees, constructed by homotopy right Kan extensions.
And the second theorem addresses Question 1.6:
Theorem 1.9.
[Wei99, Theorem 8.5] Homogeneous functors of degree $k$ taking values in pointed spaces admit a classification in terms of fibrations over $B_{k}\left(M\right)$ equipped with sections, where $B_{k}\left(M\right)$ denotes the space of unordered configurations of $k$ points in $M$.
Now notice that everything we have discussed so far makes sense for presheaves taking values in categories other than that of spaces. Moreover, it seems quite natural to consider this generalization. For example, if one is interested in (co)homology, then it is not hard to imagine that calculi of presheaves of (co)chain complexes or spectra would be very useful. As such, numerous studies consider such generalized manifold calculus. But none of them seem to address Questions 1.4, 1.5, and 1.6 in this generalized setting. We will fix this in this paper.
Here is what we will do. We will generalize Theorems 1.8 and 1.9 to presheaves on $M$ taking values in an arbitrary $\infty$-category $\mathcal{C}$ with small limits. It is straightforward to state a generalization of Theorem 1.8: The definitions of good functors, polynomial functors, polynomial approximations, and homogeneous functors directly carry over to $\mathcal{C}$-valued presheaves; we just have to replace $\operatorname{Open}\left(M\right)$ by its nerve $N\left(\operatorname{Open}\left(M\right)\right)$ and homotopy limits by $\infty$-categorical limits. So a generalization of Theorem 1.8 will be as follows.
Theorem 1.10 (Theorem 3.4).
The right Kan extension functor restricts to a categorical equivalence
$\operatorname{Fun}_{{\rm istp}}\left(N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}},\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
where:
  • $\operatorname{Fun}_{{\rm istp}}\left(N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}},\mathcal{C}\right)$ denotes the $\infty$-category of functors $N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ which carries smooth isotopy equivalences to equivalences; and
  • $\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of $k$-excisive good functors $N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$.
In particular, every good functor $F:N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ admits a $k$th polynomial approximation, which is the right Kan extension of $F|N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}}$.
To generalize Theorem 1.9, notice that the datum of a fibration over $B_{k}\left(M\right)$ equipped with a distinguished section is equivalent (via the $\infty$-categorical Grothendieck construction [Lur09, 2.2.1.2]) to that of a functor $\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{S}_{\ast}$, where $\mathcal{S}_{\ast}$ denotes the $\infty$-category of pointed spaces. With this in mind, we generalize Theorem 1.9 as follows:
Theorem 1.11 (Corollary 4.5).
Suppose that $\mathcal{C}$ is pointed. There is an equivalence of $\infty$-categories
$\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\simeq\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
where $\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of good functors $N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ that are of homogeneous of degree $k$.
Remark 1.12.
Notice that most of the statements of Theorems 1.10 and 1.11 make sense even when $M$ does not have a smooth structure. Because of this, we will work mostly with topological manifolds in the main body of the paper, indicating how various statements will be affected in the presence of smooth structures here and there.
Remark 1.13.
In [BdBW13], Boavida de Brito and Weiss established an analog of Theorem 1.9 for context-free [^1] The word “context” here has no relation with the title of this paper. manifold calculus, i.e., calculus of functors defined on the category of all smooth manifolds of a fixed dimension, not just on the poset of open sets on a single manifold. We will also state and prove a generalization of this in Section 5.
Remark 1.14.
In [STS19], Songhafouo Tsopméné and Stanley considered the classification of homogeneous functors in the case where $\mathcal{C}$ is the nerve of an ordinary category. Their result is a special case of Theorem 1.11, although our proof method is quite different from theirs.
At first glance, one may think that proofs of Theorem 1.10 and 1.11 are straightforward, using a minor modification of Weiss’s argument. However, Weiss’s original argument does not seem to generalize to the case at hand, for it relies critically on concrete constructions of spaces. We will get around this problem by using a deep theorem on localizations with respect to isotopy equivalences, due to Ayala and Francis [AF15, Proposition 2.19] [^2] Since the original proof of [AF15, Proposition 2.19] is missing some important justifications, we will record a detailed proof in Subsection 2.3. The author learned the proof from David Ayala and is grateful to him for explaining the proof..
Theorem 1.10 and 1.11 have a wide array of applications outside what has traditionally been regarded as part of manifold calculus. For example, (part of) the theory of factorization homology [AF20] deals with functors of the form
$H:\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}\to\mathcal{C},$
where $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denotes the $\infty$-category of smooth $n$-manifolds and smooth embeddings. By precomposing the functor $N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, we obtain a $\mathcal{C}^{\mathrm{op}}$-valued precosheaf $F:N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}^{\mathrm{op}}$. An observation of Ayala and Francis [AF20, Corollary 2.2.24] (which will be reproved in Subsection 3.4) essentially asserts that the functor $F$ is the limit of its Taylor tower, in the sense that the map $F\to\operatorname{lim}_{k}T_{k}F$ is an equivalence. This means that invariants of manifolds realized by factorization homology, which include very interesting class of invariants [AF20], can be approximated by their Taylor approximations. As another example, we will apply Theorem 1.11 in our future work to a functor taking values in the $\infty$-category $\mathcal{O}\mathsf{p}$ of $\infty$-operads. We will see that Theorem 1.11 is just the right tool to establish that this functor enjoys a “local-to-global” property. In light of the ubiquity of pre(co)sheaves on manifolds, we expect that there will be many other applications.
Outline of the Paper
In Section 2, we collect miscellaneous results that will be used in later sections. The statements of the results in later sections can be understood without reading this section, so we encourage readers to read this section only when the need arises. Section 3 concerns the existence and classification of polynomial functors, as well as the convergence of Taylor towers. Section 4 classifies homogeneous functors. In Section 5, we prove a generalization of the results in Section 3 for context-free manifold calculus. Appendix A contains a discussion of the context-free analog of the localization theorem of Ayala and Francis. The results in the appendix will not be used elsewhere, but it has some independent interest.
Acknowledgment
I appreciate David Ayala for kindly sharing the proof of Theorem 2.24, which plays a central role in this paper. Alexander Kupers generously answered my question on MathOverflow [sku]. His answer is presented as Proposition 2.7. I am also indebted to Daisuke Kishimoto and Mitsunobu Tsutaya for their constant support and encouragement.
Notation and Terminology
  • We say that a simplicial category $\mathcal{C}$ is locally Kan if its hom-simplicial sets are Kan complexes. A full simplicial subcategory $\mathcal{C}^{0}\subset\mathcal{C}$ is homotopically replete if $X\to Y$ is a homotopy equivalence of $\mathcal{C}$ and $X\in\mathcal{C}^{0}$, then $Y\in\mathcal{C}^{0}$.
  • A functor $\mathcal{C}\to\mathcal{D}$ of simplicial categories is called a weak equivalence if it induces a categorical equivalence $\pi_{0}\left(\mathcal{C}\right)\to\pi_{0}\left(\mathcal{D}\right)$ and a weak homotopy equivalence $\mathcal{C}\left(X,Y\right)\to\mathcal{D}\left(fX,fY\right)$ for every pair of objects $X,Y\in\mathcal{C}$.
  • The term $\infty$-category” will be used as a synonym of quasi-category of Joyal [Joy02]. We will mainly follow [Lur09] for notation and terminology for $\infty$-categories, with the following exceptions:
    • We say that a morphism $f:S\to T$ of simplicial sets is final if it is cofinal in the sense of [Lur09, Definition 4.1.1.1 ]. We say that $f$ is initial if its opposite is final.
    • We say that a functor $f:\mathcal{C}\to\mathcal{D}$ of $\infty$-categories is a localization with respect to a set $S$ of morphisms of $\mathcal{C}$ if it is initial among functors inverting the morphisms in $S$. (See Subsection 2.3 for a precise definition.) This definition is more general than the definition of [Lur09].
  • If $\mathcal{C}$ is an $\infty$-category, we denote its maximal sub Kan complex by $\mathcal{C}^{\simeq}$ and refer to it as the core of $\mathcal{C}$.
  • Let $n\geq 0$. An $n$-manifold, or a manifold of dimension $n$, will always mean a topological manifold without boundary, i.e., a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$. If $n\geq 1$, we define an $n$-manifold with boundary to be a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$ or $[0,\infty)\times\mathbb{R}^{n-1}$. In particular, not every $n$-manifold with boundary is an $n$-manifold.
  • Let $n\geq 0$. Given $n$-manifolds $M$ and $N$, we let $\operatorname{Emb}\left(M,N\right)$ denote the topological space of embeddings $M\to N$, topologized by the compact-open topology. We let $\mathcal{M}\mathsf{fld}_{n}^{T}$ denote the topological category whose objects are $n$-manifolds and whose hom spaces are given by $\operatorname{Emb}\left(-,-\right)$, and let $\mathcal{M}\mathsf{fld}_{n}^{\Delta}$ denote the simplicial category obtained from $\mathcal{M}\mathsf{fld}_{n}^{T}$ by applying the singular complex functor to the hom spaces. We let $\mathcal{M}\mathsf{fld}_{n}$ denote the homotopy coherent nerve of the topological category $\mathcal{M}\mathsf{fld}_{n}^{T}$, and let $\mathsf{Mfld}_{n}$ denote the nerve of the the ordinary category obtained from $\mathcal{M}\mathsf{fld}_{n}^{T}$ by forgetting the topology of the mapping spaces. Equivalences of $\mathcal{M}\mathsf{fld}_{n}$ are called isotopy equivalences.
    We let $\mathcal{D}\mathsf{isk}_{n}\subset\mathcal{M}\mathsf{fld}_{n}$ and $\mathsf{Disk}_{n}\subset\mathsf{Mfld}_{n}$ denote the full subcategories spanned by the objects that are homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$. For each $k\geq 0$, we let $\mathcal{D}\mathsf{isk}_{n}^{\leq k}\subset\mathcal{D}\mathsf{isk}_{n}$ and $\mathsf{Disk}_{n}^{\leq k}\subset\mathsf{Disk}_{n}$ denote the full subcategories spanned by the objects with at most $k$ components.
  • Let $n\geq 0$. We let $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the homotopy coherent nerve of the topological category of smooth $n$-manifolds and smooth embeddings, whose mapping spaces $\operatorname{Emb}_{\mathrm{sm}}\left(-,-\right)$ are topologized by the weak topology (also called the compact-open $C^{\infty}$ topology) of [Hir76, Chapter 2]. We let $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}\subset\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the full subcategory spanned by the smooth manifolds diffeomorphic to a finite disjoint union of $\mathbb{R}^{n}$ (with the standard smooth structure). Equivalences of $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ will be called smooth isotopy equivalences. The $\infty$-categories $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, $\mathsf{Mfld}_{\mathrm{sm},n}$, $\mathsf{Disk}_{\mathrm{sm},n}$, and $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}$ are defined as in the previous point.
  • Let $n,k\geq 0$. If $M$ is an $n$-manifold, we will write $\operatorname{Disj}\left(M\right)\subset\operatorname{Open}\left(M\right)$ for the full subcategory spanned by the open sets homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$, and let $\operatorname{Disj}^{\leq k}\left(M\right)$ denote its full subposet spanned by the open sets with at most $k$ components. If $M$ is smooth, we let $\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\subset\operatorname{Disj}\left(M\right)$ denote the full subcategory spanned by open sets diffeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$ (with the standard smooth structure), and write $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)=\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\cap\operatorname{Disj}^{\leq k}\left(M\right)$.
  • Let $n\geq 0$ and let $M$ be an $n$-manifold. For each finite set $S$, we let $\operatorname{Conf}\left(S,M\right)\subset M^{S}$ denote the subspace consisting of the injections $S\to M$. For each $k\geq 0$, we will write $\operatorname{Conf}\left(k,M\right)=\operatorname{Conf}\left(\{1,\dots,k\},M\right)$. (Thus $\operatorname{Conf}\left(0,M\right)$ is a point.) The $k$th symmetric group $\Sigma_{k}$ acts on $\operatorname{Conf}\left(k,M\right)$ by precomposition. We will write $B_{k}\left(M\right)=\operatorname{Conf}\left(k,M\right)/\Sigma_{k}$ for the orbit space of this action. The points of $B_{k}\left(M\right)$ will be identified with subsets of $M$ of cardinality $k$.
2. Preliminaries
In this section, we develop machineries we will use in later sections. Subsection 2.1 deals with germ bundles, which is an analog of tangent bundles in the topological setting. In Subsection 2.2, we consider the relation between Kan extensions and colimit diagrams, using Lurie’s path fibrations. In Subsection 2.3, we review Ayala and Francis’s localization theorems for isotopy equivalences. Finally, in Subsection 2.4, we consider the problem of zero extensions of functors.
The first three subsections are written in a linear order. The final subsection is independent of the other subsections.
2.1. Germs
[…]
Definition 2.1.
[Lur17, Definition 5.4.1.6] Let $n\geq 0$. For each positive number $r>0$, let $B^{n}\left(r\right)\subset\mathbb{R}^{n}$ denote the open ball of radius $r$ centered at the origin. Given a finite set $S$ and an $n$-manifold $M$, we define the simplicial set $\operatorname{Germ}\left(S,M\right)$ of $S$-germs of $M$ as the colimit
$\operatorname{Germ}\left(S,M\right)=\operatorname{colim}_{r\in\left(\mathbb{R}_{>0}\right)^{\mathrm{op}}}\operatorname{Sing}\operatorname{Emb}\left(S\times B^{n}\left(r\right),M\right).$
If $S$ is a singleton, we simply write $\operatorname{Germ}\left(S,M\right)=\operatorname{Germ}\left(M\right)$. The evaluation at the origin determines a map $\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$. For each $p\in M$, we set $\operatorname{Germ}_{p}\left(M\right)=\operatorname{Germ}\left(M\right)\times_{\operatorname{Sing}M}\{p\}$.
Remark 2.2.
Let $n\geq 0$. The simplicial set $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$ has the structure of a simplicial monoid. To see this, note that its $k$-simplex is represented by a map $\sigma:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to\mathbb{R}^{n}$, where $r>0$ is a positive number and for each point $x\in\left|\Delta^{k}\right|$, the map $\sigma\left(x,-\right):B^{n}\left(r\right)\to\mathbb{R}^{n}$ is an embedding which fixes the origin. Given another such map $\tau:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to\mathbb{R}^{n}$, the product $[\tau]\circ[\sigma]$ is represented by the map
[…]
where $0<r^{\prime\prime}<r$ is chosen so that $\sigma|\Delta^{k}\times B^{n}\left(r^{\prime\prime}\right)$ takes values in $B^{n}\left(r^{\prime}\right)$. (Such a number $r^{\prime\prime}$ exists because $\left|\Delta^{k}\right|$ is compact.)
More generally, if $M$ is an $n$-manifold and $S$ is a finite set, then $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$ acts on $\operatorname{Germ}\left(S,M\right)$ (from the right) as follows: A $k$-simplex of $\operatorname{Germ}\left(S,M\right)\times\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$ is represented by maps $\tau:\left|\Delta^{k}\right|\times\coprod_{s\in S}B^{n}\left(r\right)\to M$ and $\left(\sigma_{s}:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to\mathbb{R}^{n}\right)_{s\in S}$ where $r,r^{\prime}>0$. Its image in $\operatorname{Germ}\left(S,M\right)$ is represented by the composite
[…]
where $r^{\prime\prime}>0$ is chosen so that the above composite is defined.
Remark 2.3.
Let $n\geq 0$, let $M$ be an $n$-manifold, and let $S$ be a finite set. As a filtered colimit of Kan complexes, the simplicial set $\operatorname{Germ}\left(S,M\right)$ is a Kan complex. Moreover, the map
$\operatorname{Sing}\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\to\operatorname{Germ}\left(S,M\right)$
is a homotopy equivalence of Kan complexes [Lur17, Proposition 5.4.1.8].
The importance of the germ construction lies in the following propositions:
Proposition 2.4.
Let $n\geq 0$. The simplicial monoid $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$ is a simplicial group.
Proposition 2.5.
Let $n\geq 0$, let $M$ be an $n$-manifold, and let $S$ be a finite set. The map
$\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$
is a principal $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$-fibration.
The proofs of Propositions 2.4 and 2.5 require a few preliminaries.
Lemma 2.6.
Let $n\geq 1$, let $X$ be a locally path-connected topological space, let $M$ be an $n$-manifold, and let
$f,g:X\times D^{n}\to M$
be continuous maps. Suppose that, for each $x\in X$, the map $f_{x}=f\left(x,-\right):D^{n}\to M$ is injective. If there is a point $x_{0}\in X$ such that $g_{x_{0}}\left(D^{n}\right)\subset f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$, then there is a neighborhood $U$ of $x_{0}$ such that $g_{x}\left(D^{n}\right)\subset f_{x}\left(\operatorname{Int}D^{n}\right)$ for all $x\in U$.
Proof.
Since $f_{x_{0}}|\operatorname{Int}D^{n}$ is an injective continuous map between $n$-manifolds, it is open. In particular, the set $V=f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$ is open. Thus, replacing $X$ by a neighborhood of $x_{0}$ if necessary, we may assume that $g$ takes values in $f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$. The subset $f_{x_{0}}^{-1}\left(g_{x_{0}}\left(D^{n}\right)\right)\subset D^{n}$ is compact and lies in the interior of $D^{n}$, so we can find some number $10<r<$ 1 such that $f_{x_{0}}^{-1}\left(g_{x_{0}}\left(D^{n}\right)\right)\subset B^{n}\left(r\right)$. Replacing $X$ by a neighborhood of $x_{0}$ if necessary, we may assume that $f\left(X\times\overline{B^{n}\left(r\right)}\right)\subset V$.
Choose a metric on $V$ and set $\varepsilon=\operatorname{dist}\left(V\setminus f_{x_{0}}\left(B^{n}\left(r\right)\right),g{}_{x_{0}}\left(D^{n}\right)\right)$. Since $D^{n}$ and $\partial\overline{B^{n}\left(r\right)}$ are compact, and since $X$ is locally path-connected, there is a path-connected neighborhood $U$ of $x_{0}$ such that
$\sup_{x\in U}\operatorname{dist}\left(f_{x}\left(\partial\overline{B^{n}\left(r\right)}\right),f_{x_{0}}\left(\partial\overline{B^{n}\left(r\right)}\right)\right)$ $<\varepsilon/2,$
$\sup_{x\in U}\operatorname{dist}\left(g_{x}\left(D^{n}\right),g_{x_{0}}\left(D^{n}\right)\right)$ $<\varepsilon/2.$
We claim that $U$ has the desired properties. In fact, we will show that $g_{x}\left(D^{n}\right)\subset f_{x}\left(B^{n}\left(r\right)\right)$ for all $x\in U$.
Let $x_{1}\in U$. We will derive a contradiction by assuming that there is a point $p\in D^{n}$ such that $q=g_{x_{1}}\left(p\right)\not\in f_{x_{1}}\left(B^{n}\left(r\right)\right)$. By construction, for each $x\in U$, the point $q$ does not belong to $f_{x}\left(\partial\overline{B^{n}\left(r\right)}\right)$. Since $U$ is path-connected, the maps $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ and $f_{x_{1}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ are homotopic. The latter map is null homotopic, for it extends to all of $\overline{B^{n}\left(r\right)}$. Hence the map $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ is also nullhomotopic. On the other hand, the distance from $q$ to $g_{x_{0}}\left(D^{n}\right)$ is less than $\varepsilon/2$, so $q$ belongs to $f_{x_{0}}\left(B_{n}\left(r\right)\right)$. It follows that the map $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ is a homotopy equivalence. This is a contradiction, for $\partial\overline{B^{n}\left(r\right)}$ is not contractible. ∎
Proposition 2.7.
Let $n\geq 0$. For any $n$-manifold $M$ and any finite set $S$, the evaluation at the origin determines a Serre fibration
$\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\to\operatorname{Conf}\left(S,M\right).$
The following proof of Proposition 2.7 is due to Alexander Kupers [sku].
Proof.
Since Serre fibrations can be recognized locally [tD08, 6.3.3], it suffices to prove the following:
  • (A)
    For any collection $\{U_{s}\}_{s\in S}$ of pairwise disjoint open sets of $M$ such that each $U_{s}$ is homeomorphic to $\mathbb{R}^{n}$, the map
    $\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\times_{\operatorname{Conf}\left(S,M\right)}\prod_{s\in S}U_{s}\to\prod_{s\in S}U_{s}$
    is a Serre fibration.
We begin with a preliminary assertion. Let $\operatorname{Homeo}\left(M\right)\subset\operatorname{Emb}\left(M,M\right)$ denote the subspace consisting of the self-homeomorphisms of $M$. We prove the following:
  • (B)
    Let $U\subset M$ be an open set homeomorphic to $\mathbb{R}^{n}$ and let $K\subset U$ be a compact subset. There is a continuous map
    $\chi:K\times K\to\operatorname{Homeo}\left(M\right)$
    which has the following properties:
    1. $\bullet$
      For each $\left(x,y\right)\in K\times K$, the support of $\chi\left(x,y\right)$ is compact and lies in $U$ and, and moreover $\chi\left(x,y\right)\left(x\right)=y$.
    2. $\bullet$
      For each $x\in K$, we have $\chi\left(x,x\right)=\operatorname{id}_{\mathbb{R}^{n}}$.
To prove (B), it suffices to consider the case where $U=M=\mathbb{R}^{n}$ and $K=D^{n}$. Choose a compactly supported smooth function $\phi:\mathbb{R}\to\mathbb{R}$ such that $\phi\left(t\right)=1$ if $\left|t\right|\leq 1$ and $\sup_{t\in\mathbb{R}}\phi^{\prime}\left(t\right)>-1/2$. We define $\chi:D^{n}\times D^{n}\to\operatorname{Homeo}\left(\mathbb{R}^{n}\right)$ by
$\chi\left(x,y\right)\left(z\right)=\left(\left(1-\phi\left(z_{i}\right)\right)z_{i}+\phi\left(z_{i}\right)\left(z_{i}+y_{i}-x_{i}\right)\right)_{i=1}^{n}.$
Note that $\chi\left(x,y\right)$ is a homeomorphism because it is the product of self-homeomorphism of $\mathbb{R}$ (which is ensured by the choice of $\phi$).
Now we prove assertion (A). Let $k\geq 0$ and consider a solid commutative diagram
[…]
We wish to find a map $I^{k}\times I\to\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\times_{\operatorname{Conf}\left(S,M\right)}\prod_{s\in S}U_{s}$ rendering the diagram commutative. Using assertion (B), we can construct a map
$\Phi:I^{k}\times I\to\operatorname{Homeo}\left(M\right)$
such that $\Phi\left(x,0\right)=\operatorname{id}_{M}$ and $\Phi\left(x,t\right)\left(G_{s}\left(x,0\right)\right)=G_{s}\left(x,t\right)$ for all $s\in S$ and $\left(x,t\right)\in I^{k}\times I$; here $G_{s}$ denotes the $s$th component of $G$. The map
$I^{k}\times I\to\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right),\,\left(x,t\right)\mapsto\Phi\left(x,t\right)\circ F\left(x,0\right)$
gives the desired filler. ∎
Corollary 2.8.
Let $n\geq 0$. For any $n$-manifold $M$ and any finite set $S$, the map
$\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$
is a Kan fibration.
Proof.
This follows from Proposition 2.7, for Kan fibrations are stable under filtered colimits. ∎
We can now prove Propositions 2.4 and 2.5.
Proof of Proposition 2.4.
The claim is trivial if $n=0$, so we will assume that $n\geq 1$. Let $\sigma:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to\mathbb{R}^{n}$ be a representative of a $k$-simplex $[\sigma]\in\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$. We wish to construct an inverse of $[\sigma]$.
For each $x\in\left|\Delta^{k}\right|$, let $\sigma_{x}=\sigma\left(x,-\right)$ denote the restriction of $\sigma$. We will first show that there is some $s>0$ which satisfies $\overline{B^{n}\left(s\right)}\subset\sigma_{x}\left(\overline{B^{n}\left(r/2\right)}\right)$ for every $x\in\left|\Delta^{k}\right|$. Since $\left|\Delta^{k}\right|$ is compact, it will suffice to prove the following:
  • ($\ast$)
    For each $x_{0}\in\left|\Delta^{k}\right|$, there are neighborhoods $U_{x_{0}}$ of $x_{0}$ and some number $s\left(x_{0}\right)>0$ such that $\overline{B^{n}\left(s\left(x_{0}\right)\right)}\subset\sigma_{x}\left(\overline{B^{n}\left(r/2\right)}\right)$ for every $x\in U_{x_{0}}$.
Assertion ($\ast$) follows from Lemma 2.6.
Now we construct the inverse of $[\sigma]$. Choose $s>0$ as in the previous paragraph. The map
$F:\sigma^{-1}\left(\overline{B^{n}\left(s\right)}\right)\to\left|\Delta^{k}\right|\times\overline{B^{n}\left(s\right)},\,\left(x,p\right)\mapsto\left(x,\sigma_{x}\left(p\right)\right)$
is a continuous bijection of compact Hausdorff spaces, so it is a homeomorphism. Consider the composite
$\tau:\left|\Delta^{k}\right|\times B^{n}\left(s\right)\xrightarrow{F^{-1}}\sigma^{-1}\left(\overline{B^{n}\left(s\right)}\right)\xrightarrow{\operatorname{pr}}\mathbb{R}^{n}.$
We have $[\tau][\sigma]=1=[\sigma][\tau]$ in $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}$, so $[\tau]$ is the desired inverse of $[\sigma]$. ∎
Proof of Proposition 2.5.
Since the map $\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\left(\operatorname{Conf}\left(S,M\right)\right)$ is a pullback of the map $\operatorname{Germ}\left(M\right)^{S}\to\operatorname{Sing}M^{S}$, it suffices to consider the case where $S$ is a singleton. Let $k\geq 0$. The action of $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}$ on $\operatorname{Germ}\left(M\right)_{k}$ is clearly free. It will therefore suffice to show that the induced map
$\theta:\operatorname{Germ}\left(M\right)_{k}/\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}\to\operatorname{Sing}\left(M\right)_{k}$
is bijective. Surjectivity of $\theta$ is immediate from Corollary 2.8. To prove that $\theta$ is injective, we must prove the following:
  • ($\ast$)
    Let $r>0$ and let $\sigma,\tau:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to M$ be continuous maps such that, for each $x\in\left|\Delta^{k}\right|$, the maps $\sigma_{x}=\sigma\left(x,0\right):B^{n}\left(r\right)\to M$ and $\tau_{x}=\tau\left(x,0\right):B^{n}\left(r\right)\to M$ are embeddings satisfying $\sigma_{x}\left(0\right)=\tau_{x}\left(0\right)$. Then there are some number $r^{\prime}>0$ and a map $g:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to B^{n}\left(r\right)$ such that
    $\sigma\left(x,g\left(x,p\right)\right)=\tau\left(x,p\right)$
    for every $\left(x,p\right)\in\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)$.
Using Lemma 2.6 and the compactness of $\left|\Delta^{k}\right|$, we can find some $0<r^{\prime}<r$ which satisfies $\tau_{x}\left(\overline{B^{n}\left(r^{\prime}\right)}\right)\subset\sigma_{x}\left(B^{n}\left(r/2\right)\right)$ for every $x\in\left|\Delta^{k}\right|$. The map
$F:\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}$ $\to\left|\Delta^{k}\right|\times M$
$\left(x,p\right)$ $\mapsto\left(x,\sigma_{x}\left(p\right)\right)$
is a continuous injection from a compact space to a Hausdorff space, so it is a homeomorphism onto its image. We define $g:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to B^{n}\left(r\right)$ as the composite
$\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\xrightarrow{\tau^{\prime}}F\left(\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}\right)\xrightarrow{F^{-1}}\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}\xrightarrow{\operatorname{pr}}B^{n}\left(r\right),$
where $\tau^{\prime}\left(x,p\right)=\left(x,\tau_{x}\left(p\right)\right)$. The map $g$ has the desired properties. ∎
We conclude this subsection with another consequence of Corollary 2.8.
Proposition 2.9.
Let $n\geq 0$, let $S$ be a finite set, let $M$ and $N$ be $n$-manifolds, and let $\phi:M\to N$ be an embedding. The square
${{\operatorname{Emb}(\mathbb{R}^{n}\times S,M)}}$${{\operatorname{Emb}(\mathbb{R}^{n}\times S,N)}}$${{\operatorname{Conf}(S,M)}}$${{\operatorname{Conf}(S,N)}}$
determined by the evaluation at the origin, is homotopy cartesian (with respect to the Quillen model structure for topological spaces).
Proof.
It will suffice to show that the square becomes homotopy cartesian after applying the singular complex functor. Using Remark 2.3, we are reduced to showing that the square
${{\operatorname{Germ}(S,M)}}$${{\operatorname{Germ}(S,N)}}$${{\operatorname{Conf}(S,M)}}$${{\operatorname{Conf}(S,N)}}$
is homotopy cartesian. By Corollary 2.8, the vertical arrows of this diagram are Kan fibrations. It will therefore suffice to show that the square is strictly cartesian, which is clear. ∎
Remark 2.10.
Proposition 2.9 remains valid if $\phi$ is a smooth embedding of smooth manifolds of the same dimension and $\operatorname{Emb}$ is replaced by $\operatorname{Emb}_{\mathrm{sm}}$. This is because the smooth version of Proposition 2.7 is also true, with the same proof.
2.2. Path Fibrations and Kan Extensions
Consider the following problem: Suppose we are given a functor $f:\mathcal{C}\to\mathcal{D}$ of $\infty$-categories and a full subcategory $\mathcal{C}_{0}\subset\mathcal{C}$ such that $f$ is a left Kan extension of $f|\mathcal{C}_{0}$. Suppose we are given a diagram $\overline{g}:K^{\triangleright}\to\mathcal{C}$. Under what condition is the composite diagram $f\overline{g}:K^{\triangleright}\to\mathcal{C}$ a colimit diagram? In this section, we will give a sufficient condition, using path fibrations.
We start by recalling the definition of oriented fiber products.
Definition 2.11.
[Lur24, Tag 01KF] The oriented fiber product of maps of simplicial sets $X\to Z\leftarrow Y$ is the simplicial set
$X\overrightarrow{\times}_{Z}Y=X\times_{\operatorname{Fun}\left(\{0\},Z\right)}\operatorname{Fun}\left(\Delta^{1},Z\right)\times_{\operatorname{Fun}\left(\{1\},Z\right)}Y.$
[…]
Proposition 2.12.
[Lur24, Tag 0478] Let $\mathcal{A}\to\mathcal{C}\leftarrow K$ be maps of simplicial sets, where $\mathcal{A}$ and $\mathcal{C}$ are $\infty$-categories. The map
$q:\mathcal{A}\overrightarrow{\times}_{\mathcal{C}}K\to K$
is a cocartesian fibration, and a morphism of $\mathcal{A}\overrightarrow{\times}_{\mathcal{C}}K$ is $q$-cocartesian if and only if its image in $\mathcal{A}$ is an equivalence.
Our criterion will rely on the following construction involving the oriented fiber product.
Construction 2.13.
Let $g:K\to\mathcal{C}$ be a map of simplicial sets, where $\mathcal{C}$ is an $\infty$-category, and let $\mathcal{C}_{0}\subset\mathcal{C}$ be a full subcategory. Let $C\in\mathcal{C}$ be an object and suppose we are given a natural transformation $\alpha:g\to\underline{C}$. Let $p:\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to\mathcal{C}_{0}$ and $q:\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to K$ denote the projections. We define a map
$\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}\{C\}=\mathcal{C}_{0}^{/C}$
as follows: The evaluation map $\operatorname{Fun}\left(\Delta^{1},\mathcal{C}\right)\times\Delta^{1}\to\mathcal{C}$ induces a natural transformation $\beta:p\to gq$. The following diagrams summarizes the situation so far:
[…]
Let $\gamma$ denote a composition
[…]
in the $\infty$-category $\operatorname{Fun}\left(\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K,\mathcal{C}\right)$. The natural transformation $\gamma$ determines a map $\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}\{C\}=\mathcal{C}_{0}^{/C}$.
Remark 2.14.
In the situation of Construction 2.13, if $p_{X}:X\to\mathcal{C}_{0}$ is a map of simplicial sets, there is an isomorphism of simplicial sets
$\operatorname{Hom}_{\operatorname{Fun}\left(X,\mathcal{C}\right)}\left(p_{X},\underline{C}\right)\cong\operatorname{Fun}_{\mathcal{C}_{0}}\left(X,\mathcal{C}_{0}^{/C}\right).$
Consequently, the map $\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to\mathcal{C}_{0}^{/C}$ is well-defined up to natural equivalence over $\mathcal{C}_{0}$. Also, if we are given a commutative diagram of $\infty$-categories
${{\mathcal{C}_{0}}}$${{\mathcal{D}_{0}}}$${{\mathcal{C}}}$${{\mathcal{D}}}$
where $\mathcal{D}_{0}\subset\mathcal{D}$ is a full subcategory, the induced diagram
${{\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K}}$${{(\mathcal{C}_{0})^{/C}}}$${{\mathcal{D}_{0}\overrightarrow{\times}_{\mathcal{D}}K}}$${{(\mathcal{D}_{0})^{/D}}}$
commutes up to natural equivalence. Here $D$ denotes the image of $C$ in $\mathcal{D}$.
We now come to the main result of this section.
Proposition 2.15.
[…]
The proof of Proposition 2.15 requires a lemma.
Lemma 2.16.
[Lur24, Tag 02ZM] Let $\mathcal{C},\mathcal{D},$ and $\mathcal{E}$ be $\infty$-categories. Let $p:\mathcal{C}\to\mathcal{D}$, $f:\mathcal{C}\to\mathcal{E}$, and $l:\mathcal{D}\to\mathcal{E}$ be functors, and let $\alpha:f\to lp$ be a natural transformation. Suppose that $p$ is a cocartesian fibration. The following conditions are equivalent:
  1. (1)
    The natural transformation $\alpha$ exhibits $l$ as a left Kan extension of $f$ along $p$.
  2. (2)
    For each object $D\in\mathcal{D}$, the map $\alpha$ restricts to a colimit diagram $f|\mathcal{C}_{D}\to\underline{lp\left(D\right)}$, where we set $\mathcal{C}_{D}=\mathcal{C}\times_{\mathcal{D}}\{D\}$.
Proof of Proposition 2.15.
By [Lur09, Corollary 3.3.1.4], we may assume that $K$ is an $\infty$-category. Set $\mathcal{E}=\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K$. We will write $p:\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to K$ and $q:\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K\to K$ for the projections, and $\beta:p\to gq$ for the natural transformation obtained from the evaluation map $\operatorname{Fun}\left(\Delta^{1},\mathcal{C}\right)\times\Delta^{1}\to\mathcal{C}$. We let $\gamma:p\to\underline{C}$ denote a composition of $\alpha q:gq\to\underline{C}$ and $\beta:p\to gq$ in $\operatorname{Fun}\left(\mathcal{E},\mathcal{C}\right)$, and let $h:\mathcal{E}\to\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}\{C\}\cong\mathcal{C}_{0}^{/C}$ denote the induced map. The situation can be summarized by the following diagram:
${{\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K}}$${{\mathcal{C}_{0}}}$${{\mathcal{C}_{0}\overrightarrow{\times}_{\mathcal{C}}K}}$${K}$${{\mathcal{C}}}$${{\mathcal{D}}}$${{\mathcal{C}_{0}^{/C}}}$${{\mathcal{C}_{0}}}${${{\{C\}}}${${{\{C\}}}$${{\mathcal{C}.}}$$\scriptstyle{q}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{h}$$\scriptstyle{p}$$\scriptstyle{\alpha}$$\scriptstyle{\beta}$$\scriptstyle{\delta}$
By construction, we have $\delta h=\gamma$. Thus $f\gamma$ is a colimit diagram. Therefore, it will suffice to show that the natural transformation $f\beta:fp\to fgq$ exhibits $fg$ as a left Kan extension of $fp$ along $q$. According to Lemma 2.16, this amounts to showing that, for each vertex $v\in K$, the natural transformation $f\beta|\mathcal{E}_{v}:fp|\mathcal{E}_{v}\to\underline{fg\left(v\right)}$ is a colimit diagram. Under the isomorphism of simplicial sets $\mathcal{E}_{v}\cong\mathcal{C}_{0}^{/g\left(v\right)}$, we can identify the natural transformation $f\beta|\mathcal{E}_{v}$ with the natural transformation $f|\mathcal{C}_{0}^{/g\left(v\right)}\to fg\left(v\right)$ obtained from the evaluation map $\mathcal{C}_{0}^{/g\left(v\right)}\times\Delta^{1}\to\mathcal{C}$. So the claim is a consequence of the hypothesis that $f$ is a left Kan extension of $f|\mathcal{C}_{0}$. ∎
To apply Proposition 2.15, we need to know when a map into slice $\infty$-categories is final. The following proposition will be useful for this purpose.
Proposition 2.17.
Let $\mathcal{C}$ be a locally Kan simplicial category, let $\mathcal{C}_{0}\subset\mathcal{C}$ be a homotopically replete full simplicial subcategory, let $X\in\mathcal{C}$ be an object, and let $\mathcal{I}$ be a small category. Let $f:\mathcal{I}^{\triangleright}\to\mathcal{C}$ be a functor which carries $\mathcal{I}$ into $\mathcal{C}_{0}$ and the cone point to $X$. The following conditions are equivalent:
  1. (1)
    The functor $f^{\prime}:N\left(\mathcal{I}\right)\to N\left(\mathcal{C}_{0}\right)_{/X}$ which is adjoint to $f$ is final.
  2. (2)
    For each object $C\in\mathcal{C}_{0}$ which admits a morphism $C\to X$ in $\mathcal{C}$, the map
    $\operatorname{hocolim}_{I\in\mathcal{I}}\mathcal{C}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
    is a weak homotopy equivalence of simplicial sets.
For the proof of Proposition 2.17, we introduce a somewhat ad-hoc notion.
Definition 2.18.
[…]
  1. (1)
    The maps $i$ and $q$ are homotopy equivalences.
  2. (2)
    For every $C\in\mathcal{C}$, the maps
    $p_{\ast}:\mathcal{C}\left(C,A\right)\to\mathcal{C}\left(C,Y\right),\,j^{*}:\mathcal{C}\left(B,C\right)\to\mathcal{C}\left(X,C\right)$
    are Kan fibrations.
Given a simplicial category $\mathcal{C}$ and an object $X\in\mathcal{C}$, we define another simplicial category $\mathcal{C}_{/X}$ by declaring that its objects are the morphisms $C\to X$ of the underlying category of $\mathcal{C}$, with mapping spaces defined by the pullback
$\mathcal{C}_{/X}\left(f:C\to X,g:C^{\prime}\to X\right)=\mathcal{C}\left(C,C^{\prime}\right)\times_{\mathcal{C}\left(C,X\right)}\{f\}.$
The simplicial category $\mathcal{C}_{/X}$ does not usually get along well with the homotopy coherent nerve functor. (It may not even be locally Kan). However, the problem disappears in the presence of model factorizations:
Lemma 2.19.
Let $\mathcal{C}$ be a locally Kan simplicial cateogry with model factorizations, and let $Z\in\mathcal{C}$ be an object. Let $\mathcal{E}\subset\mathcal{C}_{/X}$ be the full simplicial subcategory spanned by the objects $C\to X$ such that, for each object $C^{\prime}\in\mathcal{C}$, the map
$\mathcal{C}\left(C^{\prime},C\right)\to\mathcal{C}\left(C^{\prime},X\right)$
is a Kan fibration. Then:
  1. (1)
    The simplicial category $\mathcal{E}$ is locally Kan.
  2. (2)
    The composite $N\left(\mathcal{E}\right)\to N\left(\mathcal{C}_{/X}\right)\to N\left(\mathcal{C}\right)_{/X}$ is an equivalence of $\infty$-categories.
Proof.
This follows from [Lur24, Tag 01ZT]. (Beware that the convention of homotopy coherent nerve in [Lur24] is different from that of [Lur09].) ∎
The following lemma says that in many cases, locally Kan simplicial categories can be replaced by a one which admits model factorizations:
Lemma 2.20.
For every locally Kan simplicial category $\mathcal{C}$, there are a locally Kan simplicial category $\mathcal{D}$ with model factorizations and a weak equivalence of simplicial categories $\mathcal{C}\xrightarrow{\simeq}\mathcal{D}$.
Proof.
Let $\mathbf{A}$ denote the simplicial category of simplicial functors $SS\mathcal{C}\to\SS$, equipped with the projective model structure. Since $\mathcal{C}$ is locally Kan, the Yoneda embedding $j:\mathcal{C}\to\mathbf{A}$ factors through the full simplicial subcategory $\mathbf{A}^{\circ}\subset\mathbf{A}$ of fibrant-cofibrant object. Let $\mathbf{B}\subset\mathbf{A}^{\circ}$ denote the full simplicial subcategory spanned by the objects that are homotopy equivalent to objects lying in the image of $j$. Then the simplicial functor $\mathcal{C}\to\mathbf{B}$ is a weak equivalence of simplicial categories, and $\mathbf{B}$ admits model factorizations. ∎
We now arrive at the proof of Proposition 2.17.
Proof of Proposition 2.17.
By Lemma 2.20, we may assume that $\mathcal{C}$ admits model factorizations. Define a full simplicial subcategory $\mathcal{E}\subset\mathcal{C}_{/X}$ as in Lemma 2.19 and let $\mathcal{E}_{0}\subset\mathcal{E}$ denote the full simplicial subcategory spanned by the objects $C\to X$ such that $C\in\mathcal{C}_{0}$. Since $\mathcal{C}_{0}$ is homotopically replete, the functor $N\left(\mathcal{E}_{0}\right)\to N\left(\mathcal{C}_{0}\right)_{/X}$ is a categorical equivalence. Using the functorial factorization of morphisms in $\mathcal{C}$, we may assume that $f^{\prime}$ factors as a composite
$N\left(\mathcal{I}\right)\xrightarrow{}N\left(\mathcal{E}_{0}\right)\xrightarrow{\simeq}N\left(\mathcal{C}_{0}\right)_{/X}.$
Condition (1) is equivalent to the condition that, for each object $\left(\alpha:C\to X\right)\in\mathcal{E}_{0}$, the $\infty$-category $N\left(\mathcal{I}\right)\times_{N\left(\mathcal{E}_{0}\right)}N\left(\mathcal{E}_{0}\right)_{\alpha/}$ is weakly contractible. According to [Lur09, Corollary 3.3.4.6] and [Lur09, Theorem 4.2.4.1], this is equivalent to the condition that $\operatorname{hocolim}_{I\in\mathcal{I}}\mathcal{C}_{/X}\left(\alpha,f\left(I\right)\right)$ is weakly contractible. Now for each $I\in\mathcal{I}$, the cartesian square
${{\mathcal{C}_{/X}(\alpha,f(I))}}$${{\mathcal{C}(C,f(I))}}${${{\{\alpha\}}}$${{\mathcal{C}(C,X)}}$
is homotopy cartesian, for the right vertical map is a Kan fibration. Since colimits in $\mathcal{S}$ are universal [Lur09, Lemma 6.1.3.14], it follows that $\operatorname{hocolim}_{I\in\mathcal{I}}\mathcal{C}_{/X}\left(\alpha,f\left(I\right)\right)$ has the weak homotopy type of the homotopy fiber of the map
$\operatorname{hocolim}_{I\in\mathcal{I}}\mathcal{C}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
over $\alpha$. The claim follows. ∎
For later reference, we record a variant of Proposition 2.17.
Notation 2.21.
Let $\mathcal{C}$ be a locally Kan simplicial category. We will write $\mathcal{C}^{\simeq}\subset\mathcal{C}$ for the simplicial subcategory with the same object as $\mathcal{C}$. If $X,Y\in\mathcal{C}$ are objects of $\mathcal{C}$, the hom-simplicial set $\mathcal{C}^{\simeq}\left(X,Y\right)\subset\mathcal{C}\left(X,Y\right)$ is the union of the components of $\mathcal{C}\left(X,Y\right)$ whose vertices are the homotopy equivalences from $X$ to $Y$.
Proposition 2.22.
Let $\mathcal{C}$ be a locally Kan simplicial category, let $\mathcal{C}_{0}\subset\mathcal{C}$ be a homotopically replete full simplicial subcategory, let $X\in\mathcal{C}$ be an object, and let $\mathcal{I}$ be a small category. Let $f:\mathcal{I}^{\triangleright}\to\mathcal{C}$ be a functor which carries $\mathcal{I}$ into $\mathcal{C}_{0}^{\simeq}$ and the cone point to $X$. The following conditions are equivalent:
  1. (1)
    The functor $f^{\prime}:N\left(\mathcal{I}\right)\to\left(N\left(\mathcal{C}_{0}\right)_{/X}\right)^{\simeq}$ which is adjoint to $f$ is a weak homotopy equivalence.
  2. (2)
    For each object $C\in\mathcal{C}_{0}$ which admits a morphism $C\to X$ in $\mathcal{C}$, the map
    $\operatorname{hocolim}_{I\in\mathcal{I}}\mathcal{C}^{\simeq}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
    is a weak homotopy equivalence of simplicial sets.
The proof of Proposition 2.22 is nearly identical to that of Proposition 2.17, noting that for a Kan complex $X$, a map $K\to X$ of simplicial sets is final if and only if it is a weak homotopy equivalence. Details are left to the reader.
We conclude this section by recording another consequence of Lemma 2.16:
Proposition 2.23.
Let
${{\mathcal{A}}}$${{\mathcal{B}}}$${{\mathcal{C}}}$$\scriptstyle{f}$$\scriptstyle{q}$$\scriptstyle{p}$
be a commutative diagram of $\infty$-categories. Suppose that $p$ and $q$ are cocartesian fibrations and that, for each object $C\in\mathcal{C}$, the map
$\mathcal{A}\times_{\mathcal{C}}\{C\}=\mathcal{A}_{C}\to\mathcal{B}_{C}$
induced by $f$ is final. Then $f$ is final.
Proof.
[…]
$\mathcal{S}^{g/}\to\mathcal{S}^{gf/}$
is a covariant equivalence over $\mathcal{S}$.
Find a functor $l:\mathcal{C}\to\mathcal{S}$ and a natural transformation $\alpha:g\to lq$ which exhibits $l$ as a left Kan extension of $g$ along $q$, and find also a Kan complex $K$ and a natural transformation $\beta:l\to\underline{K}$ which exhibits $K$ as a colimit of $l$. Then a composite natural transformation
[…]
exhibits $K$ as a colimit of $g$ (by the transitivity of Kan extensions[Lur24, Tag 031M]), and we wish to show that the natural transformation $\theta f:gf\to\underline{K}$ exhibits $K$ as a colimit of $gf$. For this, it suffices to show that the natural transformation $\alpha f:gf\to lp$ exhibits $l$ as a left Kan extension of $gf$ along $p$. By Lemma 2.16, we must show that, for each $C\in\mathcal{C}$, the natural transformation
$\alpha f|\mathcal{A}_{C}:gf|\mathcal{A}_{C}\to\underline{l\left(C\right)}$
exhibits $l\left(C\right)$ as a colimit of $gf|\mathcal{A}_{C}$. Since $f|\mathcal{A}_{C}:\mathcal{A}_{C}\to\mathcal{B}_{C}$ is final, we are reduced to showing that the natural transformation
$\alpha|\mathcal{B}_{C}:g|\mathcal{B}_{C}\to\underline{l\left(C\right)}$
is a colimit diagram. This follows from Lemma 2.16. ∎
2.3. Localizing with respect to Isotopy Equivalences
Let $L:\mathcal{C}\to\mathcal{D}$ be a functor of $\infty$-categories and let $S$ be a set of morphisms of $\mathcal{C}$. Recall ([Lan21, Definition 2.4.2]) that $L$ is said to exhibit $\mathcal{D}$ as a (Dwyer–Kan) localization of $\mathcal{C}$ with respect to $S$ it satisfies the following conditions:
  • The functor $L$ carries every morphism in $S$ to an equivalence.
  • For every $\infty$-category $\mathcal{E}$, the functor
    $\operatorname{Fun}\left(\mathcal{D},\mathcal{E}\right)\to\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{E}\right)$
    is a categorical equivalence, where $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{E}\right)$ denotes the full subcategory of $\operatorname{Fun}\left(\mathcal{C},\mathcal{E}\right)$ spanned by the functors $\mathcal{C}\to\mathcal{E}$ carrying every morphism in $S$ to an equivalence.
A recurring theme in homotopy theory is that many nontrivial $\infty$-categories arise as localizations of ordinary categories. A famous theorem of Ayala and Francis [AF15, Proposition 2.19] is one manifestation of this: It asserts that for each $n$-manifold $M$, the functor
$\mathsf{Disk}_{n/M}\to\mathcal{D}\mathsf{isk}_{n/M}$
exhibits $\mathcal{D}\mathsf{isk}_{n/M}$ as a localization of $\mathsf{Disk}_{n/M}$ with respect to isotopy equivalences. The goal of this subsection is to present a detailed proof of their theorem, which the author learned from David Ayala.
For later applications, we will also need a variant of Ayala and Francis’s theorem. We state both of them as a single theorem.
Theorem 2.24.
Let $n\geq 0$ and let $M$ be an $n$-manifold.
  1. (1)
    The functor $\mathsf{Disk}_{n/M}\to\mathcal{D}\mathsf{isk}_{n/M}$ exhibits $\mathcal{D}\mathsf{isk}_{n/M}$ as a localization of $\mathsf{Disk}_{n/M}$ with respect to isotopy equivalences.
  2. (2)
    For every $k\geq 0$, the functor $\mathsf{Disk}_{n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ exhibits $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ as a localization of $\mathsf{Disk}_{n/M}^{\leq k}$ with respect to isotopy equivalences.
We will establish Theorem 2.24 after a few preliminaries. We begin with a variation of Quillen’s Theorem B.
Definition 2.25.
[Cis19, Definition 4.6.3] A map $E\to B$ of simplicial sets is said to be locally constant if for each morphism $X\to B$ of simplicial sets, the square
${{X\times_{B}E}}$${E}$${X}$${B}$
is homotopy cartesian in the Kan–Quillen model structure.
Proposition 2.26 (A Variation of Quillen’s Theorem B).
Let $\pi:E\to B$ be a cocartesian fibration of simplicial sets. The following conditions are equivalent:
  1. (1)
    The map $\pi$ is locally constant.
  2. (2)
    For each edge $\alpha:b\to b^{\prime}$ in $B$, the induced functor
    $\alpha_{!}:E_{b}=E\times_{B}\{b\}\to E_{b^{\prime}}$
    is a weak homotopy equivalence.
Proof.
[…]
$K\times_{B}E\to L\times_{B}E$
is a weak homotopy equivalence. Let $\mathscr{{M}}$ denote the class of morphisms $S\to T$ of simplicial sets such that, for any map $T\to B$, the induced map $S\times_{B}E\to T\times_{B}E$ is a weak homotopy equivalence. We must show that $\mathscr{{M}}$ contains all weak homotopy equivalences. By [Cis19, Proposition 4.6.1], it suffices to show that $\mathscr{{M}}$ contains all morphisms of the form $\Delta^{0}\to\Delta^{n}$, for any $n\geq 0$.
Let $0\leq i\leq n$ be integers and let $\sigma:\Delta^{n}\to B$ be an $n$-simplex, which we depict as $b_{0}\to\cdots\to b_{n}$. We wish to show that the map $E_{b_{i}}\to\Delta^{n}\times_{B}E$ is a weak homotopy equivalence. Choose a cocartesian natural transformation $h:\Delta^{\{i,n\}}\times E_{b_{i}}\to\Delta^{n}\times_{B}E$ rendering the diagram
{${{\{i\}\times E_{b_{i}}}}$${{\Delta^{n}\times_{B}E}}$${{\Delta^{\{i,n\}}\times E_{b_{i}}}}$${{\Delta^{n}}}$
commutative. It will suffice to show that the restriction $h|\{n\}\times E_{b_{i}}:\{n\}\times E_{b_{i}}\to\Delta^{n}\times_{B}E$ is a weak homotopy equivalence. We can factor this map as
$\{n\}\times E_{b_{i}}\xrightarrow{h^{\prime}}E_{b_{n}}\xrightarrow{h^{\prime\prime}}\Delta^{n}\times_{B}E,$
where $h^{\prime\prime}$ denotes the inclusion. The map $h^{\prime}$ is a weak homotopy equivalence by hypothesis. Also, since $p$ is a cocartesian fibration and the inclusion $\{n\}\subset\Delta^{n}$ is final, [Lur09, Proposition 4.1.2.15] shows that $h^{\prime\prime}$ is final. In particular, $h^{\prime\prime}$ is a weak homotopy equivalence. It follows that the map $h|\{n\}\times E_{b_{i}}$ is a composite of weak homotopy equivalences, for it is the composite of two weak homotopy equivalences. The claim follows. ∎
We use Proposition 2.26 to prove the following recognition result for localizations.
Proposition 2.27.
Let $f:\mathcal{C}\to\mathcal{D}$ be a functor of $\infty$-categories. Set $\mathcal{W}=\mathcal{C}\times_{\mathcal{D}}\mathcal{D}^{\simeq}$. Suppose that, for each object $C\in\mathcal{C}$, the map $\mathcal{C}_{/C}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}_{/f\left(C\right)}\right)^{\simeq}$ is a weak homotopy equivalence. The following conditions are equivalent:
  1. (1)
    The functor $f$ exhibits $\mathcal{D}$ as a localization of $\mathcal{C}$ with respect to morphisms in $\mathcal{W}$.
  2. (2)
    The map $\mathcal{W}\to\mathcal{D}^{\simeq}$ is a weak homotopy equivalence.
Proof.
We will use the localization criterion using Rezk’s classification diagram, due to Mazel-Gee. For each simplicial set $K$, let $\operatorname{Fun}^{\mathcal{W}}\left(K,\mathcal{C}\right)\subset\operatorname{Fun}\left(K,\mathcal{C}\right)$ denote the subcategory spanned by the natural transformations whose components are morphisms of $\mathcal{W}$. Recall that the classification diagram $N\left(\mathcal{C},\mathcal{W}\right)$ is the simplicial object in $SS\SS$ whose $n$th simplicial set is given by
$N\left(\mathcal{C},\mathcal{W}\right)_{n}=\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{n},\mathcal{C}\right)=\operatorname{Fun}\left(\Delta^{n},\mathcal{C}\right)\times_{\mathcal{C}^{n+1}}\mathcal{W}^{n+1}.$
We will write $N\left(\mathcal{D}\right)=N\left(\mathcal{D},\mathcal{D}^{\simeq}\right)$; it is a complete Segal space [Rez01], called the classifying diagram of $\mathcal{D}$. According to [MG19, Theorem 3.8] (see also [Ara23, Corollary 4.6]), condition (1) is equivalent to the following condition:
  • (1)
    The map $N\left(\mathcal{C},\mathcal{W}\right)\to N\left(\mathcal{D}\right)$ is a weak equivalence of the complete Segal space model structure.
We will show that condition (1) is equivalent to condition (2).
First we show that the map $d_{0}:N\left(\mathcal{C},\mathcal{W}\right)_{1}\to N\left(\mathcal{C},\mathcal{W}\right)_{0}$ is locally constant. Consider the commutative diagram
[…]
We wish to show that $\pi$ is locally constant. According to Proposition 2.12, the maps $\pi$ and $\pi^{\prime}$ are cocartesian fibrations and $\phi$ carries $\pi$-cocartesian morphisms to $\pi^{\prime}$-cocartesian morphisms. Therefore, for each morphism $\alpha:C\to C^{\prime}$, the square
${{\mathcal{C}^{/C}\times_{\mathcal{C}}\mathcal{W}}}$${{(\mathcal{D}^{/f(C)})^{\simeq}}}$${{\mathcal{C}^{/C^{\prime}}\times_{\mathcal{C}}\mathcal{W}}}$${{(\mathcal{D}^{/f(C^{\prime})})^{\simeq}}}$$\scriptstyle{\alpha_{!}}$$\scriptstyle{f(\alpha)_{!}}$
consisting of fibers of $\pi$ and $\pi^{\prime}$ commutes up to natural equivalence. By hypothesis (and the equivalence of the upper slices and lower slices [Lur09, Proposition 4.2.1.5]) the horizontal arrows are weak homotopy equivalences. Hence $\alpha_{!}$ is a weak homotopy equivalence. It follows from Proposition 2.26 that $\pi$ is locally constant, as desired.
Next, we show that the Reedy fibrant replacement of $N\left(\mathcal{C},\mathcal{W}\right)$ is a Segal space. We must show that, for each $k\geq 2$, the square
${{\operatorname{Fun}^{\mathcal{W}}(\Delta^{k},\mathcal{C})}}$${{\operatorname{Fun}^{\mathcal{W}}(\Delta^{1},\mathcal{C})}}$${{\operatorname{Fun}^{\mathcal{W}}(\Delta^{\{1,\dotsm,k\}},\mathcal{C})}}$${{\operatorname{Fun}^{\mathcal{W}}(\{1\},\mathcal{C})}}$$\scriptstyle{\pi}$
is homotopy cartesian in the Kan–Quillen model structure. Since $\pi$ is locally constant, this is equivalent to the assertion that the map
$\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{k},\mathcal{C}\right)\to\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1}\amalg_{\{1\}}\Delta^{\{1,\dots,k\}},\mathcal{C}\right)$
be a weak homotopy equivalence. But this map is a trivial fibration, for the inclusion $\Delta^{1}\amalg_{\{1\}}\Delta^{\{1,\dots,k\}}\subset\Delta^{k}$ is a weak categorical equivalence [Lur09, Lemma 5.4.5.10].
Now we show that (1$\implies$(2). Suppose that condition (1) holds. We wish to show that the map $N\left(\mathcal{C},\mathcal{W}\right)_{0}\to N\left(\mathcal{D}\right)_{0}$ is a weak homotopy equivalence. Choose a trivial Reedy cofibration $i:N\left(\mathcal{C},\mathcal{W}\right)\to N^{f}\left(\mathcal{C},\mathcal{W}\right)$ with $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ Reedy fibrant (with respect to the Kan–Quillen model structure on $SS\SS$). Since $N\left(\mathcal{D}\right)$ is a complete Segal space, it is Reedy fibrant, so the map $N\left(\mathcal{C},\mathcal{W}\right)\to N\left(\mathcal{D}\right)$ can be factored as a composite
[…]
Since $i$ is a levelwise weak homotopy equivalence, it suffices to show that the map $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{0}\to N\left(\mathcal{D}\right)_{0}$ is a weak homotopy equivalence. By hypothesis, the map $F$ is a weak equivalence of the complete Segal space model structure. Therefore, it suffices to show that the Segal space $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ is a complete Segal space (for weak equivalences of the complete Segal space model structure between complete Segal spaces are nothing but levelwise homotopy equivalences).
Let $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}\subset\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$ denote the union of the components whose vertices are homotopy equivalences of the Segal space $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$. We must show that the map
$\theta:\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{0}\to\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}$
is a homotopy equivalence. Since $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ is a Segal space, the map $F$ is a Dwyer–Kan equivalence of Segal spaces [Rez01, Theorem 7.1]. Therefore, given a morphism $\alpha$ of $\mathcal{C}$, the morphism $i\left(\alpha\right)$ is a homotopy equivalence of $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ if and only if $Fi\left(\alpha\right)$ is a homotopy equivalence of $N\left(\mathcal{D}\right)$. By the definition of $\mathcal{W}$, the latter condition holds if and only if $\alpha$ belongs to $\mathcal{W}$. Therefore, the inverse image of $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}\subset\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$ under the weak homotopy equivalence
$i_{1}:\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1},\mathcal{C}\right)\xrightarrow{\simeq}\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$
is the simplicial subset $\operatorname{Fun}\left(\Delta^{1},\mathcal{W}\right)\subset\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1},\mathcal{C}\right)$. In particular (since $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}$ is a union of components) the map $i_{1}$ restricts to a weak homotopy equivalence
$i_{1}^{\prime}:\operatorname{Fun}\left(\Delta^{1},\mathcal{W}\right)\xrightarrow{\simeq}\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}.$
Now consider the commutative diagram
[…]
The maps $i_{0},i_{1}^{\prime},i_{1}$ are weak homotopy equivalences, and so is the map $\theta^{\prime}$ (for it is a left (and a right) adjoint). Therefore, the map $\theta$ is a weak homotopy equivalence, as required.
Next we show (2$\implies$(1). It will suffice to show that, for each $n\geq 0$, the map $N\left(f\right)_{n}:N\left(\mathcal{C},\mathcal{W}\right)_{n}\to N\left(\mathcal{D}\right)_{n}$ is a weak homotopy equivalence. Since $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ is a Segal space, we only need to prove this in the case where $n\in\{0,1\}$. If $n=0$, the claim follows from our hypothesis (2). If $n=1$, we consider the commutative diagram
[…]
As we already saw in the second paragraph of the proof, the vertical maps are locally constant cocartesian fibrations. (The map $\pi^{\prime}$ is even a Kan fibration). Also, our hypothesis ensures that, for each $C\in\operatorname{Fun}^{\mathcal{W}}\left(\{1\},\mathcal{C}\right)$, the induced map $\pi^{-1}\left(C\right)\to\left(\pi^{\prime}\right)^{-1}\left(f\left(C\right)\right)$ between the fibers is a weak homotopy equivalence. Hence the square is homotopy cartesian in the Kan–Quillen model structure. Since $N\left(f\right)_{0}$ is a weak homotopy equivalence, so must be $N\left(f\right)_{1}$, and the proof is complete. ∎
Corollary 2.28.
Let $\mathcal{C}$ and $\mathcal{D}$ be $\infty$-categories, let $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{D}^{\prime}\subset\mathcal{D}$ be full subcategories, and let $f:\mathcal{C}\to\mathcal{D}$ be a functor which carries $\mathcal{C}^{\prime}$ into $\mathcal{D}^{\prime}$. Set $\mathcal{W}=\mathcal{C}\times_{\mathcal{D}}\mathcal{D}^{\simeq}$. Suppose that, for each $C\in\mathcal{C}$, the functor
$\mathcal{C}^{\prime}_{/C}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}^{\prime}_{/f\left(C\right)}\right)^{\simeq}$
is a weak homotopy equivalence. Then for each $C\in\mathcal{C}$, the functor
$\mathcal{C}^{\prime}_{/C}\to\mathcal{D}^{\prime}_{/f\left(C\right)}$
is a localization with respect to the morphisms whose images in $\mathcal{D}^{\prime}_{f/\left(C\right)}$ are equivalences.
Proof.
We apply Proposition 2.27 to the functor $\mathcal{C}^{\prime}_{/C}\to\mathcal{D}^{\prime}_{f/\left(C\right)}$. It suffices to show that, for each object $\left(\alpha:C^{\prime}\to C\right)\in\mathcal{C}^{\prime}_{/C}$, the map
$\theta:\mathcal{C}^{\prime}_{/\alpha}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}^{\prime}_{/f\left(\alpha\right)}\right)^{\simeq}$
is a weak homotopy equivalence. Consider the commutative diagram
${{\mathcal{C}^{\prime}_{/\alpha}\times_{\mathcal{C}}\mathcal{W}}}$${{(\mathcal{D}^{\prime}_{/f(\alpha)})^{\simeq}}}$${{\mathcal{C}^{\prime}_{/C^{\prime}}\times_{\mathcal{C}}\mathcal{W}}}$${{(\mathcal{D}^{\prime}_{/f(C^{\prime})})^{\simeq}.}}$$\scriptstyle{\theta}$$\scriptstyle{\theta^{\prime}}$
The vertical maps are trivial fibrations, and $\theta^{\prime}$ is a weak homotopy equivalence by hypothesis. Hence $\theta$ is a weak homotopy equivalence, as desired. ∎
To apply Corollary 2.28 to our context, we need a lemma. For each $n,k\geq 0$, write $\mathsf{Disk}_{n}^{=k}$ and $\mathcal{D}\mathsf{isk}_{n}^{=k}$ for the full subcategories of $\mathsf{Disk}_{n}$ and $\mathsf{Disk}_{n}$ spanned by the objects homeomorphic to $\mathbb{R}^{n}\times\{1,\dots,k\}$. Note that, by Kister’s theorem [Kis64], the subcategory $\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\subset\mathsf{Disk}_{n/M}^{=p}$ is spanned by the morphisms which induce bijections between the sets of components.
Lemma 2.29.
Let $n,k\geq 0$ and let $M$ be an $n$-manifold. The map
$\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{=k}\right)^{\simeq}$
is a weak homotopy equivalence.
For the proof of Lemma 2.29, we introduce a bit of notation.
Notation 2.30.
Let $k\geq 0$ and let $M$ be a manifold.
  • We will write $B^{\prime}_{k}\left(M\right)\subset B_{k}\left(M\right)$ for the open set consisting of those subsets $S\subset M$ of cardinality $k$ such that the map $\pi_{0}\left(S\right)\to\pi_{0}\left(M\right)$ is injective.
  • Let $k\geq 0$ and let $M$ be a manifold. We let $I_{M}^{=k}$ denote the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between the elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components.
Proof of Lemma 2.29.
The assignment $U\mapsto\left(U\hookrightarrow M\right)$ induces a categorical equivalence $N\left(I_{M}^{=k}\right)\xrightarrow{\simeq}\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}$, so it suffices to show that the composite
$\theta:N\left(I_{M}^{=k}\right)\to\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{=k}\right)^{\simeq}$
is a weak homotopy equivalence.
Given $n$-manifolds $X,Y$, let $\operatorname{Emb}^{\simeq}\left(X,Y\right)\subset\operatorname{Emb}\left(X,Y\right)$ denote the subspace consisting of the isotopy equivalences from $X$ to $Y$. According to Proposition 2.22 and [Lur09, Proposition 4.2.4.1], the map $\theta$ is a weak homotopy equivalence if and only if the following condition holds:
  • ($\ast$)
    For each object $U\in\mathcal{D}\mathsf{isk}_{n}^{=k}$, the map
    $\operatorname{hocolim}_{V\in I_{M}^{=k}}\operatorname{Sing}\operatorname{Emb}^{\simeq}\left(U,V\right)\to\operatorname{Sing}\operatorname{Emb}\left(U,M\right)$
    is a weak homotopy equivalence.
We will prove ($\ast$). The claim is obvious if $k=0$, so assume that $k\geq 1$. Fix a homeomorphism $U\cong\mathbb{R}^{n}\times\{1,\dots,k\}$ and let $V_{1},\dots,V_{k}\subset V$ denote the components of $V$. Evaluation at the origin gives us a commutative diagram
${{\operatorname{Emb}^{\simeq}(U,V)}}$${{\operatorname{Emb}(U,V)}}$${{\operatorname{Emb}(U,M)}}$${{\coprod_{\sigma\in\Sigma_{k}}\prod_{i=1}^{k}V_{\sigma(i)}}}$${{\operatorname{Conf}(k,V)}}$${{\operatorname{Conf}(k,M)}}$${{B^{\prime}_{k}(V)}}$${{B_{k}(V)}}$${{B_{k}(M)}}$$\scriptstyle{(1)}$$\scriptstyle{(3)}$$\scriptstyle{(4)}$$\scriptstyle{(2)}$
of topological spaces. Proposition 2.9 shows that the square (1) is homotopy cartesian. The squares (2), (3), (4) are homotopy cartesian because they are strictly cartesian and their vertical arrows are Serre fibrations. It follows that the outer square is homotopy cartesian. Since colimits in $\mathcal{S}$ are universal [Lur09, Lemma 6.1.3.14], we are reduced to showing that the map
$\operatorname{hocolim}_{V\in I_{M}^{=p}}\operatorname{Sing}B^{\prime}_{k}\left(V\right)\to\operatorname{Sing}B_{p}\left(M\right)$
is a weak homotopy equivalence. This follows from [Lur17, Theoerm A.3.1]. ∎
We now arrive at the proof of Theorem 2.24.
Proof of Theorem 2.24.
We will apply Corollary 2.28 to the functor $\mathsf{Mfld}_{n}\to\mathcal{M}\mathsf{fld}_{n}$ and the full subcategories $\mathsf{Disk}_{n}\subset\mathsf{Mfld}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}\subset\mathcal{M}\mathsf{fld}_{n}$ for part (1), and the full subcategories $\mathsf{Disk}_{n}^{\leq k}\subset\mathsf{Mfld}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}\subset\mathcal{M}\mathsf{fld}_{n}$ for part (2). For part (1), we must show that for each $n$-manifold $M$, the functor
$\mathsf{Disk}_{n/M}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}\right)^{\simeq}$
is a weak homotopy equivalence, and for part (2), we must show that the functor
$\mathsf{Disk}_{n/M}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)^{\simeq}$
is a weak homotopy equivalence. Both of these assertions follow from Lemma 2.29. ∎
Remark 2.31.
The proof of Theorem 2.24 carries over to the smooth case (using Remark 2.10). In other words, for each $n,k\geq 0$ and each smooth $n$-manifold $M$, the functors $\mathsf{Disk}_{\mathrm{sm},n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n/M}^{\leq k}$ and $\mathsf{Disk}_{\mathrm{sm},n/M}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n/M}$ are localization with respect to smooth isotopy equivalences.
Remark 2.32.
In [KSW24, Lemma A.1, A.2], Karlsson, Scheimbauer, and Walde independently proved Proposition 2.26 and the implication (2$\implies$(1) of Proposition 2.27.
2.4. Extension by Zero Objects
Consider the following problem: Let $\mathcal{C}$ be an $\infty$-category and let $\mathcal{C}_{0},\mathcal{C}_{1}\subset\mathcal{C}$ be full subcategories such that every object of $\mathcal{C}$ belongs to exactly one of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$. Let $\mathcal{D}$ be a pointed $\infty$-category and suppose we are given a functor $f:\mathcal{C}\to\mathcal{D}$ carrying each object of $\mathcal{C}_{0}$ to a zero object. One would expect that $f$ is equivalent to a smaller piece of data. In other words, we would expect that there is a categorical equivalence
$\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\simeq\{\text{something smaller}\}$
where $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\subset\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)$ denotes the full subcategory spanned by the objects carrying each object in $\mathcal{C}_{0}$ to a zero object. We want to figure out what the smaller thing is.
A naïve guess of the “smaller thing” is $\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$, but this is too naïve. Indeed, there could be a morphism in $\mathcal{C}_{1}$, say $\alpha$, which factors through a morphism of $\mathcal{C}_{0}$, and there is no reason why an arbitrary functor $\mathcal{C}_{1}\to\mathcal{D}$ carries $\alpha$ to a zero object. So the restriction $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$ is, in general, not essentially surjective.
The next guess is that the essential image of the restriction functor $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$ will suffice. If $\mathcal{C}$ and $\mathcal{D}$ are nerves of ordinary categories, this is true. But for arbitrary $\infty$-categories, this is false; the problem is that there could be a nontrivial homotopy between zero maps, as one can see from the particular case where $\mathcal{C}=\Delta^{2}$ and $\mathcal{C}_{1}=\{1\}$.
In this subsection, we will give one answer to this question in the case where $\mathcal{C}_{1}$ is “isolated” in $\mathcal{C}$.
To state our main result, we need a bit of terminology.
Definition 2.33.
Let $\mathcal{C}$ be an $\infty$-category and let $\mathcal{C}^{\prime}\subset\mathcal{C}$ be a full subcategory. A morphism $X\to Y$ of $\mathcal{C}^{\prime}$ is said to be isolated in $\mathcal{C}$ if it does not factor through any object of $\mathcal{C}$ which lies outside of $\mathcal{C}^{\prime}$. We say that $\mathcal{C}^{\prime}$ is isolated in $\mathcal{C}$ if the morphisms of $\mathcal{C}^{\prime}$ that are isolated in $\mathcal{C}$ are closed under compositions.
Here is the main result of this subsection.
Proposition 2.34.
Let $\mathcal{C}$ be an $\infty$-category and let $\mathcal{C}_{0},\mathcal{C}_{1}\subset\mathcal{C}$ be full subcategories such that every object of $\mathcal{C}$ belongs to exactly one of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$. Let $\mathcal{D}$ be a pointed $\infty$-category and let $\mathcal{Z}\subset\mathcal{D}$ denote the full subcategory spanned by the zero objects. Suppose that $\mathcal{C}_{1}$ is isolated in $\mathcal{C}$, and let $\mathcal{C}_{1}^{{\rm isltd}}\subset\mathcal{C}_{1}$ denote the subcategory spanned by the morphisms isolated in $\mathcal{C}$. The functor
$\theta:\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{Z}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1}^{{\rm isltd}},\mathcal{D}\right)$
is a trivial fibration.
Proof.
Let us say that a morphism $f$ of $\mathcal{C}$ is basic if either $f$ is a morphism of $\mathcal{C}_{1}^{{\rm isltd}}$ or $f$ is not a morphism of $\mathcal{C}^{1}$. We will say that a simplex $\sigma:\Delta^{d}\to\mathcal{C}$ is basic if for each $0\leq i<d$, the restriction $\sigma|\Delta^{\{i,i+1\}}$ is basic. We let $\{\sigma_{\alpha}:\Delta^{d_{\alpha}}\to\mathcal{C}\}_{\alpha\in A}$ denote the set of all basic simplices of $\mathcal{C}$. Note that the map $\coprod_{\alpha\in A}\Delta^{d_{\alpha}}\to\mathcal{C}$ is an epimorphism of simplicial sets because every simplex of $\mathcal{C}$ is a face of a basic simplex.
For each $\alpha\in A$, set $P_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{0}$, $Q_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{1}$, and $Q^{\prime}_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{1}^{{\rm isltd}}$. Since $\theta$ is a pullback of the map
$\operatorname{Fun}\left(\coprod_{\alpha\in A}\Delta^{d_{\alpha}},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\coprod_{\alpha\in A}P_{\alpha},\mathcal{D}\right)}\operatorname{Fun}\left(\coprod_{\alpha\in A}P_{\alpha},\mathcal{Z}\right)\to\operatorname{Fun}\left(\coprod_{\alpha\in A}Q_{\alpha}^{\prime},\mathcal{D}\right),$
it suffices to show that for each $\alpha\in A$, the map
$\theta_{\alpha}:\operatorname{Fun}\left(\Delta^{d_{\alpha}},\mathcal{D}\right)\times_{\operatorname{Fun}\left(P_{\alpha},\mathcal{D}\right)}\operatorname{Fun}\left(P_{\alpha},\mathcal{Z}\right)\to\operatorname{Fun}\left(Q^{\prime}_{\alpha},\mathcal{D}\right)$
is a trivial fibration. Now since $\sigma_{\alpha}$ is isolated, the subcategory $Q^{\prime}_{\alpha}\subset Q_{\alpha}$ is spanned by the morphisms of $Q_{\alpha}$ that are isolated in $\Delta^{d_{\alpha}}$. Therefore, we are reduced to the case where $\mathcal{C}=\Delta^{d}$ for some $d\geq 0$. In this case, we will prove the assertion by induction on $d$.
If $d=0$, the claim is obvious. For the inductive step, suppose we have proved the assertion up to $d-1$. If $\mathcal{C}_{0}$ is empty, the claim is trivial, so assume that $\mathcal{C}_{0}$ is nonempty. Let $m$ be the maximal integer which belongs to $\mathcal{C}_{0}$. There are two cases to consider.
Suppose first that $m=d$. In this case, [Lur09, Proposition 4.3.2.15] shows that the functor
$\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{Z}\right)\to\operatorname{Fun}\left(\Delta^{d-1},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{Z}\right)$
is a trivial fibration. By the induction hypothesis, the functor
$\operatorname{Fun}\left(\mathcal{C}\cap\Delta^{d-1},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{Z}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1}^{{\rm isltd}},\mathcal{D}\right)$
is a trivial fibration. It follows that $\theta$ is a composition of trivial fibrations and hence is itself a trivial fibration.
Suppose next that $m<d$. Set $X=\Delta^{\{0,\dots,m\}}\amalg_{\{m\}}\Delta^{\{m\dots,d\}}$. The map $\theta$ factors as
[…]
The map $f$ is a trivial fibration because the inclusion $X\to\mathcal{C}$ is a weak categorical equivalence [Lur09, Lemma 5.4.5.10]. It will therefore suffice to show that $g$ is a trivial fibration. Since $g$ is a categorical fibration, it suffices to show that it is a categorical equivalence. We may identify $g$ with the map between the pullbacks of the rows of the following commutative diagram:
${{\operatorname{Fun}(\Delta^{m},\mathcal{D})\times_{\operatorname{Fun}(\mathcal{C}_{0}\cap\Delta^{m},\mathcal{D})}\operatorname{Fun}(\mathcal{C}_{0}\cap\Delta^{m},\mathcal{Z})}}$${{\operatorname{Fun}(\{m\},\mathcal{Z})}}$${{\operatorname{Fun}(\Delta^{\{m,\dots,d\}},\mathcal{D})\times_{\operatorname{Fun}(\{m\},\mathcal{D})}\operatorname{Fun}(\{m\},\mathcal{Z})}}$${{\operatorname{Fun}(\Delta^{m}\cap\mathcal{C}_{1}^{\mathrm{isltd}},\mathcal{D})}}$${{\Delta^{0}}}$${{\operatorname{Fun}(\Delta^{\{m,\dots,d\}}\cap\mathcal{C}_{1}^{\mathrm{isltd}},\mathcal{D}).}}$
By the induction hypothesis, the vertical arrows of the above diagrams are trivial fibrations. Moreover, the horizontal arrows are categorical fibrations of $\infty$-categories. Hence $g$ is also a categorical equivalence, and the proof is complete. ∎
3. Polynomial Functors
In this section, we will prove the existence of polynomial approximations (Theorem 3.4). In Subsection 3.1, we reduce this to a lemma on Kan extensions of isotopy functors (Lemma 3.5) and a description of polynomial functors in terms of Kan extensions and sheaves (Theorem 3.10). Subsections 3.2 and 3.3 are devoted to the proofs of these results. In Subsection 3.4, we will explain how polynomial approximations can be organized into a tower, and discuss its convergence.
Warning 3.1.
To avoid writing “$\mathrm{op}$” repeatedly, we will work with covariant functors, so that polynomial functors are described in terms of cosheaves and left Kan extensions. The theory of contravariant polynomial functors can be recovered by replacing every functor by its opposite.
Remark 3.2.
Using Remark 2.31, we can show that everything in this section is valid in the smooth case if we make the following replacements:
  • manifolds by smooth manifolds;
  • isotopy equivalences by smooth isotopy equivalences;
  • $\operatorname{Disj}\left(-\right)$ and $\operatorname{Disj}^{\leq k}\left(-\right)$ by $\operatorname{Disj}_{\mathrm{sm}}\left(-\right)$ and $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(-\right)$;
  • $\mathcal{M}\mathsf{fld}_{n}$, $\mathcal{D}\mathsf{isk}_{n}$, and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ by $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$, and $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$.
3.1. Statement of the Main Result
In this subsection, we state the main result of this section and explain how we are going to prove it.
Definition 3.3.
Let $k\geq 0$, let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with finite colimits. Let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor.
  1. (1)
    Let $\mathcal{J}\subset N\left(\operatorname{Open}\left(M\right)\right)$ be a subcategory and let $G:\mathcal{J}\to\mathcal{C}$ be a functor. We say that $G$ is isotopy invariant if it carries each morphism in $\mathcal{J}$ which is an isotopy equivalence to an equivalence of $\mathcal{C}$. Isotopy invariant functors will be called isotopy functors. We will write $\operatorname{Fun}_{\mathrm{istp}}\left(\mathcal{J},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{J},\mathcal{C}\right)$ for the full subcategory spanned by the isotopy functors.
  2. (2)
    We say that $F$ is exhaustive if for each increasing sequence $U_{0}\subset U_{1}\subset\cdots$ of open sets of $M$, the map
    $\operatorname{colim}_{i}F\left(U_{i}\right)\to F\left(\bigcup_{i\geq 0}U_{i}\right)$
    is an equivalence.
  3. (3)
    We say that $F$ is $k$-excisive, or polynomial of degree $\leq k$, if for each open set $U$ of $M$ and each pairwise disjoint closed subsets $A_{0},\dots,A_{k}$ of $U$, the map
    $\operatorname{colim}_{\emptyset\neq S\subset\{0,\dots,k\}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)\to F\left(U\right)$
    is an equivalence.
  4. (4)
    We let $\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\subset\operatorname{Fun}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ denote the full subcategory spanned by the $k$-excisive, exhaustive isotopy functors.
Here is the main result of this section.
Theorem 3.4.
Let $k\geq 0$, let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. The left Kan extension functor $\operatorname{Fun}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\to\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ restricts to a categorical equivalence
$\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{\mathrm{istp},\,\mathrm{exh}}^{k}\left(M;\mathcal{C}\right).$
In particular, the inclusion
$\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\hookrightarrow\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$
admits a right adjoint, which carries each object $F\in\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ to a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$.
We will deduce Theorem 3.4 from two results: The first one is the following lemma, which we prove in Subsection 3.2.
Lemma 3.5.
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor which is a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. If $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$ is an isotopy functor, so is $F$.
The second one concerns the identification of polynomial functors with functors enjoying certain gluing properties. The gluing properties are best expressed using cosheaves with respect to Grothendieck topologies on $N\left(\operatorname{Open}\left(M\right)\right)$, which we now introduce.
Definition 3.6.
Let $\mathcal{C}$ be an $\infty$-category and let $\mathcal{A}$ be another $\infty$-category equipped with a Grothendieck topology [Lur09, Definition 6.2.2.1]. We will say that a functor $F:\mathcal{A}\to\mathcal{C}$ is a cosheaf (with respect to the Grothendieck topology) if for each object $A\in\mathcal{A}$ and each covering sieve $\mathcal{A}_{/A}^{0}\subset\mathcal{A}_{/A}$ of $A$, the composite
[…]
is a colimit diagram.
Definition 3.7.
Let $X$ be a topological space, and let $k\geq 0$. A subset $\mathcal{U}\subset\operatorname{Open}\left(X\right)$ of $X$ is called a Weiss $k$-cover if for each subset $S\subset X$ of cardinality $\leq k$, there is an element of $\mathcal{U}$ which contains $S$. The Weiss $k$-topology on $X$ is the Grothendieck topology on $N\left(\operatorname{Open}\left(X\right)\right)$ whose covering sieves of each open set $U\in\operatorname{Open}\left(X\right)$ are precisely the Weiss $k$-covers $\mathcal{U}$ of $U$ that are also sieves on $U$. If $\mathcal{C}$ is an $\infty$-category, a functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is called a Weiss $k$-cosheaf if it is a cosheaf with respect to the Weiss $k$-topology. The Weiss topology on $X$ is the intersection of the Weiss $k$-topologies as $k$ ranges over all nonnegative integers. A Weiss cosheaf is a cosheaf with respect to the Weiss topology.
Example 3.8.
Let $X$ be a topological space and $\mathcal{C}$ an $\infty$-category.
  1. (1)
    Any nonempty collection of open sets of $X$ is a Weiss 00-cover of $X$. Consequently, a functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is a Weiss 00-cosheaf if and only if it is essentially constant, i.e., it factors through a contractible Kan complex. Since $N\left(\operatorname{Open}\left(X\right)\right)$ is already weakly contractible, this is equivalent to the requirement that $F$ carries all morphisms to equivalences.
  2. (2)
    A functor $F:N\left(\operatorname{Open}\left(X\right)\right)\to\mathcal{C}$ is a Weiss $1$-cosheaf if and only if for each nonempty sieve $\mathcal{U}\subset\operatorname{Open}\left(X\right)$, the map
    $\operatorname{colim}_{U\in\mathcal{U}}F\left(U\right)\to F\left(\bigcup_{U\in\mathcal{U}}U\right)$
    is an equivalence. Thus a Weiss $1$-cosheaf is a cosheaf on $X$ (that is, a cosheaf with respect to the standard Grothendieck topology on $\operatorname{Open}\left(X\right)$) if and only if $F\left(\emptyset\right)$ is an initial object.
Definition 3.9.
Let $k\geq 0$, let $X$ be a topological space, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(X\right)\right)$ be a functor of small $\infty$-categories. For each finite set $S\subset X$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $S\subset\chi\left(I\right)$.
  1. (1)
    We say that $\chi$ satisfies the Weiss $k$-condition if for each subset $S\subset M$ of cardinality at most $k$, the $\infty$-category $\mathcal{I}_{S}$ is weakly contractible.
  2. (2)
    We say that $\chi$ satisfies the Weiss condition if for every finite subset $S\subset M$, the $\infty$-category $\mathcal{I}_{S}$ is weakly contractible.
We now arrive at the identification result of polynomial functors with cosheaves and Kan extensions, which we prove in Subsection 3.3.
Theorem 3.10.
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$.
  2. (2)
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor satisfying the Weiss $k$-condition. Then the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(U\right)$
    is an equivalence of $\mathcal{C}$.
  3. (3)
    The functor $F$ is a Weiss $k$-cosheaf.
  4. (4)
    The functor $F$ is $k$-excisive and exhaustive.
For later discussions, we will also prove the following limit case of Theorem 3.10:
Theorem 3.11.
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$.
  2. (2)
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor which satisfies the Weiss condition. Then the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(U\right)$
    is an equivalence of $\mathcal{C}$.
  3. (3)
    The functor $F$ is a Weiss cosheaf.
Assuming Lemma 3.5 and Theorem 3.10, we can prove Theorem 3.4 as follows:
Proof of Theorem 3.4.
By Lemma 3.5, the left Kan extension functor restricts to a left adjoint
$L:\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right),$
which is fully faithful by [Lur09, Proposition 4.3.2.15]. By Theorem 3.10, the essential image of $L$ is $\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$. The claim follows. ∎
3.2. Proof of Lemma 3.5
In this subsection, we prove Lemma 3.5.
Notation 3.12.
Let $n,k\geq 0$, let $M$ be an $n$-manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. We consider the following $\infty$-categories.
  1. (1)
    The full subcategory $\operatorname{Fun}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$.
  2. (2)
    The full subcategory $\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $N\left(\operatorname{Disj}\left(M\right)\right)$.
  3. (3)
    We let $\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ denote the intersection of $\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ and $\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$, and define $\operatorname{Fun}_{\mathrm{istp}}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ similarly.
  4. (4)
    The full subcategory $\operatorname{Fun}^{\leq k}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$.
  5. (5)
    The full subcategory $\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)$ spanned by the functors that are left Kan extensions of their restrictions to $\mathcal{D}\mathsf{isk}_{n/M}$.
Lemma 3.13.
Let $n,k\geq 0$, let $M$ be an $n$-manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. Then:
  1. (1)
    The functor $\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is well-defined and is a categorical equivalence.
  2. (2)
    The functor $\operatorname{Fun}^{\leq k}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is well-defined and is a categorical equivalence.
  3. (3)
    A functor $F\in\operatorname{Fun}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is isotopy invariant if and only if $F|N\left(\operatorname{Disj}\left(M\right)\right)$ is isotopy invariant.
  4. (4)
    A functor $F\in\operatorname{Fun}^{\leq k}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ is isotopy invariant if and only if $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$ is isotopy invariant.
Proof.
We will prove assertions (1) and (3); assertions (2) and (4) can be proved similarly.
First we prove the “well-defined” part of (1). Let $F:\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ be a functor which is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n/M}$. We wish to show that the functor $F|N\left(\operatorname{Open}\left(M\right)\right)$ is a left Kan extension of $N\left(\operatorname{Disj}\left(M\right)\right)$ and that it is isotopy invariant. The latter assertion is obvious. For the former, let $U\subset M$ be an open set. We must show that the composite
[…]
is a colimit diagram. Let $\iota:U\to M$ denote the inclusion. By hypothesis, the composite
[…]
is a colimit diagram. It will therefore suffice to show that the functor
$\theta:N\left(\operatorname{Disj}\left(M\right)\right)_{/U}\to\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota}$
is final. Since the functor $\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota}\to\mathcal{D}\mathsf{isk}_{n/U}$ is a trivial fibration, it suffices to show that the functor $N\left(\operatorname{Disj}\left(M\right)\right)_{/U}\cong N\left(\operatorname{Disj}\left(U\right)\right)\to\mathcal{D}\mathsf{isk}_{n/U}$ is final. This follows from Theorem 2.24, since localization functors are final [Lur24, Tag 02N9].
We now prove (3). We must show that for any isotopy invariant functor $F:N\left(\operatorname{Disj}\left(M\right)\right)\to\mathcal{C}$, its left Kan extension is also isotopy invariant. By Theorem 2.24, we may assume that $F$ extends to a functor $G:\mathcal{D}\mathsf{isk}_{n/M}\to\mathcal{C}$. Let $G^{\prime}:\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ denote a left Kan extension of $G$. By the result in the previous paragraph, $G^{\prime}|N\left(\operatorname{Open}\left(M\right)\right)$ is a left Kan extension of $F=G^{\prime}|N\left(\operatorname{Disj}\left(M\right)\right)$. Since $G^{\prime}|N\left(\operatorname{Open}\left(M\right)\right)$ is isotopy invariant, we are done.
Finally, we complete the proof of (1). We must show that the functor
$f:\operatorname{Fun}^{\infty}\left(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}^{\infty}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$
is a categorical equivalence. Consider the commutative diagram
${{\operatorname{Fun}^{\infty}(\mathcal{M}\mathsf{fld}_{n/M},\mathcal{C})}}$${{\operatorname{Fun}^{\infty}_{\mathrm{istp}}(N(\operatorname{Open}(M)),\mathcal{C})}}$${{\operatorname{Fun}(\mathcal{D}\mathsf{isk}_{n/M},\mathcal{C})}}$${{\operatorname{Fun}_{\mathrm{istp}}(N(\operatorname{Disj}(M)),\mathcal{C}).}}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{g^{\prime}}$$\scriptstyle{f^{\prime}}$
The functors $g$ and $g^{\prime}$ are trivial fibrations by [Lur09, Proposition 4.3.2.15], and the functor $f^{\prime}$ is a categorical equivalence by Theorem 2.24. Hence $f$ is also a categorical equivalence. The proof is now complete. ∎
Proof of Lemma 3.5.
This follows from Lemma 3.13. ∎
3.3. Proofs of Theorems 3.10 and 3.11
We now turn to the proofs of Theorems 3.10 and 3.11. We need a few lemmas.
The first lemma gives a sufficient condition for a functor into $\mathcal{D}\mathsf{isk}_{n/M}$ (or $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ to be final.
Lemma 3.14.
Let $n,k\geq 0$, let $\mathcal{I}$ be a (small) category and let $\overline{f}:\mathcal{I}^{\triangleright}\to\mathcal{M}\mathsf{fld}_{n}^{\Delta}$ be a functor. Set $\overline{f}\left(I\right)=U_{I}$ for $I\in\mathcal{I}$ and $\overline{f}\left(\infty\right)=M$. For each subset $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $U_{I}$ contains $S$.
  1. (1)
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$, then the functor $N\left(\mathcal{I}\right)\to\mathcal{D}\mathsf{isk}_{n}{}_{/M}$ is final.
  2. (2)
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$ of cardinality $\leq k$, then the functor $N\left(\mathcal{I}\right)\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ is final.
Proof.
We prove part (1); the proof of part (2) is similar. By Proposition 2.17 and [Lur09, Theorem 4.2.4.1], it will suffice to show that, for each object $V\in\mathcal{D}\mathsf{isk}_{n}$, the map
$\operatorname{hocolim}_{I\in\mathcal{I}}\operatorname{Sing}\operatorname{Emb}\left(V,U_{I}\right)\to\operatorname{Sing}\operatorname{Emb}\left(V,M\right)$
is a weak homotopy equivalence. If $V=\emptyset$, the claim is obvious because $\mathcal{I}=\mathcal{I}_{\emptyset}$ is weakly contractible by hypothesis. So suppose that $V$ is nonempty. Let $p$ denote the cardinality of the set of components of $V$ and fix a homeomorphism $V\cong\mathbb{R}^{n}\times\{1,\dots,p\}$. By Proposition 2.9, the evaluation at the origin of $\mathbb{R}^{n}$ determines a homotopy cartesian square
${{\operatorname{Sing}\operatorname{Emb}(V,U_{I})}}$${{\operatorname{Sing}\operatorname{Emb}(V,M)}}$${{\operatorname{Sing}\operatorname{Conf}(p,U_{I})}}$${{\operatorname{Sing}\operatorname{Conf}(p,M)}}$
for every $I\in\mathcal{I}$. Since colimits in $\mathcal{S}$ are universal, we are reduced to showing that the map
$\operatorname{hocolim}_{I\in\mathcal{I}}\operatorname{Sing}\operatorname{Conf}\left(p,U_{I}\right)\to\operatorname{Sing}\operatorname{Conf}\left(p,M\right)$
is a weak homotopy equivalence. This is a direct consequence of our hypothesis and [Lur17, Theorem A.3.1]. ∎
[…]
Lemma 3.15.
Let $n\geq 1$, $k\geq 0$, and let $-1<c_{0}<\cdots<c_{k}<1$ and $r>0$ be real numbers. Set
$C_{i}=\{x\in\mathbb{R}^{n}\mid\left(x_{1}-c_{i}\right)^{2}+\sum_{1<i\leq n}x_{i}^{2}\leq r^{2}\},$
and suppose that the sets $C_{0},\dots,C_{k}$ are mutually disjoint and are contained in the interior of $D^{n}$. Then $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)$ is obtained from $\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching an $\left(n-1\right)$-handle $k$ times. (See Figure 3.1.)
Figure 3.1. Picture of Lemma 3.15.
Proof.
Using bump functions, construct a smooth function $\phi:\mathbb{R}\to\mathbb{R}$ with the following properties (Figure 3.2):
  1. (1)
    For each $0\leq i\leq k$, the restriction $\phi|[c_{i}-r,c_{i}+r]$ is given by $\phi\left(x\right)=-\left(x-c_{i}\right)^{2}$.
  2. (2)
    The derivative of $\phi$ is positive on $\left(-\infty,c_{1}\right)$ and is negative on $\left(c_{k},\infty\right)$.
  3. (3)
    The function $\phi$ agrees with $-x^{2}$ outside $[-1,1]$
  4. (4)
    For each $1\leq i\leq k$, there is a unique point $a_{i}\in\left(c_{i-1},c_{i}\right)$ such that $\phi^{\prime}\left(a_{i}\right)=0$.
  5. (5)
    For each $0<i\leq k$, we have $\phi^{\prime\prime}\left(a_{i}\right)>0$.
  6. (6)
    We have $-1<\phi\left(a_{1}\right)<\cdots<\phi\left(a_{k}\right)$.
Figure 3.2. Graph of $\phi$.
We then define $F:\mathbb{R}^{n}\to\mathbb{R}$ by $F\left(x\right)=\phi\left(x_{1}\right)-\sum_{1<i\leq n}x_{i}^{2}$. By construction, the critical points of $F$ in $F^{-1}\left((-\infty,r^{2}]\right)$ are the points $\{\left(a_{i},0,\dots,0\right)\}_{1\leq i\leq k}$, and all of them are nondegenerate and have index $n-1$. Therefore, the Morse theory (see, e.g., [Kos93, Proposition VII.2.2]) shows that the set $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)=F^{-1}\left((-\infty,-r^{2}]\right)$ is obtained from $F^{-1}\left((-\infty,-1]\right)=\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching $\left(n-1\right)$-handles $k$ times. ∎
Lemma 3.16.
Let $n,k\geq 0$ and let $M$ be an $n$-manifold. Let $\mathcal{U}$ be a set of open sets of $M$ which satisfies the following conditions:
  1. (1)
    The set $\mathcal{U}$ contains $\operatorname{Disj}^{\leq k}\left(M\right)$.
  2. (2)
    Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of elements in $\mathcal{U}$. Then $\bigcup_{i\geq 0}U_{i}$ belongs to $\mathcal{U}$.
  3. (3)
    Let $U\subset M$ be an open set and let $A_{0},\dots,A_{k}$ be pairwise disjoint closed sets of $U$. Suppose that, for each nonempty subset $S\subset\{0,\dots,k\}$, the open set $U\setminus\bigcup_{i\in S}A_{i}$ belongs to $\mathcal{U}$. Then $U$ belongs to $\mathcal{U}$.
Then $\mathcal{U}=\operatorname{Open}\left(M\right)$.
Proof.
The proof proceeds in several steps.
  1. (Step 1)
    We show that $\mathcal{U}$ contains $\operatorname{Disj}\left(M\right)$. It suffices to show that $\operatorname{Disj}^{\leq p}\left(M\right)\subset\mathcal{U}$ for every $p\geq 0$. We prove this by induction on $p$. If $p\leq k$, the claim follows from our assumption (1). Suppose we have proved that $\operatorname{Disj}^{\leq p-1}\left(M\right)\subset\mathcal{U}$ for some $p>k$, and let $U\in\operatorname{Disj}^{\leq p}\left(M\right)$. We wish to show that $U\in\mathcal{U}$. Let $A_{0},\dots,A_{k}$ be distinct components of $U$. (There may be other components, but choose $k+1$ one of them.) For every nonempty subset $S\subset\{0,\dots,k\}$, the open set $U\setminus\bigcup_{i\in S}A_{i}$ belongs to $\operatorname{Disj}^{\leq p-1}\left(M\right)$ and hence to $\mathcal{U}$. It follows from the induction hypothesis and condition (3) that $U\in\mathcal{U}$, completing the induction.
  2. (Step 2)
    Let $U\subset M$ be an open set which has the form $U=\operatorname{Int}N$ for some compact manifold $N$ with boundary admitting a handle decomposition. We will show that $U$ belongs to $\mathcal{U}$.
    Let $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ be an $\left(n+1\right)$-tuple of nonnegative integers. We say that a handle decomposition $\emptyset=N_{0}\subset\cdots\subset N_{\sum_{i=0}^{n}a_{i}}=N$ of $N$ is of type $\left(a_{0},\dots,a_{n}\right)$ if for each $0\leq d\leq n$, there are exactly $a_{d}$ integers $i$ such that $N_{i-1}$ is obtained from $N_{i}$ by attaching a $d$-handle. We define the handle type of $N$ to be the minimal element $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ for which $P$ admits a handle decomposition of type $\left(a_{0},\dots,a_{n}\right)$, where $\mathbb{Z}_{\geq 0}^{n+1}$ is endowed with the lexicographic ordering read from right to left. (In other words, given distinct elements $\left(a_{0},\dots,a_{n}\right),\left(b_{0},\dots,b_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$, we declare that $\left(b_{0},\dots,b_{n}\right)<\left(a_{0},\dots,a_{n}\right)$ if and only if $b_{i}<a_{i}$, where $i$ is the maximal integer such that $a_{i}\neq b_{i}$.) We will show by transfinite induction on the handle type of $N$ that $U$ belongs to $\mathcal{U}$.
    Let $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ be the handle type of $N$. Suppose that the claim has been proved for every element $\left(a^{\prime}_{0},\dots,a^{\prime}_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ smaller than $\left(a_{0},\dots,a_{n}\right)$. We must show that $U$ belongs to $\mathcal{U}$. If $a_{1}=\cdots=a_{n}=0$, then $U$ belongs to $\operatorname{Disj}\left(M\right)$, and the claim follows from Step 1. So suppose that $a_{i}>0$ for some $i$. Let
    $\emptyset=N_{0}\subset\cdots\subset N_{a}=N$
    be a handle decomposition of $N$ of type $\left(a_{0},\dots,a_{n}\right)$, where we wrote $a=\sum_{i=0}^{n}a_{i}$. Since attaching a 00-handle is equivalent to adding a disjoint copy of $D^{n}$, by rearranging the order of the handle attachment, we may assume that $N_{a}$ is obtained from $N_{a-1}$ by attaching a handle of positive index, say $\lambda$. Let $e:D^{\lambda}\times D^{n-\lambda}\to N_{a}$ denote the corresponding embedding. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{\lambda}$ and set $A_{j}=C_{j}\times D^{n-\lambda}$. If $S\subset\{0,\dots,k\}$ is a nonempty subset, then we have
    $U\setminus\bigcup_{j\in S}A_{j}$ $=\operatorname{Int}\left(N_{a}\setminus\bigcup_{j\in S}e\left(\operatorname{Int}\left(C_{j}\right)\times D^{n-\lambda}\right)\right)$
    $=\operatorname{Int}\left(N_{a-1}\amalg_{S^{\lambda-1}\times D^{n-\lambda}}e\left(\left(D^{\lambda}\setminus\bigcup_{j\in S}\operatorname{Int}\left(C_{j}\right)\right)\times D^{n-\lambda}\right)\right).$
    Now according to Lemma 3.15, the manifold with boundary $N_{a}\left(S\right)=N_{a-1}\amalg_{S^{\lambda-1}\times D^{n-\lambda}}e\left(\left(D^{\lambda}\setminus\bigcup_{j\in S}\operatorname{Int}\left(C_{j}\right)\right)\times D^{n-\lambda}\right)$ admits a handle decomposition of type $\left(b_{0},\dots,b_{n}\right)$, where
    $b_{i}=\begin{cases}a_{i}&\text{if }i\neq\lambda,\lambda-1,\\ a_{\lambda-1}+\left|S\right|-1&\text{if }i=\lambda-1,\\ a_{\lambda}-1&\text{if }i=\lambda.\end{cases}$
    Since $\left(b_{0},\dots,b_{n}\right)<\left(a_{0},\dots,a_{n}\right)$, the induction hypothesis implies that $\operatorname{Int}\left(N_{a}\left(S\right)\right)$ belongs to $\mathcal{U}$. It follows from condition (3) that $U$ belongs to $\mathcal{U}$, as desired.
  3. (Step 3)
    We show that if $U$ is smoothable or $n\neq 4$, then $U$ belongs to $\mathcal{U}$. By Step 2 and assumption (2), it suffices to show that $U$ has a (possibly infinite) handle decomposition. If $U$ is smoothable, then it has a handle decomposition by Morse theory [Wal16, Lemma 5.1.8]. If $n=5$, this is [Qui82, Theorem 2.3.1], and if $n\geq 6$, this is [KS77, Essay III, Theorem 2.1]. If $n\leq 3$, then it suffices to show that $U$ has a smooth structure. If $n=1$, this follows from the classification of $1$-manifolds [Lee11, Theorem 5.27]. If $n=2$, this is [Hat22, Theorem A]. If $n=3$, then by [Thu97, Theorem 3.10.8], it suffices to show that $U$ has a PL structure. This is [Ham76, Theorem 2].
  4. (Step 4)
    We finish off the proof by considering the case where $n=4$. Let $U\subset M$ be a nonempty finite set. For each path component $V\subset U$, choose $\left(k+1\right)$ points $p_{0}^{V},\dots,p_{k}^{V}\in V$, and set $A_{j}=\{p_{j}^{V}\mid V\in\pi_{0}\left(U\right)\}$ for $0\leq j\leq k$. Then $A_{0},\dots,A_{k}$ are disjoint closed subsets of $U$, and moreover for each nonempty finite set $S\subset\{0,\dots,k\}$, the components of the open set $U\setminus\bigcup_{i\in S}A_{i}$ are noncompact. Since connected, non-compact $4$-manifolds are smoothable [FQ90, 8.2], we deduce from Step 2 that $\bigcap_{i\in S}U\setminus A_{i}$ belongs to $\mathcal{U}$. Condition (3) now shows that $U$ belongs to $\mathcal{U}$, and the proof is complete.
Remark 3.17.
Our proof of Lemma 3.16 is a modification of Weiss’s proof of [Wei99, Theorem 5.1], where he essentially established the lemma in the case where $M$ is smoothable.
We now arrive at the proof of Theorems 3.10 and 3.11.
Proof of Theorem 3.10.
Clearly (2$\implies$(3$\implies$(4). It will therefore suffice to show that (1$\implies$(2) and (4$\implies$(1).
First we show that (1$\implies$(2). Suppose that $F$ is a left Kan extension of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Let $\mathcal{I}^{\prime}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}^{\prime}\to N\left(\operatorname{Open}\left(U\right)\right)$ be a functor satisfying the Weiss $k$-condition. For each $I\in\mathcal{I}^{\prime}$, set $U_{I}=\chi\left(I\right)$. We must show that the map
$\operatorname{colim}_{I\in\mathcal{I}^{\prime}}F\left(U_{I}\right)\to F\left(U\right)$
is an equivalence of $\mathcal{C}$. According to [Lur24, Tag 02MD] and [Lur24, Tag 02N9], there is a small category $\mathcal{I}$ and a final functor $N\left(\mathcal{I}\right)\to\mathcal{I}^{\prime}$ whose pullbacks are all final. Then the composite $\chi\circ p$ satisfies the Weiss $k$-condition, so it suffices to consider the case where $\mathcal{I}^{\prime}=N\left(\mathcal{I}\right)$.
Set $n=\dim M$. Using Lemma 3.13, we may assume that $F$ extends to a functor $\mathcal{M}\mathsf{fld}_{n/M}\to\mathcal{C}$ which is left Kan extended from its restriction to $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$. By Proposition 2.15, it suffices to show that the functor
$f:\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\overrightarrow{\times}_{\mathcal{M}\mathsf{fld}_{n/M}}N\left(\mathcal{I}\right)\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)^{/\iota}$
is final, where $\iota:U\hookrightarrow M$ denotes the inclusion. By Remark 2.14, the square
${{N(\operatorname{Disj}^{\leq k}(M))\overrightarrow{\times}_{N(\operatorname{Open}(M))}N(\mathcal{I})}}$${{N(\operatorname{Open}(M))^{/U}}}$${{\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\overrightarrow{\times}_{\mathcal{M}\mathsf{fld}_{n/M}}N(\mathcal{I})}}$${{(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k})^{/\iota}}}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{g^{\prime}}$$\scriptstyle{f^{\prime}}$
commutes up to natural equivalence. Since final maps have the right cancellation property [Cis19, Corollary 4.1.9], it will suffice to show that the functors $g$ and $g^{\prime}f^{\prime}$ are final functors.
We start by showing that that $g$ is final. By Propositions 2.12 and 2.23 (and using the equivalence of the ordinary and alternative slices [Lur09, Proposition 4.2.1.5]), it suffices to show that for each object $I\in\mathcal{I}$, the functor
$N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)_{/U_{I}}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)_{/\iota_{I}}$
is final, where $\iota_{I}:U_{I}\hookrightarrow M$ denotes the inclusion. Since the forgetful functor $\left(\mathcal{D}\mathsf{isk}_{n/M}\right)_{/\iota_{I}}\to\mathcal{D}\mathsf{isk}_{n/U_{I}}$ is a trivial fibration, we are reduced to showing that the functor
$N\left(\operatorname{Disj}^{\leq k}\left(U_{I}\right)\right)\to\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}$
is final. This is a consequence of Lemma 3.14.
Next, we show that the composite $g^{\prime}f^{\prime}$ is final. Set $\mathcal{X}=N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)\overrightarrow{\times}_{N\left(\operatorname{Open}\left(M\right)\right)}N\left(\mathcal{I}\right)$. By the equivalence of the ordinary and alternative slice, it suffices to show that the functor
$\mathcal{X}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
defined by $\left(V,I\right)\mapsto\left(V\hookrightarrow U\right)$ is final. For this, we use Lemma 3.14: It suffices to show that, for each finite subset $S\subset U$ of cardinality $\leq k$, the full subcategory $\mathcal{X}_{S}\subset\mathcal{X}$ spanned by the objects $\left(V,I\right)\in\mathcal{X}$ such that $S\subset V$ is weakly contractible. Consider the functor
$p:\mathcal{X}_{S}\to N\left(\mathcal{I}_{S}\right).$
The functor $p$ is a cocartesian fibration whose fibers are the nerves of cofiltered posets. It follows from [Lur09, Proposition 4.1.2.15] that $p$ is an initial functor. In particular, $p$ is a weak homotopy equivalence. Since $N\left(\mathcal{I}_{S}\right)$ is weakly contractible by hypothesis, this shows that $\mathcal{X}_{S}$ is weakly contractible. This completes the proof of (1$\implies$(2).
We now complete the proof by showing that (4$\implies$(1). Let $G$ be a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$, and let $\alpha:G\to F$ be a natural transformation which extends the identity natural transformation of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. We must show that $\alpha$ is a natural equivalence.
Call an open set $U\subset M$ good if the map $\alpha_{U}$ is an equivalence. We wish to show that every open set of $M$ is good. Since $F$ and $G$ are both $k$-excisive and exhaustive (for we already know that (1$\implies$(4)), the collection of good open subsets of $M$ satisfies the hypotheses of Lemma 3.16. Therefore, every open set of $M$ is good, and we are done. ∎
Proof of Theorem 3.11.
Clearly (2$\implies$(3). Also, we can prove (1$\implies$(2) in exactly the same way as we proved the implication (1$\implies$(2) of Theorem 3.10. It will therefore suffice to show that (3$\implies$(1).
Suppose that $F$ is a Weiss cosheaf. Let $G$ be a left Kan extension of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. Note that $G$ is a Weiss cosheaf because we have already shown that (1$\implies$(3). Let $\alpha:G\to F$ be a natural transformation which extends the identity natural transformation of $F|N\left(\operatorname{Disj}\left(M\right)\right)$. We must show that $\alpha$ is a natural equivalence.
Call an open set $U\subset M$ good if the map $\alpha_{U}$ is an equivalence. We wish to show that every open set of $M$ is good. We prove this in two steps.
Let $U\subset M$ be an open set homeomorphic to an open subset of $\mathbb{R}^{n}$, where $n=\dim M$. We claim that $U$ is good. Choose an embedding $U\hookrightarrow\mathbb{R}^{n}$ and identify $U$ with a subset of $\mathbb{R}^{n}$. Let $\operatorname{Cube}\left(U\right)$ denote the poset of open subsets of $U$ that are finite disjoint unions of open cubes lying in $U$, and let $\overline{\operatorname{Cube}\left(U\right)}$ denote the poset of open subsets of elements of $\operatorname{Cube}\left(U\right)$. Since $\operatorname{Cube}\left(U\right)$ is closed under finite intersection, the inclusion $N\left(\operatorname{Cube}\left(U\right)\right)\to N\left(\overline{\operatorname{Cube}\left(U\right)}\right)$ is final. Moreover, $\overline{\operatorname{Cube}\left(U\right)}$ is a covering sieve of $U$ in the Weiss topology on $\operatorname{Open}\left(M\right)$. Since every element of $\operatorname{Cube}\left(U\right)$ is good, it follows that $\alpha_{U}$ is an equivalence.
Next, let $U\subset M$ be an arbitrary subset. We claim that $U$ is good. Let $\mathcal{U}$ denote the set of all open sets of $U$ that are homeomorphic to an open subset of $\mathbb{R}^{n}$. By the result in the previous paragraph, every element in $\mathcal{U}$ is good. Moreover, $\mathcal{U}$ is a covering sieve of $U$ in the Weiss topology. It follows that $\alpha_{U}$ is an equivalence, and the proof is complete. ∎
3.4. Taylor Towers and Their Convergence
Now that we have completed the proof of Theorem 3.4, let us use it to construct the Taylor (co)tower of an isotopy functor and discuss when the tower converges to the original functor.
Let $M$ be a manifold and let $\mathcal{C}$ be an $\infty$-category with small colimits. Theorem 3.4 says that for each isotopy functor $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ and each $k\geq 0$, there is a best approximation of $F$ by $k$-excisive, exhaustive isotopy functor, namely, the left Kan extension $T_{k}F$ of $F|N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. This is the $k$th polynomial approximation of $F$. We now organize these approximations into a cotower
$T_{0}F\to T_{1}F\to\cdots\to F.$
For later discussions, we work in a slightly more general setting.
Definition 3.18.
Let $\mathcal{A}$ be an $\infty$-category equipped with a sequence $\mathcal{A}^{0}\subset\mathcal{A}^{1}\subset\cdots$ of full subcategories. Set $\mathcal{A}^{\infty}=\bigcup_{i}\mathcal{A}^{i}$. We let $\mathbb{T}\left(\mathcal{A}\right)\subset N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects $\left(i,A\right)$, where $i\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}_{i}$, together with the objects $\left(\infty,A\right)$, where $A\in\mathcal{A}$.
Let $\mathcal{C}$ be another $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. A functor $\mathbb{T}\left(F\right):N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}\to\mathcal{C}$ is called a Taylor cotower of $F$ (with respect to the subcategories $\{\mathcal{A}^{i}\}_{i\geq 0}$) if it is a left Kan extension of the composite
[…]
For each $k\in\mathbb{Z}_{\geq 0}\cup\{\infty\}$, we set $T_{k}F=\mathbb{T}\left(F\right)|\{k\}\times\mathcal{A}$. Note that for each $k\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}$, the inclusion
$\{\operatorname{id}_{k}\}\times\mathcal{A}_{/A}^{k}\hookrightarrow\mathbb{T}\left(\mathcal{A}\right)_{/\left(k,A\right)}$
is final (for it is a right adjoint), so the restriction $T_{k}F$ is a left Kan extension of $F|\mathcal{A}^{k}$. By convention, we will write $T_{-1}F$ for any functor which carries all objects of $\mathcal{A}$ to an initial object of $\mathcal{C}$.
A Taylor cotower $\mathbb{T}\left(F\right)$ of $F$ is said to be convergent if the functor $N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\to\operatorname{Fun}\left(\mathcal{A},\mathcal{C}\right)$ adjoint to $\mathbb{T}\left(F\right)$ is a colimit diagram.
Example 3.19.
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor. A Taylor cotower of $F$ is a Taylor cotower of $F$ with respect to the subcategories $\{\operatorname{Disj}^{\leq i}\left(M\right)\}_{i\geq 0}$. If $F$ is an isotopy functor, then for each $k\in\mathbb{Z}$, the $k$th level $T_{k}F$ of a Taylor cotower of $F$ is polynomial of degree $\leq k$ (Theorem 3.4).
Warning 3.20.
Despite the name, the constituents of Taylor cotowers of non-isotopy functors may fail to be polynomial. For example, consider the functor $F:N\left(\operatorname{Open}\left(\mathbb{R}\right)\right)\to N\left(\mathsf{Set}\right)$ defined by $F\left(U\right)=U\times U$. It is easy to check that $T_{1}F$ is not excisive.
The proposition below gives the characterization of convergent functors.
Proposition 3.21.
Let $\mathcal{A}$ be an $\infty$-category and let $\mathcal{A}_{0}\subset\mathcal{A}_{1}\subset\cdots$ be a sequence of full subcategories of $\mathcal{A}$. Set $\mathcal{A}_{\infty}=\bigcup_{i\geq 0}\mathcal{A}_{i}$. Let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ has a convergent Taylor cotower.
  2. (2)
    The functor $F$ is a left Kan extension of $F|\mathcal{A}_{\infty}$.
Proof.
We will use the transitivity of Kan extensions [Lur09, Proposition 4.3.2.8]. Let $T=\mathbb{T}\left(F\right)$ be a Taylor cotower of $F$. Since colimits in functor categories can be formed pointwise [Lur24, Tag 02X9], an easy finality argument shows that condition (1) is equivalent to the following condition:
  • (1)
    $T$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$.
Let $\mathcal{X}\subset N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects in $N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$ and the objects in {}×$\{\infty\}\times\mathcal{A}_{\infty}$. Since every object $A\in\mathcal{A}_{\infty}$ belongs to some $\mathcal{A}_{i}$, the functor $T|\mathcal{X}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$. Therefore, condition (1) is equivalent to the following condition:
  • (1′′)
    $T$ is a left Kan extension of $T|\mathcal{X}$.
On the other hand, $T|\mathcal{X}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}_{\infty}$, so (1′′) is equivalent to the following condition:
  • (1′′′)
    $T$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\cup\{\infty\}\right)\times\mathcal{A}_{\infty}$.
Since $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}$ is a left Kan extension of $T|N\left(\mathbb{Z}_{\geq 0}\right)\times\mathcal{A}_{\infty}$, condition (1′′′) is equivalent to condition (2), and we are done. ∎
Corollary 3.22.
Let $M$ be a manifold and let $\mathcal{C}$ be an $\infty$-category with small colimits. An isotopy functor $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ has a convergent Taylor cotower if and only if $F$ satisfies the equivalent conditions of Theorem 3.11.
Proof.
This follows from Proposition 3.21 and Theorem 3.11. ∎
4. Classification of Homogeneous Functors
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor which is polynomial of degree $\leq k$. We say that $F$ is homogeneous of degree $k$ if $F$ evaluates to a zero object at each element $U\in\operatorname{Disj}\left(M\right)$ with fewer than $k$ components, or equivalently, if the $\left(k-1\right)$th level $T_{k-1}F$ of a Taylor cotower of $F$ (Example 3.19) carries every object to a zero object. For example, every functor of polynomial $\leq 0$ is homogeneous of degree 00; a $1$-excisive isotopy functor is homogeneous of degree $1$ if and only if it is a cosheaf (by Theorem 3.10 and Example 3.8).
We are interested in homogeneous functors for the following reason: Suppose that $F$ is an isotopy functor. A cofiber of $T_{k-1}F\to T_{k}F$, called a $k$th homogeneous layer of $F$, is then an isotopy functor which is homogeneous of degree $\leq k$: It is an isotopy functor by Lemma 3.5; it is polynomial of degree $\leq k$ by Theorem 3.10; homogeneity is clear. Therefore, if $F$ has a convergent Taylor cotower, then up to extension problem we can reduce the analysis of $F$ to those of the homogeneous layers of $F$.
In this section, we prove that homogeneous functors are classified by unordered configuration spaces and give an interpretation of the homogeneous layer in terms of higher derivatives.
Notation 4.1.
Let $M$ be a manifold, and let $\mathcal{C}$ be a pointed $\infty$-category with small colimits. We let $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ denote the full subcategory of $\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the exhaustive isotopy functors that are homogeneous of degree $k$.
Recall that if $M$ is an $n$-manifold, then $I_{M}^{=k}$ denotes the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components (Notation 2.30). The nerve of $I_{M}^{=k}$ classifies homogeneous functors in the following sense:
Proposition 4.2.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $k\geq 0$. The inclusion $I_{M}^{=k}\subset\operatorname{Open}\left(M\right)$ induces a trivial fibration
$\theta:\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right).$
Proof.
By Theorem 3.4, the restriction functor
$p:\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$
is a trivial fibration. Let $\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\subset\operatorname{Fun}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$ denote the full subcategory spanned by the functors which carry every object in $N\left(\operatorname{Disj}^{\leq k-1}\left(M\right)\right)$ to a zero object. The functor
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)$
is a pullback of $p$, so it is a trivial fibration. On the other hand, Proposition 2.34 shows that the functor
$\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right)$
is a trivial fibration. Therefore, $\theta$ is a composition of trivial fibrations and hence is itself a trivial fibration. ∎
Now if $M$ is a manifold and $k\geq 0$, then a functor on $N\left(I_{M}^{=k}\right)$ is an isotopy functor if and only if it maps every morphism of $N\left(I_{M}^{=k}\right)$ to an equivalence. Therefore, Proposition 4.2 says that homogeneous functors of degree $k$ on a manifold $M$ can be classified in terms of the classifying space of $I_{M}^{=k}$, i.e., a localization of $N\left(I_{M}^{=k}\right)$ with respect to all morphisms [Lur24, Tag 01MY]. The following proposition identifies the homotopy type of this classifying space.
Proposition 4.3.
Let $M$ be a manifold. For each $k\geq 0$, there is a weak homotopy equivalence
$N\left(I_{M}^{=k}\right)\xrightarrow{\simeq}\operatorname{Sing}B_{k}\left(M\right)$
which carries an object $U\in I_{M}^{=k}$ to a point $\{p_{1},\dots,p_{k}\}\in B_{k}\left(M\right)$ which intersects every component of $U$.
Remark 4.4.
The identification of the homotopy type of $N\left(I_{M}^{=k}\right)$ dates back at least to [Wei99, Lemma 3.5] and has been reproved again and again (e.g., [AF15, Lemma 2.12]). The point of Proposition 4.3 is that there is an explicit weak homotopy equivalence, whose account seem to be lacking in the existing literature.
Proof.
The proposition is a consequence of the following assertion:
  • ($\ast$)
    Let $P\subset I_{M}^{=1}$ be a full subcategory which forms a basis of the topology on $M$. Then there is a weak homotopy equivalence
    $N\left(P\right)\to\operatorname{Sing}M$
    which carries each object $U\in N\left(P\right)$ to a point in $U$.
To see how ($\ast$) implies the proposition, notice that each element $U\in I_{M}^{=k}$ determines an element $B^{\prime}_{k}\left(U\right)\in I_{B_{k}\left(M\right)}^{=1}$, where $B^{\prime}_{k}\left(U\right)$ denotes the subset of $B_{k}\left(M\right)$ consisting of the points $\{p_{1},\dots,p_{k}\}\in B_{k}\left(M\right)$ which intersects every component of $U$. (This is consistent with Notation 2.30.) The map $U\mapsto B^{\prime}_{k}\left(U\right)$ identifies $I_{M}^{=k}$ with a full subcategory of $I_{B_{k}\left(M\right)}^{=1}$ which form a basis of the topology on $B_{k}\left(M\right)$. We can then apply ($\ast$) to $P=I_{M}^{=k}$ to obtain the proposition.
To prove ($\ast$), let $n$ denote the dimension of $M$. Let $\operatorname{Chart}\left(P,M\right)$ denote the category whose elements are the embeddings $\iota:\mathbb{R}^{n}\to M$ such that $\iota\left(\mathbb{R}^{n}\right)\in P$, and whose morphisms $\iota\to\iota^{\prime}$ are the continuous maps rendering the diagram
[…]
commutative. The forgetful functor $N\left(\operatorname{Chart}\left(P,M\right)\right)\to N\left(P\right)$ is a trivial fibration. It will therefore suffice to construct a weak homotopy equivalence $N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$ which carries each object $\iota\in N\left(\operatorname{Chart}\left(P,M\right)\right)$ to the point $\iota\left(0\right)\in\operatorname{Sing}M$.
Given integers $p,q\geq 0$ and continuous maps $\sigma:\left|\Delta^{p}\right|\to\mathbb{R}^{n}$ and $\tau:\left|\Delta^{q}\right|\to\mathbb{R}^{n}$, we define a map $\sigma\star\tau:\left|\Delta^{p+q+1}\right|\to\mathbb{R}^{n}$ by
$\sigma\left(s_{0},\dots,s_{p},t_{0},\dots,t_{q}\right)=\begin{cases}s\sigma\left(\frac{s_{0}}{s},\dots,\frac{s_{p}}{s}\right)+t\tau\left(\frac{t_{0}}{t},\dots,\frac{t_{q}}{t}\right)&\text{if }s,t\neq 0,\\ \tau\left(t_{0},\dots,t_{q}\right)&\text{if }s=0,\\ \sigma\left(s_{0},\dots,s_{p}\right)&\text{if }t=0,\end{cases}$
where we set $s=\sum_{i=0}^{p}s_{i}$ and $t=\sum_{i=1}^{q}t_{i}$. Note that if $f:\Delta^{p^{\prime}}\to\Delta^{p}$ and $g:\Delta^{q^{\prime}}\to\Delta^{q}$ are morphisms of simplicial sets, then $\left(\sigma\star\tau\right)\left(\left|f\star g\right|\right)=\left(\sigma\left|f\right|\right)\star\left(\tau\left|g\right|\right)$.
Define
$\Phi_{M}:N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$
as follows: Let $\sigma=\left(\iota_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{p}}\iota_{p}\right)$ be a $p$-simplex of $N\left(\operatorname{Chart}\left(P,M\right)\right)$. Define $\alpha_{\sigma}:\left|\Delta^{p}\right|\to\mathbb{R}^{n}$ inductively by
$\alpha_{\sigma}=\begin{cases}0&\text{if }p=0,\\ \left(f_{p}\alpha_{d_{p}\sigma}\right)\star 0&\text{if }p>0.\end{cases}$
One can show by induction on $p$ that for each $p>0$ and $0\leq i\leq p$, we have
$\alpha_{\sigma}\partial_{i}=\begin{cases}f_{p}\alpha_{d_{p}\sigma}&\text{if }i=p,\\ \alpha_{d_{i}\sigma}&\text{if }i<p,\end{cases}$
and that for each $0\leq i\leq p$ we have $\alpha_{\sigma}\sigma_{i}=\alpha_{s_{i}\sigma}$. Therefore, the assignment $\sigma\mapsto\iota_{p}\circ\alpha_{\sigma}$ determines a morphism of simplicial sets $N\left(\operatorname{Chart}\left(P,M\right)\right)\to\operatorname{Sing}M$, which we denote by $\Phi_{M}$. We claim that $\Phi_{M}$ is a weak homotopy equivalence.
For each $U\in P$, let $P_{U}\subset P$ denote the full subcategory spanned by the elements that are contained in $U$. Consider the diagram
$SSF:P\to\SS$
given by $U\mapsto N\left(\operatorname{Chart}\left(P_{U},U\right)\right)$. The colimit of this diagram is $N\left(\operatorname{Chart}\left(P,M\right)\right)$, so there is a commutative diagram
${{\operatorname{hocolim}_{U\in P}N(\operatorname{Chart}(P_{U},U))}}$${{N(\operatorname{Chart}(P,M))}}$${{\operatorname{hocolim}_{U\in P}\operatorname{Sing}(U)}}$${{\operatorname{Sing}(M).}}$$\scriptstyle{h}$$\scriptstyle{\Phi_{M}}$$\scriptstyle{g}$$\scriptstyle{f}$
Here $f$ is induced by the maps $\{\Phi_{U}\}_{U\in P}$. We must show that the map $\Phi$ is a weak homotopy equivalence. Since each $N\left(\operatorname{Chart}\left(P_{U},U\right)\right)$ has a terminal object, the map $f$ is a weak homotopy equivalence. The map $h$ is also a weak homotopy equivalence by [Lur17, Theorem A.3.1]. It will therefore suffice to show that the map $g$ is a weak homotopy equivalence. For this, it suffices to show that the diagram $F$ is projectively cofibrant. We will show that, in fact, for each pair of simplicial subsets $A\subset B\subset N\left(\operatorname{Chart}\left(P,M\right)\right)$, the map $A\cap F\to B\cap F$ is a projective cofibration. (Here $A\cap F$ denotes the diagram $SSP\to\SS$ given by $U\mapsto A\cap F\left(U\right)$.) Working simplex by simplex, it will suffice to consider the case where there is a pushout square
[…]
Depict $\sigma$ as $\iota_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{p}}\iota_{p}$, where each $\iota_{i}$ is an embedding $\mathbb{R}^{n}\to M$. Let $SSj:P^{\mathrm{op}}\to\operatorname{Fun}\left(P,\SS\right)$ denote the Yoneda embedding. Then the square
${{\partial\Delta^{p}\times j(\iota_{p}(\mathbb{R}^{n}))}}$${{A\cap F}}$${{\Delta^{p}\times j(\iota_{p}(\mathbb{R}^{n}))}}$${{B\cap F}}$
is cocartesian. The left vertical arrow is a projective cofibration, so the right vertical arrow is also a projective cofibration. The proof is now complete. ∎
Combining Propositions 4.2 and 4.3, we obtain:
Corollary 4.5.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $k\geq 0$. There is a zig-zag of categorical equivalences
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(N\left(I_{M}^{=k}\right),\mathcal{C}\right)\xleftarrow{\simeq}\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right).$
Propositions 4.2 and 4.3 lead to an interesting observation if we apply it to homogeneous layers of isotopy functors [^4] The author learned this interpretation from [Mun10].. Let $n\geq 0$, let $k\geq 1$, let $M$ be an $n$-manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be an isotopy functor. Let $L_{k}F$ denote a cofiber of the natural transformation $T_{k-1}F\to T_{k}F$ and let $\overline{L}_{k}F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ denote the corresponding functor. Let us compute the value of $\overline{L}_{k}F$ at a point$\{p_{1},\dots,p_{k}\}\in B_{k}M$. First, choose pairwise disjoint open sets $B_{1},\dots,B_{k}\subset M$ that are homeomorphic to $\mathbb{R}^{n}$, and set $U=B_{1}\cup\dots\cup B_{k}$. Then
$\overline{L}_{k}F\left(\{p_{1},\dots,p_{k}\}\right)$ $\simeq L_{k}F\left(U\right)$
$\simeq\mathrm{cofib}\left(T_{k-1}F\left(U\right)\to T_{k}F\left(U\right)\right).$
Since $T_{k-1}F$ is $\left(k-1\right)$-excisive, the map $T_{k-1}F\left(U\right)\to T_{k}F\left(U\right)$ can be identified with the map
$\theta:\operatorname{colim}_{S\subsetneq\{1,\dots,k\}}F\left(\bigcup_{i\in S}B_{i}\right)\to F\left(\bigcup_{i=1}^{k}B_{i}\right).$
A cofiber of $\theta$ is called a total cofiber of the $k$-dimensional cubical diagarm $S\mapsto F\left(\bigcup_{i\in S}B_{i}\right)$. Total cofibers admit an interpretation in terms of a derivative: For example if $k=2$, then the total homotopy cofiber can be written as
$\mathrm{cofib}\left(\mathrm{cofib}\left(F\left(\emptyset\right)\to F\left(B_{1}\right)\right)\to\operatorname{cofib}\left(F\left(B_{2}\right)\to F\left(B_{1}\cup B_{2}\right)\right)\right)$
which resembles the classical formula
$f^{\prime\prime}\left(0\right)=\operatorname{lim}_{h_{1},h_{2}\to 0}\frac{f\left(h_{1}+h_{2}\right)-f\left(h_{1}\right)-f\left(h_{2}\right)+f\left(0\right)}{h_{1}h_{2}}$
of the second derivative. Therefore, the values of $\overline{L}_{k}F$ are these “$k$th derivatives” of $F$, highlighting the connection with the Taylor expansion in ordinary calculus.
Remark 4.6.
Let $M$ be a smooth manifold and let $I_{\mathrm{sm},M}^{=k}$ denote the subcategory of $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)$ spanned by the smooth isotopy equivalences between objects with exactly $k$ components. Corollary 4.5 remains valid if we replace $I_{M}^{=k}$ and by $I_{\mathrm{sm},M}^{=k}$ and $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ by the full subcategory of $\operatorname{Fun}\left(N\left(\operatorname{Open}\left(M\right)\right),\mathcal{C}\right)$ spanned by the functors that are homogeneous of degree $k$, exhaustive, and carries smooth isotopy equivalences to equivalences, with essentially the same proof.
5. Context-Free Case
In many cases, functors subject to the analysis of manifold calculus are restrictions of functors defined on all of $\mathcal{M}\mathsf{fld}_{n}$ or its variants. The study of such functors is called context-free manifold calculus [Tur13]. Some of the results of Section 3 generalize to the context-free case, which we record in this section.
Definition 5.1.
Let $n,k\geq 0$. Let $\mathcal{C}$ be an an $\infty$-category and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor.
  1. (1)
    The Weiss $k$-topology on $\mathsf{Mfld}_{n}$ is the Grothendieck topology on $\mathsf{Mfld}_{n}$ such that for each object $M\in\mathsf{Mfld}_{n}$, a sieve $\mathcal{U}$ on $M$ is a covering sieve if and only if for each finite set $S\subset M$ of cardinality $\leq k$, there is an object $\left(U\to M\right)\in\mathcal{U}$ such that $S\subset U$. The intersection of the Weiss topologies on $M$ are called the Weiss topology.
  2. (2)
[…]
  3. (3)
    We say that $F$ is $k$-excisive, or polynomial of degree $\leq k$, if for each $n$-manifold $M$ and each pairwise disjoint closed sets $A_{0},\dots,A_{k}\subset M$, the map
    $\operatorname{colim}_{\emptyset\neq S\subset\{0,\dots,k\}}F\left(M\setminus\bigcup_{i\in S}A_{i}\right)\to F\left(M\right)$
    is an equivalence.
  4. (4)
    We say that $F$ is exhaustive if for each $n$-manifold $M$ and each increasing sequence $U_{0}\subset U_{1}\subset\cdots$ of open sets of $M$ which covers $M$, the map
    $\operatorname{colim}_{i}F\left(U_{i}\right)\to F\left(M\right)$
    is an equivalence.
  5. (5)
    We will write $\operatorname{Exc}_{\mathrm{exh}}^{k}\left(\mathcal{M}\mathsf{fld}_{n};\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ for the full subcategory spanned by the $k$-excisive, exhaustive functors $\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$.
  6. (6)
    The Taylor cotower of $F$ is the Taylor cotower of $F$ with respect to the sequence of full subcategories $\mathcal{D}\mathsf{isk}_{n}^{\leq 0}\subset\mathcal{D}\mathsf{isk}_{n}^{\leq 1}\subset\cdots$ (Definition 3.18).
Remark 5.2.
In light of the constructions in non-context-free manifold calculus, some readers find it more natural to work with functors on $\mathsf{Mfld}_{n}$ which carry isotopy equivalences to equivalences, rather than functors defined on $\mathcal{M}\mathsf{fld}_{n}$. There is a reason for this: The localization theorem for isotopy equivalences (Theorem 2.24), which formed a basis of our argument in non-context-free manifold calculus, is significantly more subtle in the context-free case. We will come back to this point in Appendix A.
The following theorems, which are the main result of this section, are analogs of Theorems 3.4, 3.10 and 3.11 in the context-free case.
Theorem 5.3.
Let $n,k\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss $k$-condition (Definition 3.9), the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
    is an equivalence.
  3. (3)
    $F$ is a Weiss $k$-cosheaf.
  4. (4)
    $F$ is $k$-excisive and exhaustive.
In particular, the inclusion $\operatorname{Exc}_{\mathrm{exh}}^{k}\left(\mathcal{M}\mathsf{fld}_{n};\mathcal{C}\right)\hookrightarrow\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ admits a right adjoint, which carries a functor $F\in\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ to a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$.
Theorem 5.4.
Let $n\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss condition (Definition 3.9), the map
    $\operatorname{colim}_{I\in N\left(\mathcal{I}\right)}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
    is an equivalence.
  3. (3)
    $F$ is a Weiss cosheaf.
Remark 5.5.
Theorems 5.3 and 5.4 remain valid if we replace $\mathcal{M}\mathsf{fld}_{n}$, $\mathcal{D}\mathsf{isk}_{n}$, and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ by $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$, and $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, with essentially the same proof.
The proofs of Theorems 5.3 and 5.4 are very similar, so we will focus on Theorem 5.3.
Proof of Theorem 5.3.
Clearly (2$\implies$(3$\implies$(4), so it suffices to show that (1$\implies$(2) and that (4$\implies$(1).
First we show that (1$\implies$(2). Suppose that $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$. Let $M$ be an $n$-manifold and let $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ be a functor of small categories satisfying the Weiss $k$-condition. We wish to show that the map
$\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
is an equivalence. In light of Theorem 3.10, it suffices to show that the composite
[…]
is a left Kan extension of $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Since $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$, it suffices to show that, for each open set $U\subset M$, the map
$N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)_{/U}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
is final. This follows from Lemma 3.14 (or from Theorem 2.24, for localization functors are final [Lur24, Tag 02N9]).
Next we show that (4$\implies$(1). Let $G$ denote a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$. We wish to show that the induced natural transformation $\alpha:G\to F$ is a natural equivalence. It suffices to show that, for each $n$-manifold $M$, the map $\alpha\iota:G\iota\to F\iota$ is a natural equivalence, where $\iota:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}\mathsf{fld}_{n}$ denotes the inclusion. By Theorem 3.10, the functor $F\iota$ is a left Kan extension of its restriction to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. As we saw in the previous paragraph, the functor $G\iota$ is also left Kan extensions of its restriction to $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$. Since the components of $\alpha\iota$ at each object in $N\left(\operatorname{Disj}^{\leq k}\left(M\right)\right)$ are equivalences, we deduce that $\alpha\iota$ must be an equivalence. The proof is now complete. ∎
As a corollary of the theorems, we obtain the following structural result on Taylor cotowers.
Corollary 5.6.
Let $n\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. Then:
  1. (1)
    For each $k\geq 0$, the $k$th level $T_{k}F$ of the Taylor cotower of $F$ is polynomial of degree $\leq k$ and exhaustive.
  2. (2)
    The Taylor cotower of $F$ is convergent if and only if it satisfies the equivalent conditions of Theorem 5.4.
Proof.
Part (1) follows from Theorem 5.3. Part (2) is a consequence of Proposition 3.21. ∎
Remark 5.7.
Theorem 5.3 admits a following variant: Suppose we are given a right fibration $\pi:\mathcal{M}\to\mathcal{M}\mathsf{fld}_{n}$. Set $\mathcal{D}=\mathcal{D}\mathsf{isk}_{n}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$ and $\mathcal{D}^{\leq k}=\mathcal{D}\mathsf{isk}_{n}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$. Given an $\infty$-category $\mathcal{C}$ with small colimits, we say that a functor $F:\mathcal{M}\to\mathcal{C}$ is a Weiss $k$-cosheaf if for each $n$-manifold $M$ and each functor $N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}$ rendering the diagram
${{\mathcal{M}}}$${{N(\operatorname{Open}(M))}}$${{\mathcal{M}\mathsf{fld}_{n}}}$
[…]
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}^{\leq k}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss $k$-condition (Definition 3.9), the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(f\left(\chi\left(I\right)\right)\right)\to F\left(f\left(M\right)\right)$
    is an equivalence, where $f:\operatorname{Open}\left(M\right)\to\mathcal{M}$ is any functor lifting $\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{n}$.
  3. (3)
    $F$ is a Weiss $k$-cosheaf.
  4. (4)
    $F$ is $k$-excisive and exhaustive.
The proof is essentially the same as that of Theorem 5.3, using the following observations:
  • For each object $\overline{M}\in\mathcal{M}$ with image $M\in\mathcal{M}\mathsf{fld}_{n}$, the functor $\mathcal{D}_{/\overline{M}}\to\mathcal{D}\mathsf{isk}_{n/M}$ is a trivial fibration (because $\pi$ is a right fibration).
[…]
We can similarly prove a variant of Theorem 5.4 for $\mathcal{M}$ and $\mathcal{D}$; we can also consider right fibrations over $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$.
One prominent source of right fibrations over $\mathcal{M}\mathsf{fld}_{n}$ comes from framed manifolds. Let $B\mathrm{Top}\left(n\right)\subset\mathcal{M}\mathsf{fld}_{n}$ denote the full subcategory spanned by $\mathbb{R}^{n}$. Given a map $p:B\to B\mathrm{Top}\left(n\right)$ of Kan complexes, the $\infty$-category of $B$-framed $n$-manifold [AF15, Definition 2.7] is modeled by the fiber product
$\mathcal{M}\mathsf{fld}_{n}^{B}=\mathcal{M}\mathsf{fld}_{n}\times_{\mathcal{S}_{/B\mathrm{Top}\left(n\right)}}\mathcal{S}_{/p}.$
Since the functor $\mathcal{S}_{/p}\to\mathcal{S}_{/B\mathrm{Top}\left(n\right)}$ is a right fibration, so is its pullback $\mathcal{M}\mathsf{fld}_{n}^{B}\to\mathcal{M}\mathsf{fld}_{n}$. Therefore, we have analogs of Theorem 5.3 and 5.4 for $B$-framed manifolds and $B$-framed disks.
Appendix A (Non)-Localization Theorems
The localization theorem for isotopy equivalences (Theorem 2.24) states that, for any $n$-manifold $M$, the functors
$\mathsf{Disk}_{n/M}\to\mathcal{D}\mathsf{isk}_{n/M},\,\mathsf{Disk}_{n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$
are localizations with respect to isotopy equivalences; there is also a similar result in the smooth category (Remark 2.31). As we saw in Section 3, these results are crucially important in the development of non-context-free manifold calculus. In this section, we consider the context-free analog of the localization theorem. The results in this section will not be used elsewhere in this paper.
It turns out that the localization theorem is a bit nuanced in the context-free case. (To the author’s knowledge, this was first observed by David Ayala and John Francis in [AF20, 2.2.13].) The validity of the theorem depends on which category (topological or smooth) we work in, as the following theorem shows.
Theorem A.1.
Let $n,k\geq 1$.
  1. (1)
    Both of the functors $\mathsf{Disk}_{n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ and $\mathsf{Disk}_{n}\to\mathcal{D}\mathsf{isk}_{n}$ are localizations with respect to isotopy equivalences.
  2. (2)
    Neither of the functors $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$ and $\mathsf{Disk}_{\mathrm{sm},n}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$ is a localization with respect to smooth isotopy equivalences.
Proof.
We begin with (1). We will focus on the functor $\mathsf{Disk}_{n}\to\mathcal{D}\mathsf{isk}_{n}$; the proof that the functor $\mathsf{Disk}_{n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ is a localization is similar. By Proposition 2.27 and Lemma 2.29, it suffices to show that the functor
$\mathsf{Disk}_{n}\times_{\mathcal{D}\mathsf{isk}_{n}}\mathcal{D}\mathsf{isk}_{n}^{\simeq}\to\mathcal{D}\mathsf{isk}_{n}^{\simeq}$
is a weak homotopy equivalence. For convenience, in this proof we will replace $\mathsf{Disk}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}$ by their full subcategories spanned by the objects $\{\mathbb{R}^{n}\times\{1,\dots,k\}\}_{k=0,1,\dots}$ and still denote them by $\mathsf{Disk}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}$.
Regard the $k$th symmetric group $\Sigma_{k}$ as a category with one object $\{1,\dots,k\}$ with morphism given by bijections, and let $N\Sigma_{k}$ denote its nerve. We consider the commutative diagram
${{\mathsf{Disk}_{n}\times_{\mathcal{D}\mathsf{isk}_{n}}\mathcal{D}\mathsf{isk}_{n}^{\simeq}}}$${{\mathcal{D}\mathsf{isk}_{n}^{\simeq}}}$${{\coprod_{k\geq 0}N\Sigma_{k}}}$$\scriptstyle{q}$$\scriptstyle{p}$
where $p$ and $q$ are given by $f\mapsto\pi_{0}\left(f\right)$. The functors $p$ and $q$ are cartesian fibrations. Since every morphism of $\coprod_{k\geq 0}N\Sigma_{k}$ is an equivalence, both $p$ and $q$ are locally constant (Definition 2.25). Therefore, it suffices to show that for each $k\geq 0$, the map
$\theta_{k}:p^{-1}\left(\{1,\dots,k\}\right)\to q^{-1}\left(\{1,\dots,k\}\right)$
is a weak homotopy equivalence.
We can identify the map $\theta_{k}$ with the map
$\left(N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\right)^{k}\to\left(N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)\right)^{k},$
where $N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$ denotes the homotopy coherent nerve of the topological monoid $\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$ (regarded as a topological category with a single object) and $N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}$ denotes the nerve of the same category with the discrete topology. It thus suffices to show that the map
$N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\to N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$
is a weak homotopy equivalence. This is a consequence of McDuff’s theorem [McD80, Corollary 2.15]. (For a comparison between the homotopy coherent nerve and McDuff’s model of classifying spaces, see [Ara24, Corollary 4.2].)
Next, we prove (2). As in the previous paragraph, it suffices to show that the map
$N\operatorname{Emb}_{\mathrm{sm}}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\to N\operatorname{Emb}_{\mathrm{sm}}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$
is not a weak homotopy equivalence. Arguing as in [McD80, Proof of Theorem 1.1], we can reduce this to showing that the map $B\Gamma_{n}^{\infty}\to BGL_{n}\left(\mathbb{R}\right)$ is not a weak homotopy equivalence of topological spaces, where $\Gamma_{n}^{\infty}$ denotes Haefliger’s groupoid [Hae71, p.143] for codimension $n$ smooth foliations. (See also Subsection A.1.) This is a consequence of Bott’s theorem [Bot70] (see [Hae71, p. 143, I.8 (a)]). ∎
A.1. A Remark on Classifying Spaces of Groupoids Internal to $\mathsf{Top}$
A little care is necessary in the final step of Theorem A.1, for in [McD80] and [Hae71], McDuff and Haefliger use different models of classifying spaces of groupoids internal to the category $\mathsf{Top}$ of topological spaces. The equivalence of the two models is probably well-known, but the author is not aware of a convenient reference for this. Thus we record a proof below.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. Abusing notation, let $\mathcal{G}$ denote the simplicial topological space corresponding to $\mathcal{G}$. (Thus $\mathcal{G}_{1}=\operatorname{mor}\mathcal{G}$, $\mathcal{G}_{2}=\operatorname{mor}\mathcal{G}\times_{\operatorname{ob}\mathcal{G}}\operatorname{mor}\mathcal{G}$, etc.) McDuff and Haefliger define the classifying space of $\mathcal{G}$ as follows:
  • In [McD80], McDuff defines the classifying space of $\mathcal{G}$ to be the fat realization of the simplicial topological space $\mathcal{G}$. We denote this topological space by $B_{{\rm MD}}\left(\mathcal{G}\right)$ [^5] This model was introduced by Segal [Seg68]..
  • Haefliger’s model is slightly more complicated. First, consider the space $E_{{\rm Hae}}\mathcal{G}$ whose points are the equivalence classes of formal expressions of the form
    $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$
    where $g_{0},g_{1},\dots$ are morphisms of $\mathcal{G}$ with a common codomain and $t_{0},t_{1},\dots$ are non-negative real numbers, all but finitely many of which are zero, and $\sum_{i=0}^{\infty}t_{i}=1$. Two expressions $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$ and $\left(t^{\prime}_{0},g^{\prime}_{0},t^{\prime}_{1},g^{\prime}_{1},\dots\right)$ are equivalent if $t_{i}=t_{i}^{\prime}$ for all $i$ and $g_{i}=g^{\prime}_{i}$ whenever $t_{i}>0$. The equivalence class of $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$ will be denoted by $\bigoplus_{i\geq 0}t_{i}g_{i}=t_{0}g_{0}\oplus t_{1}g_{1}\oplus\cdots$. We will write $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$ for the set map $\tau_{i}\left(\bigoplus_{j\geq 0}t_{j}g_{j}\right)=t_{i}$, and define $p_{i}:\tau_{i}^{-1}\left((0,1]\right)\to\operatorname{mor}\mathcal{G}$ by $p_{i}\left(\bigoplus_{j\geq 0}t_{j}g_{j}\right)=g_{i}$. We topologize $E_{{\rm Hae}}\mathcal{G}$ so that it has the following universal property: If $X$ is a topological space, then a set map $X\to E_{{\rm Hae}}\mathcal{G}$ is continuous if and only if the composites $\tau_{i}f:X\to[0,1]$ and $p_{i}f:\left(\tau_{i}f\right)^{-1}\left((0,1]\right)\to\operatorname{mor}\mathcal{G}$ are continuous.
    Now introduce an equivalence relation on $E_{{\rm Hae}}\mathcal{G}$ as follows: Two points $\bigoplus_{j\geq 0}t_{j}g_{j}$ and $\bigoplus_{j\geq 0}t^{\prime}_{j}g^{\prime}_{j}$ are equivalent if there is some morphism $g$ of $\mathcal{G}$ such that $\bigoplus_{j\geq 0}t_{j}gg_{j}=\bigoplus_{j\geq 0}t_{j}gg^{\prime}_{j}$. The resulting quotient space will be denoted by $B_{{\rm Hae}}\mathcal{G}$. We still denote by $\tau_{i}:B_{{\rm Hae}}\mathcal{G}\to[0,1]$ the map induced by $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$. The space $B_{{\rm Hae}}\mathcal{G}$ is the model of the classifying space of $\mathcal{G}$ used by Haefliger in [Hae71] [^6] This model was introduced by Milnor in [Mil56]..
These models are equivalent in the following sense:
Proposition A.2.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. There is a zig-zag of weak homotopy equivalences between $B_{{\rm MD}}\mathcal{G}$ and $B_{{\rm Hae}}\mathcal{G}$, which is natural in $\mathcal{G}$.
The proof of Proposition A.2 relies on a lemma. Write $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)$ for the semisimplicial set of nondegenerate simplices of the nerve of the poset $\mathbb{Z}_{\geq 0}$.
Lemma A.3.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. There is a bijective continuous map
$\phi:\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\to B_{{\rm Hae}}\mathcal{G},$
which is a homotopy equivalence.
Proof.
We will write $X=\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|$. By definition, a point of $X$ can be represented by a sequence
$\left(\left(t_{0},\dots,t_{n}\right),x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n},k_{0}<\cdots<k_{n}\right),$
where $\left(t_{0},\dots,t_{n}\right)$ is a point of $\left|\Delta^{n}\right|$, $x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n}$ is a point of $\left(N\mathcal{G}\right)_{n}$, and $k_{0}<\cdots<k_{n}$ is an element of $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)_{n}$. We declare that $\phi$ carries such a point to the point
$\left(\bigoplus_{0\leq i<k_{0}}0\right)\oplus t_{0}f_{n}\dots f_{1}\oplus\left(\bigoplus_{k_{0}<i<k_{1}}0\right)\oplus t_{1}f_{n}\dots f_{2}\oplus\cdots\oplus t_{n}\operatorname{id}_{x_{n}}\oplus\left(\bigoplus_{i<k_{n}}0\right).$
This defines a bijection $\phi:X\to B_{{\rm Hae}}\mathcal{G}$, which is continuous by the definition of the topology on $B\mathcal{G}$.
To show that $\phi$ is a homotopy equivalence, we first construct a homotopy inverse of $\phi$. Use [tD08, 13.1.7] to find a partition of unity $\left(\psi_{j}\right)_{j=0}^{\infty}$ on $B\mathcal{G}$ subordinate to the cover $\left(\tau_{j}^{-1}(0,1]\right)_{j=0}^{\infty}$. We define a set map $\psi:B_{{\rm Hae}}\mathcal{G}\to X$ as follows: Let $b=\bigoplus_{i=0}^{\infty}t_{i}g_{i}\in B_{{\rm Hae}}\mathcal{G}$ be a point. Let $k_{0}<\dots<k_{n}$ be the enumeration of the integers $i\geq 0$ such that $t_{i}>0$. Then $\psi\left(b\right)$ is represented by the point
$\left(\left(\psi_{k_{0}}\left(b\right)t_{k_{0}},\dots,\psi_{k_{n}}\left(b\right)t_{k_{n}}\right),x_{0}\xrightarrow{g_{k_{1}}^{-1}g_{k_{0}}}\cdots\xrightarrow{g_{k_{n}}^{-1}g_{k_{n-1}}}x_{n},k_{0}<\dots<k_{n}\right),$
where $x_{i}$ denotes the domain of $g_{k_{i}}$. To check the continuity of $\psi$, choose a neighborhood $U$ of $b$ such that $\psi_{j}|U=0$ for all $j\in\mathbb{Z}_{\geq 0}\setminus\{k_{0},\dots,k_{n}\}$. (This is possible because $\operatorname{supp}\psi_{j}\subset\tau_{j}^{-1}\left((0,1]\right)$ and the supports of $\{\psi_{j}\}_{j}$ are locally finite.) Without loss of generality, we may assume that $\tau_{k_{0}},\dots,\tau_{k_{n}}$ are positive on $U$. Then on $U$, the map $\psi$ can be written as a composite
$U\to\left|\Delta^{k}\right|\times\mathcal{G}_{n}\times\{k_{0}<\cdots<k_{n}\}\to X,$
each of which is continuous. This proves that $\psi$ is continuous.
We claim that $\psi$ is a homotopy inverse of $\phi$. Define a map $h:X\times[0,1]\to X$ as follows: Let $x=[\left(t_{0},\dots,t_{n}\right),\sigma,k_{0}<\cdots<k_{n}]$ be a point of $X$ and let $0\leq s\leq 1$. Then $h\left(x,s\right)$ is represented by
$\left(\left(\left(\left(1-s\right)+s\psi_{k_{0}}\left(\phi\left(x\right)\right)\right)t_{0},\dots,\left(\left(1-s\right)+s\psi_{k_{0}}\left(\phi\left(x\right)\right)\right)t_{n}\right),\sigma,k_{0}<\dots<k_{n}\right).$
This defines a homotopy from $\operatorname{id}_{X}$ to $\psi\phi$. Likewise, there is a map $B_{{\rm Hae}}\mathcal{G}\times[0,1]\to B_{{\rm Hae}}\mathcal{G}$ given by
$\left(b,s\right)=\left(\bigoplus_{i\geq 0}t_{i}g_{i},s\right)\mapsto\bigoplus_{i\geq 0}\left(\left(1-s\right)+s\psi_{i}\left(b\right)\right)t_{i}g_{i},$
which is a homotopy from $\operatorname{id}_{B\mathcal{G}}$ to $\phi\psi$. The claim follows. ∎
Proof of A.2.
Consider the maps
$B_{{\rm MD}}\mathcal{G}=\left\|\mathcal{G}\right\|\xleftarrow{\phi^{\prime}}\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\xrightarrow{\phi}B_{{\rm Hae}}\mathcal{G},$
where $\phi$ is the map of Lemma A.3 and $\phi^{\prime}$ is induced by the projection. We claim that these maps are weak homotopy equivalences. The map $\phi$ is a weak homotopy equivalence by Lemma A.3. For the map $\phi^{\prime}$, notice that the map
$\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\to\left\|\mathcal{G}\right\|\times\left\|N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|$
is a weak homotopy equivalence, for fat realizations preserve products up to weak homotopy equivalences [ERW19, Theorem 7.2]. The claim now follows from the observation that the topological space $\left\|N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\cong\left|N\left(\mathbb{Z}_{\geq 0}\right)\right|$ is contractible. ∎