### v3

# A Context for Manifold Calculus

Abstract.
We develop Weiss’s manifold calculus in the setting of $\infty$-categories, where we allow the target $\infty$-category to be any $\infty$-category with small limits. We will establish the connection between polynomial functors, Kan extensions, and Weiss sheaves, and will classify homogeneous functors. We will also generalize Weiss and Boavida de Brito’s theorem to functors taking values in arbitrary $\infty$-categories with small limits.
2020 Mathematics Subject Classification:
55U35, 57R40, 18F50
Introduction
Let $M$ be a smooth $n$-manifold, and let $\operatorname{Open}\left(M\right)$ be the poset of open sets of $M$. The idea of Weiss’s manifold calculus is to study space-valued presheaves $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ on $M$ by constructing a tower of presheaves
$F\to\cdots\to T_{k}F\to\cdots\to T_{1}F\to T_{0}F.$
Each $T_{k}F$ is the best “polynomial functor” of degree $\leq k$ that approximates $F$. Borrowing an analogy from calculus, the tower is called the Taylor tower of $F$ and the functors $T_{k}F$ the polynomial approximations of $F$.
The Taylor tower is determined by the definition of polynomials, so we must make this explicit. There is a heuristic in homotopy theory that homology behaves like linear (or affine) functions. A homology theory is more or less characterized by the excision axiom, so we make the following definition:
Definition 0.1.
[Wei99, Definition 2.2] We say that $F$ is polynomial of degree $\leq k$, or $k$-excisive, if for each open set $U\subset M$ and pairwise disjoint closed subsets $A_{0},\dots,A_{k}$ of $U$, the map
$F\left(U\right)\to\underset{\emptyset\subsetneq S\subset\{0,\dots,k\}}{\operatorname{holim}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
is a homotopy equivalence.
For example, $1$-excisivity is equivalent to the following condition: For every open set $U\subset M$ and every pair of disjoint closed sets $A_{0},A_{1}\subset U$, the square
[…]
is homotopy cartesian. This corresponds to the conventional usage of the term “excisivity,” hence justifying our terminology.
We also want to ensure that $F$ interacts well with the smooth structure on $M$, and also that $F$ is “continuous.” We thus make the following definition:
Definition 0.2.
[…]
  • Let $U\subset V$ be an inclusion of open sets of $M$ which is a smooth isotopy equivalence. Then the map $F\left(V\right)\to F\left(U\right)$ is a homotopy equivalence.
  • Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of open subsets of $M$. Then the map $F(\bigcup_{i\geq 0}U_{i})\to\operatorname{holim}_{i}F\left(U_{i}\right)$ is a homotopy equivalence.
With these definitions, we can now define polynomial approximations.
Definition 0.3.
Let $F,G:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be functors that carry smooth isotopy equivalences to homotopy equivalences. A natural transformation $\alpha:F\to G$ is said to exhibit $G$ as a $k$th polynomial approximation of $F$ if $G$ is polynomial of degree $\leq k$ and $\alpha$ is homotopically initial among the natural transformations to good polynomial functors of degree $\leq k$. In this case, we will write $G=T_{k}F$.
At this point, several questions come up naturally:
Question 0.4.
When does a good presheaf on $M$ admit polynomial approximations? If it does, how do we construct the approximations?
Question 0.5.
Can we classify polynomial functors?
Suppose now that $F$ takes values in the category of pointed spaces and that we have constructed its polynomial approximations. In good cases, the analysis of $F$ reduces, up to extension problems, to studying the fibers $L_{k}F=\mathrm{hofib}\left(T_{k}F\to T_{k-1}F\right)$. The functor $L_{k}F$ has the special property that it is homogeneous: It is polynomial of degree $\leq k$, and its $\left(k-1\right)$th polynomial approximation vanishes. This leads to the following question:
Question 0.6.
Can we classify homogeneous functors?
In [Wei99], Weiss gave complete answers to these questions. We now review his answers as two separate theorems (Theorems 0.7 and 0.8).
The first theorem answers Questions 0.4 and 0.5. We will let $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)$ denote the subset of $\operatorname{Open}\left(M\right)$ consisting of the elements that are diffeomorphic to $\mathbb{R}^{n}\times S$, where $S$ is a finite set of cardinality at most $k$.
Theorem 0.7.
[Wei99, Lemma 3.8, Theorem 4.1, Theorem 5.1] Let $G:\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be a functor carrying isotopy equivalences to homotopy equivalences. The homotopy right Kan extension of $G$ along the inclusion $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$ is good and $k$-excisive. Moreover, every good polynomial functor of degree $\leq k$ arises in this way.
In particular, every good space-valued presheaf $F$ on $M$ admits polynomial approximations in all degrees, constructed by homotopy right Kan extensions.
And the second theorem addresses Question 0.6:
Theorem 0.8.
[Wei99, Theorem 8.5] Homogeneous functors of degree $k$ taking values in pointed spaces admit a classification in terms of fibrations over $B_{k}\left(M\right)$ equipped with sections near the “fat diagonal”, where $B_{k}\left(M\right)$ denotes the space of unordered configurations of $k$ points in $M$.
Most of the things (perhaps except for Theorem 0.8) we have discussed so far make sense for presheaves taking values in categories other than that of spaces. Moreover, it seems quite natural to consider this generalization. For example, if one is interested in (co)homology, then it is not hard to imagine that presheaves of (co)chain complexes or spectra would be very useful. As such, numerous studies consider such generalized manifold calculus; see [Wei04, RW14, HK23, SAS18, TS25] for instance. In this paper, we will answer Questions 0.4, 0.5, and 0.6 in the generalized setting, and aim to set a foundation of generalized manifold calculus with arbitrary targets.
Here is what we will do. We will generalize Theorems 0.7 and 0.8 to presheaves on $M$ taking values in an arbitrary $\infty$-category $\mathcal{C}$ with small limits. It is straightforward to state a generalization of Theorem 0.7: The definitions of good functors, polynomial functors, polynomial approximations, and homogeneous functors directly carry over to $\mathcal{C}$-valued presheaves; we just have to replace homotopy limits by $\infty$-categorical limits. So a generalization of Theorem 0.7 will be as follows.
Theorem 0.9 (Theorem 1.3).
The right Kan extension functor restricts to a categorical equivalence
$\operatorname{Fun}_{{\rm istp}}\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
where:
  • $\operatorname{Fun}_{{\rm istp}}\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$ denotes the $\infty$-category of functors $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\to\mathcal{C}$ carrying smooth isotopy equivalences to equivalences; and
  • $\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of $k$-excisive good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$.
In particular, every good functor $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ admits a $k$th polynomial approximation, which is the right Kan extension of $F|\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}}$.
To generalize Theorem 0.8, recall that a fibration over $B_{k}\left(M\right)$ is equivalent (via the $\infty$-categorical Grothendieck construction [Lur09, 2.2.1.2]) to a functor $\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{S}$, where $\mathcal{S}$ denotes the $\infty$-category of spaces. With this in mind, we generalize Theorem 0.8 as follows:
Theorem 0.10 (Corollary 3.5, Theorem 3.16).
Suppose that $\mathcal{C}$ is pointed. There is an equivalence of $\infty$-categories
$\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\simeq\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
where $\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ that are of homogeneous of degree $k$.
Remark 0.11.
Theorem 0.10 says nothing about the fat diagonal, but a more precise statement involving the fat diagonal statement will follow. (See Theorem 3.16.) In particular, when applied to the case where $\mathcal{C}$ is the $\infty$-category $\mathcal{S}_{\ast}$ of pointed spaces, the theorem recovers Weiss’s classification of homogeneous functors.
Remark 0.12.
Most of the statements of Theorems 0.9 and 0.10 make sense even when $M$ does not have a smooth structure. Because of this, we will work mostly with topological manifolds, explaining how various statements will be affected in the presence of smooth and PL structures here and there.
Remark 0.13.
In [BdBW13], Boavida de Brito and Weiss established an analog of Theorem 0.8 for context-free [^1] The word “context” here has no relation with the title of this paper. manifold calculus, i.e., calculus of functors defined on the category of all smooth manifolds of a fixed dimension, not just on the poset of open sets on a single manifold. We will also state and prove a generalization of this in Section 4.
Remark 0.14.
Our paper is thematically very similar to the works of Songhafouo Tsopméné, Arnaud, and Stanley[SAS18, SAS19, TS25], in which the authors ask the extent to which manifold calculus can be developed for presheaves taking values in categories other than spaces. (They mainly work in the setting of model categories and employ very different technique than ours.) Our theorems will recover the results in their paper and solve their conjecture. More specifically:
  • By applying Theorem 0.9 and 0.10 to the case where $\mathcal{C}$ is the underlying $\infty$-category of a simplicial model category, we recover the main results of [SAS18].
  • By applying Theorem 0.10 to the case where $\mathcal{C}$ is an ordinary category, we recover the main result of [SAS19].
  • By applying Theorem 0.10 to the case where $\mathcal{C}$ is the underlying $\infty$-category of a simplicial model category $\mathbf{C}$, we recover the main result and a conjecture of [TS25]: For each object $X\in\mathbf{C}$, the homotopy classes of maps $B_{k}\left(M\right)\to B\mathrm{hAut}\left(X\right)$ are in bijection with the weak equivalence class good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathbf{C}$ that are of homogeneous of degree $k$ and carrying disjoint union of $k$ open balls to objects weakly equivalent to $X$.
At first glance, one might expect that proofs of Theorem 0.9 and 0.10 are straightforward, requiring only a minor modification of Weiss’s argument. While this is true to some extent, several parts of Weiss’s original proof rely critically on concrete constructions of spaces, which do not generalize easily. Moreover, our theorems is slightly stronger than Weiss’s results, in that it gives an equivalence of $\infty$-categories, and this added generality introduces further complications. To overcome these challenges, we develop and apply more sophisticated $\infty$-categorical techniques.
Theorem 0.9 and 0.10 have a wide array of applications outside what has traditionally been regarded as part of manifold calculus. For example, (part of) the theory of factorization homology [AF20] deals with functors of the form
$H:\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}\to\mathcal{C},$
where $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denotes the $\infty$-category of smooth $n$-manifolds and smooth embeddings. By precomposing the functor $\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, we obtain a $\mathcal{C}^{\mathrm{op}}$-valued precosheaf $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}^{\mathrm{op}}$. An observation of Ayala and Francis [AF20, Corollary 2.2.24] (which will be reproved in Subsection 2) essentially asserts that the functor $F$ is the limit of its Taylor tower, in the sense that the map $F\to\operatorname{lim}_{k}T_{k}F$ is an equivalence. This means that invariants of manifolds realized by factorization homology, which include a very interesting class of invariants [AF20], can be approximated by their Taylor approximations. It is crucial here to have a flexibility in the target category, which this paper offers, because invariants can arise in many different forms, such as spectra, chain complexes, or even $\infty$-operads [Ara24b].
Outline of the Paper
This paper consists of 5 sections. Section 1 concerns the existence and classification of polynomial functors. In Section 2, we construct and consider the convergence problem of Taylor towers. Section 3 classifies homogeneous functors. In Section 4, we prove a generalization of the results in Sections 1 and 2 for context-free manifold calculus. Section 5, discusses how these ideas extend to manifolds with boundary
The paper is also accompanied by an appendix, consisting of three sections. The purpose of the appendix is to record miscellaneous results used in the main body of the paper, so it should be consulted only when the need arises.
Acknowledgment
I thank Michael Weiss for carefully explaining some of the key ideas of his papers, David Ayala for kindly sharing the proof of Theorem A.1, and Alexander Kupers for generously answering my question on MathOverflow [sku]. I am also indebted to Daisuke Kishimoto and Mitsunobu Tsutaya for their constant support and encouragement. Last, but not least, I express my sincerest gratitude to the referee for carefully reading the manuscript and making valuable feedback, which has significantly improved this paper.
Notation and Terminology
  • The term $\infty$-category” will be used as a synonym of quasi-category of Joyal [Joy02]. We will mainly follow [Lur09] for notation and terminology for $\infty$-categories, with the following exceptions:
    • We say that a morphism $f:S\to T$ of simplicial sets is final if it is cofinal in the sense of [Lur09, Definition 4.1.1.1 ]. We say that $f$ is initial if its opposite is final.
    • We say that a functor $f:\mathcal{C}\to\mathcal{D}$ of $\infty$-categories is a localization with respect to a set $S$ of morphisms of $\mathcal{C}$ if it is initial among functors inverting the morphisms in $S$. (See Subsection A.1 for a precise definition.) This definition is more general than the definition in [Lur09].
    • Except in Appendix A, we will not distinguish between ordinary categories and their nerves.
  • If $\mathcal{C}$ is an $\infty$-category, we denote its maximal sub Kan complex by $\mathcal{C}^{\simeq}$ and refer to it as the core of $\mathcal{C}$.
  • Let $n\geq 0$. An $n$-manifold, or a manifold of dimension $n$, will always mean a topological manifold without boundary, i.e., a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$. If $n\geq 1$, we define an $n$-manifold with boundary to be a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$ or $0,\infty)\times\mathbb{R}^{n-1}$. In particular, not every $n$-manifold with boundary is an $n$-manifold.
  • Let $n\geq 0$. Given $n$-manifolds $M$ and $N$, we let $\operatorname{Emb}\left(M,N\right)$ denote the topological space of embeddings $M\to N$, topologized by the compact-open topology. We let $\mathcal{M}\mathsf{fld}_{n}^{T}$ denote the topological category whose objects are $n$-manifolds and whose hom spaces are given by $\operatorname{Emb}\left(-,-\right)$, and let $\mathcal{M}\mathsf{fld}_{n}^{\Delta}$ denote the simplicial category obtained from $\mathcal{M}\mathsf{fld}_{n}^{T}$ by applying the singular complex functor to the hom spaces. We let $\mathcal{M}\mathsf{fld}_{n}$ denote the homotopy coherent nerve of the topological category $\mathcal{M}\mathsf{fld}_{n}^{T}$, and let $\mathsf{Mfld}_{n}$ denote the nerve of the the ordinary category obtained from $\mathcal{M}\mathsf{fld}_{n}^{T}$ by forgetting the topology of the mapping spaces. Equivalences of $\mathcal{M}\mathsf{fld}_{n}$ are called isotopy equivalences.
    We let $\mathcal{D}\mathsf{isk}_{n}\subset\mathcal{M}\mathsf{fld}_{n}$ and $\mathsf{Disk}_{n}\subset\mathsf{Mfld}_{n}$ denote the full subcategories spanned by the objects that are homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$. For each $k\geq 0$, we let $\mathcal{D}\mathsf{isk}_{n}^{\leq k}\subset\mathcal{D}\mathsf{isk}_{n}$ and $\mathsf{Disk}_{n}^{\leq k}\subset\mathsf{Disk}_{n}$ denote the full subcategories spanned by the objects with at most $k$ components.
  • Let $n\geq 0$. We let $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the homotopy coherent nerve of the topological category of smooth $n$-manifolds and smooth embeddings, whose mapping spaces $\operatorname{Emb}_{\mathrm{sm}}\left(-,-\right)$ are topologized by the weak topology (also called the compact-open $C^{\infty}$ topology) of [Hir76, Chapter 2]. We let $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}\subset\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the full subcategory spanned by the smooth manifolds diffeomorphic to a finite disjoint union of $\mathbb{R}^{n}$ (with the standard smooth structure). Equivalences of $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ will be called smooth isotopy equivalences. The $\infty$-categories $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, $\mathsf{Mfld}_{\mathrm{sm},n}$, $\mathsf{Disk}_{\mathrm{sm},n}$, and $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}$ are defined as in the previous point. (We will not define a PL version of these $\infty$-categories because the author is not aware of a good topology on PL embeddings whose paths correspond to PL isotopies.)
  • Let $n,k\geq 0$. If $M$ is an $n$-manifold, we will write $\operatorname{Disj}\left(M\right)\subset\operatorname{Open}\left(M\right)$ for the full subcategory spanned by the open sets homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$, and let $\operatorname{Disj}^{\leq k}\left(M\right)$ denote its full subposet spanned by the open sets with at most $k$ components. If $M$ is smooth, we let $\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\subset\operatorname{Disj}\left(M\right)$ denote the full subcategory spanned by the open sets diffeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$ (with the standard smooth structure), and write $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)=\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\cap\operatorname{Disj}^{\leq k}\left(M\right)$.
  • Let $n\geq 0$ and let $M$ be an $n$-manifold. For each finite set $S$, we let $\operatorname{Conf}\left(S,M\right)\subset M^{S}$ denote the subspace consisting of the injections $S\to M$. For each $k\geq 0$, we will write $\operatorname{Conf}\left(k,M\right)=\operatorname{Conf}\left(\{1,\dots,k\},M\right)$. (Thus $\operatorname{Conf}\left(0,M\right)$ is a point.) The $k$th symmetric group $\Sigma_{k}$ acts on $\operatorname{Conf}\left(k,M\right)$ by precomposition. We will write $B_{k}\left(M\right)=\operatorname{Conf}\left(k,M\right)/\Sigma_{k}$ for the orbit space of this action. The points of $B_{k}\left(M\right)$ will be identified with subsets of $M$ of cardinality $k$.
[…]
  • For each finite set $S$, we will write $\mathcal{P}\left(S\right)$ for the poset of subsets of $S$, and $\mathcal{P}_{0}\left(S\right)$ for the poset of nonempty subsets of $S$.
1. Polynomial Functors
In this section, we will show that polynomial approximations exist and that they are constructed by Kan extending along $\operatorname{Disj}^{\leq k}\left(M\right)\to\operatorname{Open}\left(M\right)$ (Theorem 1.3).
Warning 1.1.
To avoid writing $\mathrm{op}$” repeatedly, we will work with covariant functors in this section, so that polynomial functors are described in terms of cosheaves and left Kan extensions. However, in later section, we sometimes work with the dual theory of presheaves and right Kan extensions (such as Theorems 2.7 and 3.16). The choice will be dictated by the ease it takes to state and prove the relevant results.
Let us start by stating the main result of this section (Theorem 1.3).
Definition 1.2.
Let $k\geq 0$, let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with finite colimits. Let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be a functor.
  1. (1)
    Let $\mathcal{J}\subset\operatorname{Open}\left(M\right)$ be a subcategory, and let $G:\mathcal{J}\to\mathcal{C}$ be a functor. We say that $G$ is isotopy invariant if it carries each morphism in $\mathcal{J}$ which is an isotopy equivalence to an equivalence of $\mathcal{C}$. Isotopy invariant functors will be called isotopy functors. We will write $\operatorname{Fun}_{\mathrm{istp}}\left(\mathcal{J},\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{J},\mathcal{C}\right)$ for the full subcategory spanned by the isotopy functors.
  2. (2)
    We say that $F$ is exhaustive if for each increasing sequence $U_{0}\subset U_{1}\subset\cdots$ of open sets of $M$, the map
    $\operatorname{colim}_{i}F\left(U_{i}\right)\to F\left(\bigcup_{i\geq 0}U_{i}\right)$
    is an equivalence.
  3. (3)
    We say that $F$ is $k$-excisive, or polynomial of degree $\leq k$, if for each open set $U$ of $M$ and each pairwise disjoint closed subsets $A_{0},\dots,A_{k}$ of $U$, the map
    $\operatorname{colim}_{\emptyset\neq S\subset\{0,\dots,k\}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)\to F\left(U\right)$
    is an equivalence.
  4. (4)
    We let $\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\subset\operatorname{Fun}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ denote the full subcategory spanned by the $k$-excisive, exhaustive isotopy functors.
Here is the main result of this section.
Theorem 1.3.
Let $k\geq 0$, let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. The left Kan extension functor $\operatorname{Fun}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\to\operatorname{Fun}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ restricts to a categorical equivalence
$\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{\mathrm{istp},\,\mathrm{exh}}^{k}\left(M;\mathcal{C}\right).$
In particular, the inclusion
$\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\hookrightarrow\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$
admits a right adjoint, which carries each object $F\in\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ to a left Kan extension of $F|\operatorname{Disj}^{\leq k}\left(M\right)$.
We will deduce Theorem 1.3 from two results: The first one is the following lemma, which is a consequence of Ayala–Francis’s localization theorem (Theorem A.1). Recall that the theorem says that the functor $\operatorname{Disj}^{\leq k}\left(M\right)\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ is a localization.
Lemma 1.4.
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{C}$ be a functor which is a left Kan extension of $F|\operatorname{Disj}^{\leq k}\left(M\right)$. If $F|\operatorname{Disj}^{\leq k}\left(M\right)$ is an isotopy functor, so is $F$.
Proof.
Since localizations are final [Lur25, Tag 02N9], Theorem A.1 gives an equivalence
$F\left(U\right)\simeq\operatorname{colim}_{V\in\operatorname{Disj}^{\leq k}\left(U\right)}F\left(V\right)\xrightarrow{\simeq}\operatorname{colim}_{V\in\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}}F\left(V\right)$
natural in $U\in\operatorname{Open}\left(M\right)$. The right-hand side is manifestly isotopy invariant, and we are done. ∎
The second one concerns the identification of polynomial functors with functors enjoying certain gluing properties. The gluing properties are best expressed using cosheaves with respect to Grothendieck topologies on $\operatorname{Open}\left(M\right)$, which we now introduce.
Definition 1.5.
Let $\mathcal{C}$ be an $\infty$-category, and let $\mathcal{A}$ be another $\infty$-category equipped with a Grothendieck topology [Lur09, Definition 6.2.2.1]. We will say that a functor $F:\mathcal{A}\to\mathcal{C}$ is a cosheaf (with respect to the Grothendieck topology) if for each object $A\in\mathcal{A}$ and each covering sieve $\mathcal{A}_{/A}^{0}\subset\mathcal{A}_{/A}$ of $A$, the composite
$\left(\mathcal{A}_{/A}^{0}\right)^{\triangleright}\to\mathcal{A}\xrightarrow{F}\mathcal{C}$
is a colimit diagram.
Definition 1.6.
Let $M$ be a manifold, and let $k\geq 0$. A subset $\mathcal{U}\subset\operatorname{Open}\left(M\right)$ of $M$ is called a Weiss $k$-cover if for each subset $S\subset M$ of cardinality $\leq k$, there is an element of $\mathcal{U}$ which contains $S$. The Weiss $k$-topology on $M$ is the Grothendieck topology on $\operatorname{Open}\left(M\right)$ whose covering sieves of each open set $U\in\operatorname{Open}\left(M\right)$ are precisely the Weiss $k$-covers $\mathcal{U}$ of $U$ that are also sieves on $U$. If $\mathcal{C}$ is an $\infty$-category, a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is called a Weiss $k$-cosheaf if it is a cosheaf with respect to the Weiss $k$-topology. The Weiss topology on $M$ is the intersection of the Weiss $k$-topologies as $k$ ranges over all nonnegative integers. A Weiss cosheaf is a cosheaf with respect to the Weiss topology.
Example 1.7.
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category.
  1. (1)
    Any nonempty collection of open sets of $M$ is a Weiss 0-cover of $X$. Consequently, a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is a Weiss 0-cosheaf if and only if it is essentially constant, i.e., it factors through a contractible Kan complex. Since $\operatorname{Open}\left(M\right)$ is already weakly contractible, this is equivalent to the requirement that $F$ carries all morphisms to equivalences.
  2. (2)
    A functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ is a Weiss $1$-cosheaf if and only if for each nonempty sieve $\mathcal{U}\subset\operatorname{Open}\left(M\right)$, the map
    $\operatorname{colim}_{U\in\mathcal{U}}F\left(U\right)\to F\left(\bigcup_{U\in\mathcal{U}}U\right)$
    is an equivalence. Thus a Weiss $1$-cosheaf is a cosheaf on $X$ (that is, a cosheaf with respect to the standard Grothendieck topology on $\operatorname{Open}\left(M\right)$) if and only if $F\left(\emptyset\right)$ is an initial object.
Definition 1.8.
Let $k\geq 0$, let $M$ be a manifold, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ be a functor of small $\infty$-categories. For each finite set $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $S\subset\chi\left(I\right)$.
  1. (1)
    We say that $\chi$ satisfies the Weiss $k$-condition if for each subset $S\subset M$ of cardinality at most $k$, the $\infty$-category $\mathcal{I}_{S}$ is weakly contractible.
  2. (2)
    We say that $\chi$ satisfies the Weiss condition if for every finite subset $S\subset M$, the $\infty$-category $\mathcal{I}_{S}$ is weakly contractible.
We now arrive at the identification result of polynomial functors with cosheaves and Kan extensions, which we prove at the end of this section.
Theorem 1.9.
Let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ is a left Kan extension of $F|\operatorname{Disj}^{\leq k}\left(M\right)$.
  2. (2)
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(U\right)$ be a functor satisfying the Weiss $k$-condition. Then the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(U\right)$
    is an equivalence of $\mathcal{C}$.
  3. (3)
    The functor $F$ is a Weiss $k$-cosheaf.
  4. (4)
    The functor $F$ is $k$-excisive and exhaustive.
For later discussions, we will also prove the following limit case of Theorem 1.9:
Theorem 1.10.
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be an isotopy functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ is a left Kan extension of $F|\operatorname{Disj}\left(M\right)$.
  2. (2)
    Let $\mathcal{I}$ be a small $\infty$-category, let $U\subset M$ be an open set, and let $\chi:\mathcal{I}\to\operatorname{Open}\left(U\right)$ be a functor which satisfies the Weiss condition. Then the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(U\right)$
    is an equivalence of $\mathcal{C}$.
  3. (3)
    The functor $F$ is a Weiss cosheaf.
Using Lemma 1.4 and Theorem 1.9, we can prove Theorem 1.3 as follows:
Proof of Theorem 1.3.
By Lemma 1.4, the left Kan extension functor restricts to a left adjoint
$L:\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right),$
which is fully faithful by [Lur09, Proposition 4.3.2.15]. By Theorem 1.9, the essential image of $L$ is $\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$. The claim follows. ∎
We now turn to the proofs of Theorems 1.9 and 1.10. We need a few lemmas.
The first lemma gives a sufficient condition for a functor into $\mathcal{D}\mathsf{isk}_{n/M}$ (or $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ to be final.
Lemma 1.11.
Let $n,k\geq 0$, let $\mathcal{I}$ be a (small) category and let $\overline{f}:\mathcal{I}^{\triangleright}\to\mathcal{M}\mathsf{fld}_{n}$ be a functor. Set $\overline{f}\left(I\right)=U_{I}$ for $I\in\mathcal{I}$ and $\overline{f}\left(\infty\right)=M$. For each subset $S\subset M$, let $\mathcal{I}_{S}\subset\mathcal{I}$ denote the full subcategory spanned by the objects $I\in\mathcal{I}$ such that $U_{I}$ contains $S$.
  1. (1)
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$, then the functor $\mathcal{I}\to\mathcal{D}\mathsf{isk}_{n}{}_{/M}$ is final.
  2. (2)
    Suppose that $U_{I}\in\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ for every $I\in\mathcal{I}$. If $\mathcal{I}_{S}$ is weakly contractible for every finite set $S\subset M$ of cardinality $\leq k$, then the functor $\mathcal{I}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ is final.
Proof.
We prove part (1); the proof of part (2) is similar. By Proposition B.3 and [Lur09, Theorem 4.2.4.1], it will suffice to show that, for each object $V\in\mathcal{D}\mathsf{isk}_{n}$, the map
$\operatorname{hocolim}_{I\in\mathcal{I}}\operatorname{Sing}\operatorname{Emb}\left(V,U_{I}\right)\to\operatorname{Sing}\operatorname{Emb}\left(V,M\right)$
is a weak homotopy equivalence of simplicial sets. If $V=\emptyset$, the claim is obvious because $\mathcal{I}=\mathcal{I}_{\emptyset}$ is weakly contractible by hypothesis. So suppose that $V$ is nonempty. Let $p$ denote the cardinality of the set of components of $V$, and fix a homeomorphism $V\cong\mathbb{R}^{n}\times\{1,\dots,p\}$. By Proposition C.9, the evaluation at the origin of $\mathbb{R}^{n}$ determines a homotopy cartesian square
$\operatorname{Sing}\operatorname{Emb}(V,U_{I})$$\operatorname{Sing}\operatorname{Emb}(V,M)$$\operatorname{Sing}\operatorname{Conf}(p,U_{I})$$\operatorname{Sing}\operatorname{Conf}(p,M)$}}
for every $I\in\mathcal{I}$. Since colimits in $\mathcal{S}$ are universal, we are reduced to showing that the map
$\operatorname{hocolim}_{I\in\mathcal{I}}\operatorname{Sing}\operatorname{Conf}\left(p,U_{I}\right)\to\operatorname{Sing}\operatorname{Conf}\left(p,M\right)$
is a weak homotopy equivalence. This is a direct consequence of our hypothesis and [Lur17, Theorem A.3.1]. ∎
Our next lemma is completely elementary in the topological case and the PL case, but we include a proof using Morse theory because we need it in the smooth version of Lemma 1.13.
Lemma 1.12.
Let $n\geq 1$, $k\geq 0$, and let $1<c_{0}<\cdots<c_{k}<1$ and $r>0$ be real numbers. Set
$C_{i}=\{x\in\mathbb{R}^{n}\mid\left(x_{1}-c_{i}\right)^{2}+\sum_{1<i\leq n}x_{i}^{2}\leq r^{2}\},$
and suppose that the sets $C_{0},\dots,C_{k}$ are mutually disjoint and are contained in the interior of $D^{n}$. Then $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)$ is obtained from $\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching an $\left(n-1\right)$-handle $k$ times. (See Figure 1.1.)
Figure $1.$ Picture of Lemma 1.12.
Proof.
Using bump functions, construct a smooth function $\phi:\mathbb{R}\to\mathbb{R}$ with the following properties (Figure 1.2):
  1. (1)
    For each $0\leq i\leq k$, the restriction $\phi|[c_{i}-r,c_{i}+r]$ is given by $\phi\left(x\right)=-\left(x-c_{i}\right)^{2}$.
  2. (2)
    The derivative of $\phi$ is positive on $\left(-\infty,c_{1}\right)$ and is negative on $\left(c_{k},\infty\right)$.
  3. (3)
    The function $\phi$ agrees with $x^{2}$ outside [1,1][-1,1]
  4. (4)
    For each $1\leq i\leq k$, there is a unique point $a_{i}\in\left(c_{i-1},c_{i}\right)$ such that $\phi^{\prime}\left(a_{i}\right)=0$.
  5. (5)
    For each $0<i\leq k$, we have $\phi^{\prime\prime}\left(a_{i}\right)>0$.
  6. (6)
    We have $1<\phi\left(a_{1}\right)<\cdots<\phi\left(a_{k}\right)$.
Figure 1.2. Graph of $\phi$.
We then define $F:\mathbb{R}^{n}\to\mathbb{R}$ by $F\left(x\right)=\phi\left(x_{1}\right)-\sum_{1<i\leq n}x_{i}^{2}$. By construction, the critical points of $F$ in $F^{-1}\left((-\infty,r^{2}]\right)$ are the points $\{\left(a_{i},0,\dots,0\right)\}_{1\leq i\leq k}$, and all of them are nondegenerate and have index n1n-1. Therefore, Morse theory (see, e.g., [Kos93, Proposition VII.2.2]) shows that the set $\mathbb{R}^{n}\setminus\operatorname{Int}\left(C_{0}\cup\cdots\cup C_{k}\right)=F^{-1}\left((-\infty,-r^{2}]\right)$ is obtained from $F^{-1}\left((-\infty,-1]\right)=\{x\in\mathbb{R}^{n}\mid\left\|x\right\|\geq 1\}$ by attaching $\left(n-1\right)$-handles $k$ times. ∎
[…]
Lemma 1.13.
Let $n,k\geq 0$, and let $M$ be an $n$-manifold. Let $\mathcal{U}$ be a set of open sets of $M$ which satisfies the following conditions:
  1. (1)
    The set $\mathcal{U}$ contains $\operatorname{Disj}^{\leq k}\left(M\right)$.
  2. (2)
    Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of elements in $\mathcal{U}$. Then $\bigcup_{i\geq 0}U_{i}$ belongs to $\mathcal{U}$.
  3. (3)
    Let $U\subset M$ be an open set and let $A_{0},\dots,A_{k}$ be pairwise disjoint closed sets of $U$. Suppose that, for each nonempty subset $S\subset\{0,\dots,k\}$, the open set $U\setminus\bigcup_{i\in S}A_{i}$ belongs to $\mathcal{U}$. Then $U$ belongs to $\mathcal{U}$.
Then $\mathcal{U}=\operatorname{Open}\left(M\right)$.
Proof.
The proof proceeds in several steps.
  1. (Step1)
    We show that $\mathcal{U}$ contains $\operatorname{Disj}\left(M\right)$. It suffices to show that $\operatorname{Disj}^{\leq p}\left(M\right)\subset\mathcal{U}$ for every $p\geq 0$. We prove this by induction on $p$. If $p\leq k$, the claim follows from our assumption (1). Suppose we have proved that $\operatorname{Disj}^{\leq p-1}\left(M\right)\subset\mathcal{U}$ for some $p>k$, and let $U\in\operatorname{Disj}^{\leq p}\left(M\right)$. We wish to show that $U\in\mathcal{U}$. Let $A_{0},\dots,A_{k}$ be distinct components of $U$. (There may be other components, but choose $k+1$ one of them.) For every nonempty subset $S\subset\{0,\dots,k\}$, the open set $U\setminus\bigcup_{i\in S}A_{i}$ belongs to $\operatorname{Disj}^{\leq p-1}\left(M\right)$ and hence to $\mathcal{U}$. It follows from the induction hypothesis and condition (3) that $U\in\mathcal{U}$, completing the induction.
  2. (Step2)
    Let $U\subset M$ be an open set which has the form $U=\operatorname{Int}N$ for some compact manifold $N$ with boundary admitting a handle decomposition. We will show that $U$ belongs to $\mathcal{U}$.
    Let $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ be an $\left(n+1\right)$-tuple of nonnegative integers. We say that a handle decomposition $\emptyset=N_{0}\subset\cdots\subset N_{\sum_{i=0}^{n}a_{i}}=N$ of $N$ is of type $\left(a_{0},\dots,a_{n}\right)$ if for each $0\leq d\leq n$, there are exactly $a_{d}$ integers $i$ such that $N_{i-1}$ is obtained from $N_{i}$ by attaching a $d$-handle. We define the handle type of $N$ to be the minimal element $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ for which $P$ admits a handle decomposition of type $\left(a_{0},\dots,a_{n}\right)$, where $\mathbb{Z}_{\geq 0}^{n+1}$ is endowed with the lexicographic ordering read from right to left. (In other words, given distinct elements $\left(a_{0},\dots,a_{n}\right),\left(b_{0},\dots,b_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$, we declare that $\left(b_{0},\dots,b_{n}\right)<\left(a_{0},\dots,a_{n}\right)$ if and only if $b_{i}<a_{i}$, where $i$ is the maximal integer such that $a_{i}\neq b_{i}$.) We will show by transfinite induction on the handle type of $N$ that $U$ belongs to $\mathcal{U}$.
    Let $\left(a_{0},\dots,a_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ be the handle type of $N$. Suppose that the claim has been proved for every element $\left(a^{\prime}_{0},\dots,a^{\prime}_{n}\right)\in\mathbb{Z}_{\geq 0}^{n+1}$ smaller than $\left(a_{0},\dots,a_{n}\right)$. We must show that $U$ belongs to $\mathcal{U}$. If $a_{1}=\cdots=a_{n}=0$, then $U$ belongs to $\operatorname{Disj}\left(M\right)$, and the claim follows from Step 1. So suppose that $a_{i}>0$ for some $i$. Let
    $\emptyset=N_{0}\subset\cdots\subset N_{a}=N$
    be a handle decomposition of $N$ of type $\left(a_{0},\dots,a_{n}\right)$, where we wrote $a=\sum_{i=0}^{n}a_{i}$. Since attaching a 0-handle is equivalent to adding a disjoint copy of $D^{n}$, by rearranging the order of the handle attachment, we may assume that $N_{a}$ is obtained from $N_{a-1}$ by attaching a handle of positive index, say $\lambda$. Let $e:D^{\lambda}\times D^{n-\lambda}\to N_{a}$ denote the corresponding embedding. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{\lambda}$ and set $A_{j}=C_{j}\times D^{n-\lambda}$. If $S\subset\{0,\dots,k\}$ is a nonempty subset, then we have
    $U\setminus\bigcup_{j\in S}A_{j}$ $=\operatorname{Int}\left(N_{a}\setminus\bigcup_{j\in S}e\left(\operatorname{Int}\left(C_{j}\right)\times D^{n-\lambda}\right)\right)$
    $=\operatorname{Int}\left(N_{a-1}\amalg_{S^{\lambda-1}\times D^{n-\lambda}}e\left(\left(D^{\lambda}\setminus\bigcup_{j\in S}\operatorname{Int}\left(C_{j}\right)\right)\times D^{n-\lambda}\right)\right).$
    Now according to Lemma 1.12, the manifold with boundary $N_{a}\left(S\right)=N_{a-1}\amalg_{S^{\lambda-1}\times D^{n-\lambda}}e\left(\left(D^{\lambda}\setminus\bigcup_{j\in S}\operatorname{Int}\left(C_{j}\right)\right)\times D^{n-\lambda}\right)$ admits a handle decomposition of type $\left(b_{0},\dots,b_{n}\right)$, where
    $b_{i}=\begin{cases}a_{i}&\text{if }i\neq\lambda,\lambda-1,\\ a_{\lambda-1}+\left|S\right|-1&\text{if }i=\lambda-1,\\ a_{\lambda}-1&\text{if }i=\lambda.\end{cases}$
    Since $\left(b_{0},\dots,b_{n}\right)<\left(a_{0},\dots,a_{n}\right)$, the induction hypothesis implies that $\operatorname{Int}\left(N_{a}\left(S\right)\right)$ belongs to $\mathcal{U}$. It follows from condition (3) that $U$ belongs to $\mathcal{U}$, as desired.
  3. (Step3)
    We show that if $U$ is smoothable or $n\neq 4$, then $U$ belongs to $\mathcal{U}$. By Step 2 and assumption (2), it suffices to show that $U$ has a (possibly infinite) handle decomposition. If $U$ is smoothable, then it has a handle decomposition by Morse theory [Wal16, Lemma 5.1.8]. If $n=5$, this is [Qui82, Theorem 2.3.1], and if $n\geq 6$, this is [KS77, Essay III, Theorem 2.1]. If $n\leq 3$, then it suffices to show that $U$ has a smooth structure. If $n=1$, this follows from the classification of $1$-manifolds [Lee11, Theorem 5.27]. If $n=2$, this is [Hat22, Theorem A]. If $n=3$, then by [Thu97, Theorem 3.10.8], it suffices to show that $U$ has a PL structure. This is [Ham76, Theorem 2].
  4. (Step4)
    We finish off the proof by considering the case where $n=4$. Let $U\subset M$ be a nonempty finite set. For each path component $V\subset U$, choose $\left(k+1\right)$ points $p_{0}^{V},\dots,p_{k}^{V}\in V$, and set $A_{j}=\{p_{j}^{V}\mid V\in\pi_{0}\left(U\right)\}$ for $0\leq j\leq k$. Then $A_{0},\dots,A_{k}$ are disjoint closed subsets of $U$, and moreover for each nonempty finite set $S\subset\{0,\dots,k\}$, the components of the open set $U\setminus\bigcup_{i\in S}A_{i}$ are noncompact. Since connected, non-compact $4$-manifolds are smoothable [FQ90, 8.2], we deduce from Step 2 that $\bigcap_{i\in S}U\setminus A_{i}$ belongs to $\mathcal{U}$. Condition (3) now shows that $U$ belongs to $\mathcal{U}$, and the proof is complete.∎
Remark 1.14.
Our proof of Lemma 1.13 is a modification of Weiss’s proof of [Wei99, Theorem 5.1], where he essentially established the lemma in the case where $M$ is smoothable. The topological case is a bit harder because we do not have a control of the order of the indices of handles.
Remark 1.15.
We have stated Lemma 1.13 to highlight its relation to exhaustivity and excisivity of functors. However, a closer inspection of the proof actually shows that we can replace condition (2) as follows:
  • (2)
    Let $U_{0}\subset U_{1}\subset\cdots$ be an increasing sequence of elements in $\mathcal{U}$, where each $U_{i}$ is has compact closure in $U$. Then $\bigcup_{i\geq 0}U_{i}$ belongs to $\mathcal{U}$.
If $M$ is a PL manifold and $\operatorname{Disj}\left(M\right)\subset\mathcal{U}$, then we can also replace condition (3) by a similar condition in which the $A_{i}$’s are compact subcomplexes of $U$ for some triangulation of $U$ coming from the PL structure, by using a handle decomposition associated to the triangulation. (See [RS72, 6.9].)
We now arrive at the proof of Theorems 1.9 and 1.10.
Proof of Theorem 1.9.
[…]
$\theta:\operatorname{colim}_{I\in\mathcal{I}^{\prime}}F\left(U_{I}\right)\to F\left(U\right)$
is an equivalence of $\mathcal{C}$. According to [Lur25, Tag 02MD] and [Lur25, Tag 02N9], there is a small category $\mathcal{I}$ and a final functor $\mathcal{I}\to\mathcal{I}^{\prime}$ whose pullbacks are all final. Then the composite $\chi\circ p$ satisfies the Weiss $k$-condition, so it suffices to consider the case where $\mathcal{I}^{\prime}=\mathcal{I}$.
Set $n=\dim M$. Using Theorem A.1 (and the fact that localizations are final), we can identify $\theta$ with the map
$\operatorname{colim}_{\left(I,V\right)\in\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}}F\left(V\right)\to\operatorname{colim}_{V\in\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}}F\left(V\right)$
induced by the functor $\phi:\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$. It will therefore suffice to show that $\phi$ is final. For this, we observe that the map $\psi:\int^{I\in\mathcal{I}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)\to\int^{I\in\mathcal{I}}\mathcal{D}\mathsf{isk}_{n/U_{I}}^{\leq k}$ is final (Proposition B.7). Since final maps have the right cancellation property [Cis19, Corollary 4.1.9], it suffices to show that the composite $\phi\circ\psi$ is final. We prove this by using Lemma 1.11: We must show that, for each finite set $S\subset U$, the category $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}$ is final. For this, observe that the projection $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}\to\mathcal{I}_{S}$ is a cocartesian fibration whose fibers are weakly contractible (as they are cofiltered), and that its base is also weakly contractible by hypothesis. It follows from Quillen’s theorem B (Proposition A.3) that $\int^{I\in\mathcal{I}_{S}}\operatorname{Disj}^{\leq k}\left(U_{I}\right)_{S}$ is also weakly contractible, as claimed.
[…]
Proof of Theorem 1.10.
[…]
Call an open set $U\subset M$ good if the map $\alpha_{U}$ is an equivalence. We wish to show that every open set of $M$ is good. We prove this in two steps.
Let $U\subset M$ be an open set homeomorphic to an open subset of $\mathbb{R}^{n}$, where $n=\dim M$. We claim that $U$ is good. Choose an embedding $U\hookrightarrow\mathbb{R}^{n}$ and identify $U$ with a subset of $\mathbb{R}^{n}$. Let $\operatorname{Cube}\left(U\right)$ denote the poset of open subsets of $U$ that are finite disjoint unions of open cubes lying in $U$, and let $\overline{\operatorname{Cube}\left(U\right)}$ denote the poset of open subsets of elements of $\operatorname{Cube}\left(U\right)$. Since $\operatorname{Cube}\left(U\right)$ is closed under finite intersection, the inclusion $\operatorname{Cube}\left(U\right)\to\overline{\operatorname{Cube}\left(U\right)}$ is final. Moreover, $\overline{\operatorname{Cube}\left(U\right)}$ is a covering sieve of $U$ in the Weiss topology on $\operatorname{Open}\left(M\right)$. Since every element of $\operatorname{Cube}\left(U\right)$ is good, it follows that $\alpha_{U}$ is an equivalence.
Next, let $U\subset M$ be an arbitrary subset. We claim that $U$ is good. Let $\mathcal{U}$ denote the set of all open sets of $U$ that are homeomorphic to an open subset of $\mathbb{R}^{n}$. By the result in the previous paragraph, every element in $\mathcal{U}$ is good. Moreover, $\mathcal{U}$ is a covering sieve of $U$ in the Weiss topology. It follows that $\alpha_{U}$ is an equivalence, and the proof is complete. ∎
Remark 1.16.
Using Variant A.10, we can show that everything in this section is valid in the smooth case if we make the following replacements:
  • manifolds by smooth manifolds;
  • isotopy equivalences by smooth isotopy equivalences;
  • $\operatorname{Disj}\left(-\right)$ and $\operatorname{Disj}^{\leq k}\left(-\right)$ by $\operatorname{Disj}_{\mathrm{sm}}\left(-\right)$ and $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(-\right)$;
A similar remark applies to the PL case.
Remark 1.17.
Let $M$ be a manifold, and let $\mathcal{C}$ be a locally small $\infty$-category with small colimits. Isotopy invariance, exhaustivity, and $k$-excisivity of a functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ can be detected jointly by the hom-functors (i.e., by the composites $\operatorname{Open}\left(M\right)^{\mathrm{op}}\xrightarrow{F}\mathcal{C}^{\mathrm{op}}\xrightarrow{\mathcal{C}\left(-,C\right)}\mathcal{S}$, where $C$ ranges over the objects in $\mathcal{C}$). Therefore, some of the key results of this section (like Lemma 1.4) could have been proved by outsourcing everything to the case of space-valued presheaves, covered in [Wei99]. We have opted for the current approach because we believe that it is more conceptual, and also because it clarifies the connection between polynomials and Weiss sheaves.
2. Taylor Towers and Their Convergence
In the previous section, we saw that polynomial approximations exist, and that they are constructed by Kan extensions. In this section, we use it to construct the Taylor (co)tower of an isotopy functor and discuss when the tower converges to the original functor.
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. Theorem 1.3 says that for each isotopy functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ and each $k\geq 0$, there is a best approximation of $F$ by $k$-excisive, exhaustive isotopy functor, namely, the left Kan extension $T_{k}F$ of $F|\operatorname{Disj}^{\leq k}\left(M\right)$. This is called the $k$th polynomial approximation of $F$. We now organize these approximations into a cotower
$T_{0}F\to T_{1}F\to\cdots\to F.$
For later discussions, we work in a slightly more general setting.
Definition 2.1.
Let $\mathcal{A}$ be an $\infty$-category equipped with a sequence $\mathcal{A}^{0}\subset\mathcal{A}^{1}\subset\cdots$ of full subcategories. Set $\mathcal{A}^{\infty}=\bigcup_{i}\mathcal{A}^{i}$. We let $\mathbb{T}\left(\mathcal{A}\right)\subset\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects $\left(i,A\right)$, where $i\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}_{i}$, together with the objects $\left(\infty,A\right)$, where $A\in\mathcal{A}$.
Let $\mathcal{C}$ be another $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. A functor $\mathbb{T}\left(F\right):\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}\to\mathcal{C}$ is called a Taylor cotower of $F$ (with respect to the subcategories $\{\mathcal{A}^{i}\}_{i\geq 0}$) if it is a left Kan extension of the composite
$\mathbb{T}\left(\mathcal{A}\right)\xrightarrow{\text{projection}}\mathcal{A}\xrightarrow{F}\mathcal{C}.$
For each $k\in\mathbb{Z}_{\geq 0}\star\{\infty\}$, we set $T_{k}F=\mathbb{T}\left(F\right)|\{k\}\times\mathcal{A}$. Note that for each $k\in\mathbb{Z}_{\geq 0}$ and $A\in\mathcal{A}$, the inclusion
$\{\operatorname{id}_{k}\}\times\mathcal{A}_{/A}^{k}\hookrightarrow\mathbb{T}\left(\mathcal{A}\right)_{/\left(k,A\right)}$
is final (for it is a right adjoint), so the restriction $T_{k}F$ is a left Kan extension of $F|\mathcal{A}^{k}$. By convention, we will write $T_{-1}F$ for any functor which carries all objects of $\mathcal{A}$ to an initial object of $\mathcal{C}$.
A Taylor cotower $\mathbb{T}\left(F\right)$ of $F$ is said to be convergent if the functor $\mathbb{Z}_{\geq 0}\star\{\infty\}\to\operatorname{Fun}\left(\mathcal{A},\mathcal{C}\right)$ adjoint to $\mathbb{T}\left(F\right)$ is a colimit diagram.
Example 2.2.
Let $M$ be a manifold, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be a functor. A Taylor cotower of $F$ is a Taylor cotower of $F$ with respect to the subcategories $\{\operatorname{Disj}^{\leq i}\left(M\right)\}_{i\geq 0}$. If $F$ is an isotopy functor, then for each $k\in\mathbb{Z}$, the $k$th level $T_{k}F$ of a Taylor cotower of $F$ is polynomial of degree $\leq k$ (Theorem 1.3).
Warning 2.3.
Despite the name, the constituents of Taylor cotowers of non-isotopy functors may fail to be polynomial. For example, consider the functor $F:\operatorname{Open}\left(\mathbb{R}\right)\to\mathsf{Set}$ defined by $F\left(U\right)=U\times U$. It is easy to check that $T_{1}F$ is not excisive.
The proposition below gives the characterization of convergent functors.
Proposition 2.4.
Let $\mathcal{A}$ be an $\infty$-category, and let $\mathcal{A}_{0}\subset\mathcal{A}_{1}\subset\cdots$ be a sequence of full subcategories of $\mathcal{A}$. Set $\mathcal{A}_{\infty}=\bigcup_{i\geq 0}\mathcal{A}_{i}$. Let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{A}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    The functor $F$ has a convergent Taylor cotower.
  2. (2)
    The functor $F$ is a left Kan extension of $F|\mathcal{A}_{\infty}$.
Proof.
We will use the transitivity of Kan extensions [Lur09, Proposition 4.3.2.8]. Let $T=\mathbb{T}\left(F\right)$ be a Taylor cotower of $F$. Since colimits in functor categories can be formed pointwise [Lur25, Tag 02X9], an easy finality argument shows that condition (1) is equivalent to the following condition:
  • (1)
    $T$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$.
Let $\mathcal{X}\subset\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}$ denote the full subcategory spanned by the objects in $\mathbb{Z}_{\geq 0}\times\mathcal{A}$ and the objects in $\{\infty\}\times\mathcal{A}_{\infty}$. Since every object $A\in\mathcal{A}_{\infty}$ belongs to some $\mathcal{A}_{i}$, the functor $T|\mathcal{X}$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$. Therefore, condition (1) is equivalent to the following condition:
  • (1′′)
    $T$ is a left Kan extension of $T|\mathcal{X}$.
On the other hand, $T|\mathcal{X}$ is a left Kan extension of $T|\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}_{\infty}$, so (1′′) is equivalent to the following condition:
  • (1′′′)
    $T$ is a left Kan extension of $T|\left(\mathbb{Z}_{\geq 0}\star\{\infty\}\right)\times\mathcal{A}_{\infty}$.
Since $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}$ is a left Kan extension of $T|\mathbb{Z}_{\geq 0}\times\mathcal{A}_{\infty}$, condition (1′′′) is equivalent to condition (2), and we are done. ∎
Corollary 2.5.
Let $M$ be a manifold, and let $\mathcal{C}$ be an $\infty$-category with small colimits. An isotopy functor $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ has a convergent Taylor cotower if and only if $F$ satisfies the equivalent conditions of Theorem 1.10.
Proof.
This follows from Proposition 2.4 and Theorem 1.10. ∎
When we can talk about connectivity of morphisms, we can give a better convergence result. We will focus on one such instance, where the target $\infty$-category is a stable $\infty$-category with t-structures. (See Section B.3 for a brief review of t-structures.) The following definition follows [GW99, Theorem 2.3].
Definition 2.6.
Let $M$ be a smooth manifold, let $\mathcal{C}$ be a stable $\infty$-category with $t$-structures, let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an exhaustive isotopy functor, and let $\rho$ and $c$ be integers. We say that $G$ is $\rho$-analytic with excess $c$ if it satisfies the following condition:
  • ($\ast$)
    Let $P\subset M$ be a smooth compact submanifold with boundary, and let $Q_{0},\dots,Q_{r}\subset M\setminus\operatorname{Int}\left(P\right)$ be pairwise disjoint, compact, codimension 0 smooth submanifolds with boundary. Suppose that each $Q_{i}$ has handle index $q_{i}<\rho$ (i.e., it is obtained from $\partial Q_{i}$ by attaching handles of indices $\leq q_{i}$). Then the $\left(r+1\right)$-cube
    $\mathcal{P}\left(\{0,\dots,r\}\right)^{\mathrm{op}}\to\mathcal{C},\,S\mapsto F\left(\operatorname{Int}\left(P\cup\bigcup_{s\in S}Q_{s}\right)\right)$
    is $\left(c+\sum_{i=0}^{r}\left(\rho-q_{i}\right)\right)$-cartesian.
The following result should be compared with [GW99, Theorem 2.3]. It also generalizes [RW14, Theorem E.5].
Theorem 2.7.
Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $M$ be a smooth manifold, let $\rho$ and $c$ be integers, and let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an exhaustive isotopy functor. Suppose that the following conditions are satisfied:
  1. (1)
    $F$ is $\rho$-analytic with excess $c$.
  2. (2)
    $\mathcal{C}$ has small limits.
  3. (3)
    The functor $\pi_{0}:\mathcal{C}\to\mathcal{C}^{\heartsuit}$ preserves countable products.
Then for every $k\geq 1$ and every open set $W\subset M$ admitting a proper Morse function whose critical points have indices less than $q$, the map
$\eta_{k-1,W}:F\left(W\right)\to T_{k-1}F\left(W\right)$
is $\left(c+k\left(\rho-q\right)-1\right)$-connected. [^2] This is slightly weaker than the estimate of [GW99], where it says $c+k\left(\rho-q\right)$ instead of $c+k\left(\rho-q\right)-1$. However, a close inspection of the proof of loc. cit. shows that it only gives the same estimate as ours. Michael Weiss informed me that this issue can probably be fixed, but the details are not written down yet. We believe that his technique will be applicable to our setting too, so we should be able to say that $\eta_{k-1,W}$ is $c+k\left(\rho-q\right)$-connected. In particular, if $\rho>q$ and $\bigcap_{n}\mathcal{C}_{\geq n}=0$, then the map
$F\left(W\right)\to\operatorname{lim}_{k}T_{k}F\left(W\right)$
is an equivalence.
Proof.
We follow [GW99, Theorem 2.3]. Shifting degrees if necessary, we may assume that $c=0$. Suppose first that $\overline{W}$ is a compact smooth submanifold with boundary of $M$ and admits a handle decomposition with handles of indices at most $q$. We will show that $\eta_{k-1,W}$ is $k\left(\rho-q\right)$-connected by induction on $q$ and the number of $q$-handles.
If $q=0$, then $W$ is the disjoint union of open balls, say $W_{1},\dots,W_{l}$. If $l\leq k-1$, then $\eta_{k-1,W}$ is an equivalence, and we are done. If $l\geq k$, it suffices to show that for each $k\leq t\leq l$, the map
$\theta_{t}:T_{t}F\left(W\right)\to T_{t-1}F\left(W\right)$
is $t\rho$-connected. For this, for each subset $S\subset\pi_{0}\left(W\right)$, let $W_{S}\subset W$ denote the union of the elements in $S$. Consider the following commutative diagram
$\coprod_{S\subset\pi_{0}(W)}\mathcal{P}_{0}(S)$$\coprod_{S\subset\pi_{0}(W),\,|S|=t}\operatorname{Disj}^{\leq t-1}(W_{S})$$\operatorname{Disj}^{\leq t-1}(W)$$\coprod_{S\subset\pi_{0}(W)}\mathcal{P}(S)$$\coprod_{S\subset\pi_{0}(W),\,|S|=t}\operatorname{Disj}^{\leq t}(W_{S})$$\operatorname{Disj}^{\leq t}(W),$$\scriptstyle{\phi_{0}}$$\scriptstyle{\phi}$
where the map $\phi|\mathcal{P}\left(S\right)$ is given by $R\mapsto W_{R}$. The right-hand square is a pushout in $\mathcal{C}\mathsf{at}_{\infty}$ (because the Joyal model structure is left proper), and the maps $\phi_{0}$ and $\phi$ are final by inspection. It follows from Remark B.10 that $\theta_{t}$ is a pullback of the map
$\prod_{S\subset\pi_{0}\left(W\right)}G\left(W_{S}\right)\to\prod_{S\subset\pi_{0}\left(W\right)}\operatorname{lim}_{R\in\mathcal{P}_{0}\left(S\right)}G\left(W_{R}\right),$
which is $t\rho$-connected by the analyticity assumption (applied to $P=\emptyset$ and $Q_{i}=\overline{W_{i}}$). Hence $\theta_{t}$ is $t\rho$-connected, as required.
For the inductive step, suppose we can write $\overline{W}=N\cup H$, where $H\cong D^{q}\times D^{n-q}$ is a $q$-handle and $N$ has one less $q$-handle than $\overline{W}$. Choose disjoint closed disks $C_{0},\dots,C_{k}\subset\operatorname{Int}D^{q}$, and set $A_{i}=C_{i}\times D^{n-q}$. By the induction hypothesis, for each $S\in\mathcal{P}_{0}\left(\{0,\dots,k\}\right)$, the map
$G\left(W\setminus\bigcup_{i\in S}A_{i}\right)\to T_{k-1}\left(W\setminus\bigcup_{i\in S}A_{i}\right)$
is $k\left(\rho-q+1\right)$-connected. It follows from Corollary B.15 that the map
$\operatorname{lim}_{S\in\mathcal{P}_{0}\left(S\right)}G\left(W\setminus\bigcup_{i\in S}A_{i}\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(S\right)}T_{k-1}\left(W\setminus\bigcup_{i\in S}A_{i}\right)\simeq T_{k-1}\left(W\right)$
is $k\left(\rho-q\right)$-connected. Moreover, the cube $S\mapsto G\left(W\setminus\bigcup_{i\in S}A_{i}\right)$ is $k\left(\rho-q\right)$-cartesian by the analyticity assumption. Hence the map $G\left(W\right)\to T_{k-1}\left(W\right)$ is $k\left(\rho-q\right)$-connected (Proposition B.12), completing the induction.
The general case and the last claim are taken care of by the Milnor exact sequence (Lemma B.16; note that conditions (2) and (3) ensure that the hypotheses of the lemma are satisfied), and we are done. ∎
3. Classification of Homogeneous Functors
In this section, we classify homogeneous functors, the building blocks of Taylor towers in manifold calculus. More formally, let $k\geq 0$, let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $F:\operatorname{Open}\left(M\right)\to\mathcal{C}$ be a functor which is polynomial of degree $\leq k$. We say that $F$ is homogeneous of degree $k$ if $F$ evaluates to a zero object at each element $U\in\operatorname{Disj}\left(M\right)$ with fewer than $k$ components; or equivalently, if the $\left(k-1\right)$th level $T_{k-1}F$ of a Taylor cotower of $F$ (Example 2.2) carries every object to a zero object. For example, every functor of polynomial of degree $\leq 0$ is homogeneous of degree 0; a $1$-excisive exhaustive isotopy functor is homogeneous of degree $1$ if and only if it is a cosheaf (by Theorem 1.9 and Example 1.7).
We are interested in homogeneous functors for the following reason: Suppose that $F$ is an isotopy functor. A cofiber of $T_{k-1}F\to T_{k}F$, called a $k$th homogeneous layer of $F$, is then an isotopy functor which is homogeneous of degree $\leq k$: It is an isotopy functor by Lemma 1.4; it is polynomial of degree $\leq k$ by Theorem 1.9; homogeneity is clear. Therefore, if $F$ has a convergent Taylor cotower, then we can reduce the analysis of $F$ to those of the homogeneous layers of $F$, at least up to extension problems.
In this section, we prove that homogeneous functors are classified by unordered configuration spaces and “objects of sections trivial near the fat diagonal” (Corollary 3.5 and Theorem 3.16). To illustrate the utility of the classification, we also give a connectivity estimate of these approximations (Corollary 3.17).
Notation 3.1.
Let $M$ be a manifold, and let $\mathcal{C}$ be a pointed $\infty$-category with small colimits. We let $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ denote the full subcategory of $\operatorname{Fun}\left(\operatorname{Open}\left(M\right),\mathcal{C}\right)$ spanned by the exhaustive isotopy functors that are homogeneous of degree $k$.
Recall that if $M$ is an $n$-manifold, then $I_{M}^{=k}$ denotes the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between the elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components (Notation A.7). The poset $I_{M}^{=k}$ classifies homogeneous functors in the following sense:
Proposition 3.2.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $k\geq 0$. The inclusion $I_{M}^{=k}\subset\operatorname{Open}\left(M\right)$ induces a trivial fibration
$\theta:\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right).$
Proof.
By Theorem 1.3, the restriction functor
$p:\operatorname{Exc}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$
is a trivial fibration. Let $\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\subset\operatorname{Fun}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$ denote the full subcategory spanned by the functors which carry every object in $\operatorname{Disj}^{\leq k-1}\left(M\right)$ to a zero object. The functor
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\to\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)$
is a pullback of $p$, so it is a trivial fibration. On the other hand, Proposition B.19 shows that the functor
$\operatorname{Fun}^{\prime}_{\mathrm{istp}}\left(\operatorname{Disj}^{\leq k}\left(M\right),\mathcal{C}\right)\to\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)$
is a trivial fibration. Therefore, $\theta$ is a composition of trivial fibrations and hence is itself a trivial fibration. ∎
Now if $M$ is a manifold and $k\geq 0$, then a functor on $I_{M}^{=k}$ is an isotopy functor if and only if it maps every morphism of $I_{M}^{=k}$ to an equivalence. Therefore, Proposition 3.2 says that homogeneous functors of degree $k$ on a manifold $M$ can be classified in terms of the classifying space of $I_{M}^{=k}$, i.e., a localization of $I_{M}^{=k}$ with respect to all morphisms [Lur25, Tag 01MY]. The following proposition identifies the homotopy type of this classifying space.
Proposition 3.3.
Let $M$ be a manifold. For each $k\geq 0$, there is a weak homotopy equivalence
$I_{M}^{=k}\xrightarrow{\simeq}\operatorname{Sing}B_{k}\left(M\right)$
which carries an object $U\in I_{M}^{=k}$ to a point $\{p_{1},\dots,p_{k}\}\in B_{k}\left(M\right)$ which intersects every component of $U$.
Remark 3.4.
The identification of the homotopy type of $I_{M}^{=k}$ dates back at least to [Wei99, Lemma 3.5] and has been reproved again and again (e.g., [AF15, Lemma 2.12]). We include a proof for later reference.
Proof.
We consider the maps
$I_{M}^{=k}\xleftarrow{\pi}\int^{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)\xrightarrow{\phi}\operatorname{Sing}B_{k}\left(M\right),$
where $B^{\prime}_{k}\left(-\right)$ is defined in Notation A.7. The map $\pi$ is a trivial fibration because it is a left fibration with contractible fibers. So it has a section $\sigma$, and $\phi\circ\sigma$ carries each element $U\in I_{M}^{=k}$ to a point in $B^{\prime}_{k}\left(U\right)$. It will therefore suffice to show that $\phi$ is a weak homotopy equivalence. According to (the argument of) [Lur09, Corolary 3.3.4.6], it will suffice to show that the map
$\operatorname{hocolim}_{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)\to\operatorname{Sing}B_{k}\left(M\right)$
is a weak homotopy equivalence. This follows from [Lur17, Theorem A.3.1]. ∎
Combining Propositions 3.2 and 3.3, we obtain:
Corollary 3.5.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, and let $k\geq 0$. There is a zig-zag of categorical equivalences
$\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)\xleftarrow{\simeq}\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right).$
Remark 3.6.
[…]
$\overline{L}_{k}F\left(\{p_{1},\dots,p_{k}\}\right)$ $\simeq L_{k}F\left(U\right)$
$\simeq\mathrm{cofib}\left(T_{k-1}F\left(U\right)\to T_{k}F\left(U\right)\right).$
Since $T_{k-1}F$ is $\left(k-1\right)$-excisive, the map $T_{k-1}F\left(U\right)\to T_{k}F\left(U\right)$ can be identified with the map
$\theta:\operatorname{colim}_{S\subsetneq\{1,\dots,k\}}F\left(\bigcup_{i\in S}B_{i}\right)\to F\left(\bigcup_{i=1}^{k}B_{i}\right).$
A cofiber of $\theta$ is called a total cofiber of the $k$-dimensional cubical diagarm $S\mapsto F\left(\bigcup_{i\in S}B_{i}\right)$. Total cofibers admit an interpretation in terms of a derivative: For example if $k=2$, then the total homotopy cofiber can be written as
$\mathrm{cofib}\left(\mathrm{cofib}\left(F\left(\emptyset\right)\to F\left(B_{1}\right)\right)\to\operatorname{cofib}\left(F\left(B_{2}\right)\to F\left(B_{1}\cup B_{2}\right)\right)\right)$
which resembles the classical formula
$f^{\prime\prime}\left(0\right)=\operatorname{lim}_{h_{1},h_{2}\to 0}\frac{f\left(h_{1}+h_{2}\right)-f\left(h_{1}\right)-f\left(h_{2}\right)+f\left(0\right)}{h_{1}h_{2}}$
of the second derivative. Therefore, the values of $\overline{L}_{k}F$ are these “$k$th derivatives” of $F$, highlighting the connection with the Taylor expansion in ordinary calculus.
In the situation of Corollary 3.5, it is relatively easy to pass from $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$ to $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)$: We just have to restrict a homogenous functor to $I_{M}^{=k}$, and then pass to the localization. We now explain how to go in the reverse direction (Theorem 3.16).
The starting point is the observation that for a manifold $M$, the equivalence $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Fun}_{\mathrm{istp}}\left(I_{M}^{=k},\mathcal{C}\right)$ coming from Proposition 3.3 is given by
$F\mapsto\left(U\mapsto\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right).$
(This follows from the proof of Proposition 3.3. Also, we used the informal notation of Remark B.6, applied to the cocartesian fibration $\int^{U\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(U\right)$). A bit more consideration then leads to the following proposition:
Proposition 3.7.
Let $M$ be a manifold, let $\mathcal{C}$ be a pointed $\infty$-category with small colimits, let $k\geq 0$, and let $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ be a functor. Define another functor $\widetilde{F}:\operatorname{Open}\left(M\right)\to\mathcal{C}$ by
$\widetilde{F}\left(U\right)=\operatorname{colim}_{\operatorname{Sing}B_{k}\left(U\right)}F.$
Then:
  1. (1)
    $\widetilde{F}$ is $k$-excisive, exhaustive, and isotopy invariant.
  2. (2)
    The assignment $F\mapsto L_{k}\widetilde{F}$ determines an inverse equivalence
    $\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M;\mathcal{C}\right)$
    of the equivalence of Proposition 3.2.
We need a few lemmas for the proof of Proposition 3.7.
Notation 3.8.
Let $M$ be a manifold and let $k\geq 0$. We write $B^{\prime\prime}_{k}\left(M\right)$ for the complement of $B^{\prime}_{k}\left(M\right)$ in $B_{k}\left(M\right)$. That is, $B^{\prime\prime}_{k}\left(M\right)$ is the subspace of $B_{k}\left(M\right)$ consisting of the points $\{p_{1},\dots,p_{k}\}$ such that the map $\{p_{1},\dots,p_{k}\}\to\pi_{0}\left(M\right)$ is not injective.
Lemma 3.9.
Let $U$ be a manifold whose components are homeomorphic to $\mathbb{R}^{n}$, and let $k\geq 1$. The map
$\int^{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{Sing}B_{k}\left(V\right)\to\operatorname{Sing}\left(B^{\prime\prime}_{k}\left(U\right)\right)$
is a weak homotopy equivalence.
Proof.
By [Lur09, Lemma 3.3.4.1 and Proposition 3.3.4.2], it suffices to show that the diagram $\{\operatorname{Sing}B_{k}\left(V\right)\to\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)\}_{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}$ is a homotopy colimit diagram of simplicial sets. According to [Lur09, Theorem A.3.1], it suffices to show that, for each point $S\in B^{\prime\prime}_{k}\left(U\right)$, the full subcategory $\operatorname{Disj}^{\leq k-1}\left(U\right)_{S}\subset\operatorname{Disj}^{\leq k-1}\left(U\right)$ spanned by the objects $V\in\operatorname{Disj}^{\leq k-1}\left(U\right)$ such that $S\subset V$ is weakly contractible. Since $S$ lies in no more than $\left(k-1\right)$ components of $U$, such a $V$ necessarily lies in the union of the components of $U$ containing points in $S$. Thus $\operatorname{Disj}^{\leq k-1}\left(U\right)_{S}$ has a final object, and is in particular weakly contractible. ∎
Proof of Proposition 3.7.
The claim is trivial if $k=0$, so we will assume that $k\geq 1$ throughout.
We start with (1). By Proposition B.8, it suffices to show that the functor
$\operatorname{Sing}B_{k}\left(-\right):\operatorname{Open}\left(M\right)\to\mathcal{S}$
is $k$-excisive, exhaustive, and isotopy invariant. The first two follows from [Lur17, Theorem A.3.1], while isotopy invariance is obvious. [^4] Incidentally, $B_{k}\left(-\right)$ is one of the “classic” polynomial functors that appeared in the original paper of Weiss; see [Wei99, Example 2.4].
For (2), let $U\in I_{M}^{=k}$ and $F\in\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right)\right)$. Since $B_{k}\left(U\right)$ is the disjoint union of $B^{\prime}_{k}\left(U\right)$ and $B^{\prime\prime}_{k}\left(U\right)$, we have an equivalence
$\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right)\vee\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right)\xrightarrow{\simeq}F\left(U\right)$
natural in $F$ and $U$. (But beware that the summands on the right-hand side are not functors of $U\in\operatorname{Open}\left(M\right)$, because an inclusion of open sets of $M$ may fail to be injective on $\pi_{0}$.) Moreover, Lemma 3.9 give us equivalences
$T_{k-1}\widetilde{F}\left(U\right)$ $=\operatorname{colim}_{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{colim}_{\operatorname{Sing}B_{k}\left(V\right)}F$
$\simeq\operatorname{colim}_{\int^{V\in\operatorname{Disj}^{\leq k-1}\left(U\right)}\operatorname{Sing}B_{k}\left(V\right)}F\left(W\right)$
$\simeq\operatorname{colim}_{\operatorname{Sing}\left(B^{\prime\prime}_{k}\left(U\right)\right)}F.$
Under these equivalences, the map $T_{k-1}\widetilde{F}\left(U\right)\to\widetilde{F}\left(U\right)$ can be identified with the inclusion of the summand
$\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right)\to\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F\right)\vee\left(\operatorname{colim}_{\operatorname{Sing}B^{\prime\prime}_{k}\left(U\right)}F\right).$
Hence we obtain equivalences
$L_{k}\widetilde{F}\left(U\right)\simeq\mathrm{cofib}\left(T_{k-1}\widetilde{F}\left(U\right)\to\widetilde{F}\left(U\right)\right)\simeq\operatorname{colim}_{\operatorname{Sing}B^{\prime}_{k}\left(U\right)}F$
natural in $F$ and $U$. As we observed right before stating Proposition 3.7, the right-hand side computes the desired inverse, so we are done. ∎
Proposition 3.7 has a refinement when there is a good supply of open sets of $B_{k}\left(M\right)$ of configurations where two or more points are “close.” To explain this in more detail, we need a few preliminaries.
Warning 3.10.
Following Warning 1.1, we will focus on manifold calculus of presheaves for the remainder of this section. If $M$ is a manifold and $\mathcal{C}$ is a pointed $\infty$-category with small limits, we write $\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)\subset\operatorname{Fun}\left(\operatorname{Open}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$ for the full subcategory spanned by the functors that are isotopy invariant, exhaustive, and homogeneous of degree $k$ when viewed as a functor $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$.
Notation 3.11.
Let $\mathcal{C}$ be an $\infty$-category with small limits, let $X$ be a topological space, and let $F:\operatorname{Sing}X\to\mathcal{C}$ be a functor. For each subset $A\subset X$, we define $\Gamma\left(F;A\right)=\operatorname{lim}_{\operatorname{Sing}A}F.$
Remark 3.12.
Notation 3.11 is motivated by the following observation: In the situation of Notation 3.11, suppose that $\mathcal{C}=\mathcal{S}$. By the straightening–unstraightening equivalence, $F$ corresponds to a Kan fibration $p:E\to\operatorname{Sing}X$. The naturality of the straightening–unstraightening correspondence gives an equivalence between $\Gamma\left(F;A\right)$ and the Kan complex $\operatorname{Fun}_{\operatorname{Sing}X}\left(\operatorname{Sing}A,E\right)$ of sections of $p$ over $\operatorname{Sing}A$.
We can also interpret $\operatorname{lim}_{\operatorname{Sing}X}F$ as a “twisted cochain.” To see this, consider the functors
$\mathbf{\Delta}\xleftarrow{\pi}\mathbf{\Delta}_{/\operatorname{Sing}X}\xrightarrow{\mathrm{fin}}\operatorname{Sing}X,$
where $\mathrm{fin}$ denotes the final vertex functor (dual of [Lan21, Lemma 3.3.6]). The functor $\mathrm{fin}$ is initial [Lan21, Theorem 3.3.8], so we have equivalences
$\operatorname{lim}_{\operatorname{Sing}X}F$ $\simeq\operatorname{lim}_{\mathbf{\Delta}_{/\operatorname{Sing}X}}\left(F\circ\mathrm{fin}\right)$
$\simeq\operatorname{lim}_{\mathbf{\Delta}}\operatorname{Ran}_{\pi}\left(F\circ\mathrm{fin}\right)$
$\simeq\operatorname{lim}_{\mathbf{\Delta}}\left([n]\mapsto\prod_{\sigma:\Delta^{n}\to X\in\left(\operatorname{Sing}X\right)_{n}}F\left(\sigma\left(n\right)\right)\right).$
Here the last equivalence is an instance of Lemma B.5. When $\mathcal{C}$ is the $\infty$-category of cochain complexes of abelian groups, the limit of a cosimplicial cochain complex is computed as the totalization of the associated bicomplex [Ara23b, Remark 1.20, Example 2.4]. If further $F$ takes values in the category of abelian groups, the resulting complex is nothing but the singular complex of $X$ with local coefficients $F$ [Whi12, VI.2].
Definition 3.13.
Let $M$ be a manifold, and let $k\geq 0$. We define the $k$-fold fat diagonal of $M$ by
$\blacktriangle_{k}\left(M\right)=\operatorname{SP}^{k}\left(M\right)\setminus B_{k}\left(M\right),$
where $\operatorname{SP}^{k}\left(M\right)=X^{k}/\Sigma_{k}$ denotes the $k$th symmetric product of $X$. We write $\operatorname{Nbd}\left(\blacktriangle_{k}\left(M\right)\right)$ for the poset of neighborhoods of $\blacktriangle_{k}\left(M\right)$ in $\operatorname{SP}^{k}\left(M\right)$.
If $M$ is triangulated and its vertices are well-ordered, $\operatorname{SP}^{k}\left(M\right)$ has a preferred triangulation containing $\blacktriangle_{k}\left(M\right)$ as its subcomplex (Example C.12). In this situation, we write $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(M\right)\right)$ for the full subposet of $\operatorname{Nbd}\left(\blacktriangle_{k}\left(M\right)\right)$ consisting of the (underlying polyhedra of) derived neighborhoods of $\blacktriangle_{k}\left(M\right)$ in $\operatorname{SP}^{k}\left(M\right)$.
Definition 3.14.
Let $\mathcal{C}$ be a pointed $\infty$-category with small limits and filtered colimits, let $M$ be a manifold that admits a PL structure, and let $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ be a functor. We define a functor $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right):\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ by
$\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(U\right)\right)=\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)^{\mathrm{op}}}\Gamma\left(F;Q\cap B_{k}\left(U\right)\right).$
Since $\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)^{\mathrm{op}}$ is weakly contractible (as it is filtered), there is a natural transformation
$\Gamma\left(F;B_{k}\left(-\right)\right)\simeq\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(-\right)\right)^{\mathrm{op}}}\Gamma\left(F;Q\cap B_{k}\left(-\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$
We write $\Gamma_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ for the fiber of this map.
Remark 3.15.
In the situation of Definition 3.14, the colimit defining $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ exists. To see this, fix a triangulation of $U$. Since the inclusion $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)\to\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)$ is initial (Lemma C.18), we only need to take the colimit over $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$. Since $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ is weakly contractible (being cofiltered by Lemma C.18) and every inclusion $Q\to Q^{\prime}$ in $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ induces a homotopy equivalence $Q\cap B_{k}\left(U\right)\xrightarrow{\simeq}Q^{\prime}\cap B_{k}\left(U\right)$, the colimit over $\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$ indeed exists. Moreover, this argument shows that the map
$\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(U\right)\right)$
is an equivalence for every $U\in\operatorname{Open}\left(M\right)$ and $Q\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$.
We can now state the explicit formula for the inverse equivalence. In the case where $\mathcal{C}$ is the $\infty$-category of pointed spaces, the formula is due to Weiss, and the proof we present below should be compared with his original argument [Wei99, Section 7].
Theorem 3.16.
Let $\mathcal{C}$ be an $\infty$-category with small limits. For every manifold $M$ admitting a PL structure and every $k\geq 0$, the functor
$\Gamma_{\blacktriangle}\left(-;B_{k}\left(-\right)\right):\operatorname{Fun}\left(\operatorname{Sing}B_{k}\left(M\right),\mathcal{C}\right)\to\operatorname{Homog}_{\mathrm{istp},\mathrm{exh}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$
is well-defined and an inverse of the equivalence of Proposition 3.2.
Proof.
In light of Proposition 3.7, it will suffice to show that for each functor $F:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$, the map $\Gamma\left(F;B_{k}\left(-\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ exhibits $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$ as a $\left(k-1\right)$th polynomial approximation of $\Gamma\left(F;B_{k}\left(-\right)\right)$. We will prove this in several steps. For notational convenience, we will write $\Phi=\Gamma\left(F;B_{k}\left(-\right)\right)$ and $\Psi=\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(-\right)\right)$. Note that these functors can take closed subsets of $M$ as its inputs by extending them by the same defining formula.
  1. (Step1)
    We show that, for each $U\in\operatorname{Disj}\left(M\right)$, the map
    $\theta:\Gamma\left(F;B^{\prime\prime}_{k}\left(U\right)\right)\to\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle^{k}\left(U\right)\right)}\Gamma\left(F;B_{k}\left(U\right)\cap Q\right)$
    is an equivalence. Identify each component of $U$ with $\mathbb{R}^{n}$ by choosing a homeomorphism. Given a decreasing continuous map $\varepsilon:[0,\infty)\to(0,1)$, let $\widetilde{Q}_{\varepsilon}\subset U^{k}$ denote the open set of the points $\left(p_{1},\dots,p_{k}\right)\in U^{k}$ satisfying the following conditions:
    • The map $\{p_{1},\dots,p_{k}\}\to\pi_{0}\left(U\right)$ is not injective.
    • For each component $U^{\prime}\subset U$ containing at least one point in $\{p_{1},\dots,p_{k}\}$, we have $\max_{p_{i},p_{j}\in U^{\prime}}\left|p_{i}-p_{j}\right|<\varepsilon\left(\max_{p_{i}\in U^{\prime}}\left|p_{i}\right|\right)$.
    Let $Q_{\varepsilon}\subset\operatorname{SP}^{k}\left(U\right)$ denote the image of $\widetilde{Q}_{\varepsilon}$. Notice that $Q_{\varepsilon}$ is a neighborhood of $\blacktriangle^{k}\left(U\right)$ in $\operatorname{SP}^{k}\left(U\right)$, and moreover that every neighborhood of $\blacktriangle^{k}\left(U\right)$ contains a neighborhood of the form $Q_{\varepsilon}$. Thus we obtain an equivalence
    $\operatorname{colim}_{\varepsilon}\Gamma\left(F;Q_{\varepsilon}\cap B_{k}\left(U\right)\right)\xrightarrow{\simeq}\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle^{k}\left(U\right)\right)}\Gamma\left(F;Q\cap B_{k}\left(U\right)\right).$
    On the other hand, the inclusion $Q_{\varepsilon}\cap B_{k}\left(U\right)\to B^{\prime\prime}_{k}\left(U\right)$ is a homotopy equivalence. To see this, choose a homeomorphism $f:[0,\infty)\to[0,1)$. By using, for each component $U^{\prime}\subset U$ intersecting $\{p_{1},\dots,p_{k}\}$, the map
    $\{p_{i}\}_{i\in U^{\prime\prime}}\mapsto\{f\left(\max_{p_{s}\in U^{\prime}}\left|p_{s}\right|\right)\varepsilon\left(\max_{p_{s}\in U^{\prime}}\left|p_{s}\right|\right)p_{i}\}_{i\in U^{\prime\prime}},$
    we obtain a map $B^{\prime\prime}_{k}\left(U\right)\to Q_{\varepsilon}\cap B_{k}\left(U\right)$. Using linear homotopies, one verifies that this is an inverse homotopy equivalence of the inclusion. (The above map and the homotopies are well-defined because $\varepsilon$ is a decreasing function with values in $\left(0,1\right)$.) So we obtain another equivalence
    $\Gamma\left(F;B^{\prime\prime}_{k}\left(U\right)\right)\xrightarrow{\simeq}\operatorname{colim}_{\varepsilon}\Gamma\left(F;Q_{\varepsilon}\cap B_{k}\left(U\right)\right).$
    The map $\theta$ is the composite of these maps, so it is an equivalence, as required.
  2. (Step2)
    Let $T_{k-1}\Phi$ and $T_{k-1}\Psi$ denote the right Kan extensions of $\Phi$ and $\Psi$ along the inclusion $\operatorname{Disj}^{\leq k-1}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$. We then have the following diagram:
    $\Phi$$\Psi$$T_{k-1}\Phi$$T_{k-1}\Psi$$\scriptstyle{\eta_{\Phi}}$$\scriptstyle{\eta_{\Psi}}$$\scriptstyle{\alpha}$
    The map $\alpha$ is an equivalence by Step 1 (because $B^{\prime\prime}_{k}\left(U\right)=B_{k}\left(U\right)$ for every $U\in\operatorname{Disj}^{\leq k-1}\left(M\right)$). Consequently, it will suffice to show that $\eta_{\Psi}$ is an equivalence. We will prove this in the next step.
  3. (Step3)
    We show that $\eta_{\Psi}$ is an equivalence. By Remark 1.15, it suffices to show that every open set $U\subset M$ equipped with a triangulation coming from a PL structure of $M$ satisfies the following conditions:
    1. (I)
      If $U\in\operatorname{Disj}\left(M\right)$, then the map $\eta_{\Psi,U}:\Psi\left(U\right)\to T_{k-1}\Psi\left(U\right)$ is an equivalence.
    2. (II)
      Let $K_{0}\subset K_{1}\subset\cdots\subset U$ be a sequence of subcomplexes whose interiors cover $U$. The map
      $\Psi\left(U\right)\to\operatorname{lim}_{i}\Psi\left(K_{i}\right)$
      is an equivalence.
    3. (III)
      Let $A_{0},\dots,A_{k-1}\subset U$ be pairwise disjoint compact subcomplexes. The map
      $\Psi\left(U\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(\{0,\dots,k-1\}\right)}\Psi\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
      is an equivalence.
    For (I), we consider the diagram
    $\Phi(U)$$\Gamma(F;B^{\prime\prime}_{k}(U))$$\Psi(U)$$\operatorname{lim}_{V\in\operatorname{Disj}^{\leq k-1}(U)}\Gamma(F;B_{k}(V))$$T_{k-1}\Psi(U).$$\scriptstyle{\eta_{\Phi,U}}\scriptstyle{\simeq}$$\scriptstyle{\psi}\scriptstyle{\simeq}$$\scriptstyle{\phi}$$\scriptstyle{\eta_{\Psi,U}}$$\scriptstyle{\alpha_{U}}\scriptstyle{\simeq}$
    We saw in the proof of Proposition 3.7 that $\phi$ is an equivalence, and in Step 1 that $\psi$ is an equivalence. Thus $\eta_{\Psi,U}$ is an equivalence, as claimed.
    For (II), use Example C.12 and triangulate $\operatorname{SP}^{k}\left(U\right)$ in such a way that it contains $\blacktriangle_{k}\left(U\right)$ as a subcomplex. As we observed in Remark 3.15, for an arbitrary element $Q\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(U\right)\right)$, the map
    $\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\Psi\left(U\right)$
    is an equivalence. A similar argument shows that the map
    $\Gamma\left(F;Q\cap B_{k}\left(K_{i}\right)\right)\to\Psi\left(K_{i}\right)$
    is an equivalence for every $i$, since $Q\cap\operatorname{SP}^{k}\left(K_{i}\right)\in\operatorname{Nbd}^{\prime}\left(\blacktriangle_{k}\left(K_{i}\right)\right)$ (Remark C.17). Consequently, we are reduced to showing that the map
    $\Gamma\left(F;Q\cap B_{k}\left(U\right)\right)\to\operatorname{lim}_{i}\Gamma\left(F;Q\cap B_{k}\left(K_{i}\right)\right)$
    is an equivalence. For this, it suffices to show that $\operatorname{Sing}\left(Q\cap B_{k}\left(U\right)\right)=\bigcup_{i\geq 0}\operatorname{Sing}\left(Q\cap B_{k}\left(K_{i}\right)\right)$ (by Remark B.10 and the fact filtered colimits of simplicial sets already compute homotopy colimits), which follows by inspection.
    For (III), subdividing $U$ if necessary, we may assume that no simplex of $U$ intersects $A_{i}$ and $A_{j}$ for distinct indices $i\neq j$. Let $C_{i}$ denote the simplicial complement of $A_{i}$ in $U$. For each subset $S\subset\{0,\dots,k\}$, the inclusion $U\setminus\bigcup_{i\in S}A_{i}\to C_{i}$ is a homotopy equivalence (see Remark C.15), so it suffices to show that the map
    $\Psi\left(U\right)\to\operatorname{lim}_{S\in\mathcal{P}_{0}\left(\{0,\dots,k-1\}\right)}\Psi\left(C_{i}\right)$
    is an equivalence. This follows from an argument similar to the one in the previous paragraph (and [Lur17, Theorem A.3.1]), noting that if $Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)$ is sufficiently small, we have $B_{k}\left(U\right)\cap Q=\bigcup_{i=0}^{k-1}B_{k}\left(C_{i}\right)\cap Q$. ∎
As a corollary of Proposition 3.7, we give a connectivity estimate of polynomial approximations. When $\mathcal{C}$ is the category of chain complexes of abelian groups or that of spectra, this is proved in [RW14, Lemma E.4] by a rather ad-hoc argument that reduces the claim to that of pointed spaces. Proposition 3.7 allows us to get away with this reduction while keeping the essence of the argument of [RW14] intact.
Corollary 3.17.
Let $M$ be an $n$-manifold admitting a PL structure, let $\mathcal{C}$ be a stable $\infty$-category with small limits and a t-structure, and let $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ be an isotopy functor. Suppose that:
  1. (i)
    For each $U\in\operatorname{Disj}\left(M\right)$, the object $F\left(U\right)\in\mathcal{C}$ is $a$-connected.
  2. (ii)
    0-connected objects are stable under small products.
Then for every $k\geq 0$, the values of $T_{k}F$ are $\left(a-kn-k\right)$-connected.
Proof.
Shifting degrees if necessary, we may assume that $a=0$. We set $\lambda_{k}=-kn-k$. The proof proceeds by induction on $k$.
For the base step where $k=0$, $T_{0}F$ is the constant diagram at $F\left(\emptyset\right)$, so the claim is trivial. For the inductive step, let $k\geq 1$, and suppose we have proved the claim for $T_{k-1}F$. We have a fiber sequence
$L_{k}F\to T_{k}F\to T_{k-1}F$
in $\operatorname{Fun}\left(\operatorname{Open}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$. By the induction hypothesis, the values of $T_{k-1}F$ are $\lambda_{k-1}$-connected, and hence are $\lambda_{k}$-connected. Thus, by the long exact sequence of homotopy groups, we are reduced to showing that $L_{k}F$ is $\lambda_{k}$-connected. We prove this in several steps.
  1. (Step1)
    We show that, for each $U\in I_{M}^{=k}$, the object $\left(L_{k}F\right)\left(U\right)\in\mathcal{C}$ is $\left(-k\right)$-connected. By Remark 3.6 and condition (i), $\left(L_{k}F\right)\left(U\right)$ can be written as a total fiber of a diagram $1]^{k}\to\mathcal{C}_{\geq 1}$. Since the fiber of maps of $a$-connected objects is $\left(a-1\right)$-connected, the claim thus follows by induction on $k$ and the inductive formula for total fibers [ACB22, Corollary 2.2].
  2. (Step2)
    We will show that the values of $L_{k}F$ are $\lambda_{k}$-connected. Let $H:\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{C}$ denote the functor corresponding to the homogeneous functor $L_{k}F$. For each $U\in\operatorname{Open}\left(M\right)$, Proposition 3.16 gives us an equivalence
    $L_{k}F\left(U\right)$ $\simeq\Gamma_{\blacktriangle}\left(H;U\right)$
    $\simeq\mathrm{fib}\left(\Gamma\left(H;B_{k}\left(U\right)\right)\to\Gamma\left(H;B_{k}\left(U\right)\cap Q\right)\right)$
    $\simeq\mathrm{fib}\left(\operatorname{lim}_{\operatorname{Sing}B_{k}\left(U\right)}H\to\operatorname{lim}_{\operatorname{Sing}\left(B_{k}\left(U\right)\cap Q\right)}H\right),$
    where $Q$ is a derived neighborhood of $\blacktriangle_{k}\left(U\right)$ in $\operatorname{SP}^{k}\left(U\right)$ with respect to some triangulation of $U$. Since $B_{k}\left(U\right)$ is obtained from $B_{k}\left(U\right)\cap Q$ by attaching cells of dimension $\leq kn$, combining Step 1 with Lemma B.17 (and condition (ii)), we deduce that the map
    $\operatorname{lim}_{\operatorname{Sing}B_{k}\left(U\right)}H\to\operatorname{lim}_{\operatorname{Sing}\left(B_{k}\left(U\right)\cap Q\right)}H$
    is $\left(\lambda_{k}+1\right)$-connected. Thus its fiber is $\lambda_{k}$-connected, completing the induction.∎
Remark 3.18.
[…]
4. Context-Free Case
In many cases, functors subject to the analysis of manifold calculus are restrictions of functors defined on all of $\mathcal{M}\mathsf{fld}_{n}$ or its variants. The study of such functors is called context-free manifold calculus [Tur13]. Some of the results of Section 1 generalize to the context-free case, which we record in this section.
Definition 4.1.
Let $n,k\geq 0$. Let $\mathcal{C}$ be an an $\infty$-category and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor.
  1. (1)
    The Weiss $k$-topology on $\mathsf{Mfld}_{n}$ is the Grothendieck topology on $\mathsf{Mfld}_{n}$ such that for each object $M\in\mathsf{Mfld}_{n}$, a sieve $\mathcal{U}$ on $M$ is a covering sieve if and only if for each finite set $S\subset M$ of cardinality $\leq k$, there is an object $\left(U\to M\right)\in\mathcal{U}$ such that $S\subset U$. The intersection of the Weiss topologies on $M$ are called the Weiss topology.
  2. (2)
    We say that $F$ is a Weiss $k$-cosheaf (resp. Weiss cosheaf) if the composite $\mathsf{Mfld}_{n}\to\mathcal{M}\mathsf{fld}_{n}\xrightarrow{F}\mathcal{C}$ is a cosheaf with respect to the Weiss $k$-topology (resp. Weiss topology).
  3. (3)
    We say that $F$ is $k$-excisive, or polynomial of degree $\leq k$, if for each $n$-manifold $M$ and each pairwise disjoint closed sets $A_{0},\dots,A_{k}\subset M$, the map
    $\operatorname{colim}_{\emptyset\neq S\subset\{0,\dots,k\}}F\left(M\setminus\bigcup_{i\in S}A_{i}\right)\to F\left(M\right)$
    is an equivalence.
  4. (4)
    We say that $F$ is exhaustive if for each $n$-manifold $M$ and each increasing sequence $U_{0}\subset U_{1}\subset\cdots$ of open sets of $M$ which covers $M$, the map
    $\operatorname{colim}_{i}F\left(U_{i}\right)\to F\left(M\right)$
    is an equivalence.
  5. (5)
    We will write $\operatorname{Exc}_{\mathrm{exh}}^{k}\left(\mathcal{M}\mathsf{fld}_{n};\mathcal{C}\right)\subset\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ for the full subcategory spanned by the $k$-excisive, exhaustive functors $\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$.
  6. (6)
    The Taylor cotower of $F$ is the Taylor cotower of $F$ with respect to the sequence of full subcategories $\mathcal{D}\mathsf{isk}_{n}^{\leq 0}\subset\mathcal{D}\mathsf{isk}_{n}^{\leq 1}\subset\cdots$ (Definition 2.1).
Remark 4.2.
In light of the constructions in non-context-free manifold calculus, some readers find it more natural to work with functors on $\mathsf{Mfld}_{n}$ which carry isotopy equivalences to equivalences, rather than functors defined on $\mathcal{M}\mathsf{fld}_{n}$. There is a reason for this: The localization theorem for isotopy equivalences (Theorem A.1), which formed a basis of our argument in non-context-free manifold calculus, is significantly more subtle in the context-free case. We will come back to this point in Appendix A.
Remark 4.3.
The Taylor cotower of $F$ coincides with the tower of Kurannich–Kupers [KK24, 5.3.3], both in the smooth case and the topological case. See [KK24, Remark 5.8].
The following theorems, which are the main result of this section, are analogs of Theorems 1.3, 1.9 and 1.10 in the context-free case.
Theorem 4.4.
Let $n,k\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss $k$-condition (Definition 1.8), the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
    is an equivalence.
  3. (3)
    $F$ is a Weiss $k$-cosheaf.
  4. (4)
    $F$ is $k$-excisive and exhaustive.
In particular, the inclusion $\operatorname{Exc}_{\mathrm{exh}}^{k}\left(\mathcal{M}\mathsf{fld}_{n};\mathcal{C}\right)\hookrightarrow\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ admits a right adjoint, which carries a functor $F\in\operatorname{Fun}\left(\mathcal{M}\mathsf{fld}_{n},\mathcal{C}\right)$ to a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$.
Theorem 4.5.
Let $n\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. The following conditions are equivalent:
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to\operatorname{Open}\left(M\right)$ satisfying the Weiss condition (Definition 1.8), the map
    $\operatorname{colim}_{I\in N\left(\mathcal{I}\right)}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
    is an equivalence.
  3. (3)
    $F$ is a Weiss cosheaf.
Remark 4.6.
Theorems 4.4 and 4.5 remain valid if we replace $\mathcal{M}\mathsf{fld}_{n}$, $\mathcal{D}\mathsf{isk}_{n}$, and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ by $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$, and $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, with essentially the same proof.
The proofs of Theorems 4.4 and 4.5 are very similar, so we will focus on Theorem 4.4.
Proof of Theorem 4.4.
[…]
$\operatorname{colim}_{I\in\mathcal{I}}F\left(\chi\left(I\right)\right)\to F\left(M\right)$
is an equivalence. In light of Theorem 1.9, it suffices to show that the composite
$\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{n}\xrightarrow{F}\mathcal{C}$
is a left Kan extension of $\operatorname{Disj}^{\leq k}\left(M\right)$. Since $F$ is a left Kan extension of $F|\mathcal{D}\mathsf{isk}_{n}^{\leq k}$, it suffices to show that, for each open set $U\subset M$, the map
$\operatorname{Disj}^{\leq k}\left(M\right)_{/U}\to\mathcal{D}\mathsf{isk}_{n/U}^{\leq k}$
is final. This follows from Lemma 1.11 (or from Theorem A.1, for localization functors are final [Lur25, Tag 02N9]).
[…]
As a corollary of the theorems, we obtain the following structural result on Taylor cotowers.
Corollary 4.7.
Let $n\geq 0$, let $\mathcal{C}$ be an $\infty$-category with small colimits, and let $F:\mathcal{M}\mathsf{fld}_{n}\to\mathcal{C}$ be a functor. Then:
  1. (1)
    For each $k\geq 0$, the $k$th level $T_{k}F$ of the Taylor cotower of $F$ is polynomial of degree $\leq k$ and exhaustive.
  2. (2)
    The Taylor cotower of $F$ is convergent if and only if it satisfies the equivalent conditions of Theorem 4.5.
Proof.
Part (1) follows from Theorem 4.4. Part (2) is a consequence of Proposition 2.4. ∎
Remark 4.8.
Theorem 4.4 admits a following variant: Suppose we are given a right fibration $\pi:\mathcal{M}\to\mathcal{M}\mathsf{fld}_{n}$. Set $\mathcal{D}=\mathcal{D}\mathsf{isk}_{n}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$ and $\mathcal{D}^{\leq k}=\mathcal{D}\mathsf{isk}_{n}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}$. Given an $\infty$-category $\mathcal{C}$ with small colimits, we say that a functor $F:\mathcal{M}\to\mathcal{C}$ is a Weiss $k$-cosheaf if for each $n$-manifold $M$ and each functor $\operatorname{Open}\left(M\right)\to\mathcal{M}$ rendering the diagram
$\mathcal{M}$$\operatorname{Open}(M)$$\mathcal{M}\mathsf{fld}_{n}$}}
commutative, the composite $\operatorname{Open}\left(M\right)\to\mathcal{M}\xrightarrow{F}\mathcal{C}$ is a Weiss $k$-cosheaf. We define what it means for a functor $\mathcal{M}\to\mathcal{C}$ to be a Weiss cosheaf, $k$-excisive, or exhaustive similarly. Then the following conditions for a functor $F:\mathcal{M}\to\mathcal{C}$ are equivalent:
  1. (1)
    $F$ is a left Kan extension of $F|\mathcal{D}^{\leq k}$.
  2. (2)
    For each $n$-manifold $M$, each small $\infty$-category $\mathcal{I}$, and each functor $\chi:\mathcal{I}\to N\left(\operatorname{Open}\left(M\right)\right)$ satisfying the Weiss $k$-condition (Definition 1.8), the map
    $\operatorname{colim}_{I\in\mathcal{I}}F\left(f\left(\chi\left(I\right)\right)\right)\to F\left(f\left(M\right)\right)$
    is an equivalence, where $f:\operatorname{Open}\left(M\right)\to\mathcal{M}$ is any functor lifting $\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{n}$.
  3. (3)
    $F$ is a Weiss $k$-cosheaf.
  4. (4)
    $F$ is $k$-excisive and exhaustive.
The proof is essentially the same as that of Theorem 4.4, using the following observations:
  • For each object $\overline{M}\in\mathcal{M}$ with image $M\in\mathcal{M}\mathsf{fld}_{n}$, the functor $\mathcal{D}_{/\overline{M}}\to\mathcal{D}\mathsf{isk}_{n/M}$ is a trivial fibration (because $\pi$ is a right fibration).
  • In point (2), the composite $\operatorname{Open}\left(M\right)\to\mathcal{M}\xrightarrow{F}\mathcal{C}$ is an isotopy functor, because $\pi$ is conservative (being a right fibration).
We can similarly prove a variant of Theorem 4.5 for $\mathcal{M}$ and $\mathcal{D}$; we can also consider right fibrations over $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$.
One prominent source of right fibrations over $\mathcal{M}\mathsf{fld}_{n}$ comes from framed manifolds. Let $B\mathrm{Top}\left(n\right)\subset\mathcal{M}\mathsf{fld}_{n}$ denote the full subcategory spanned by $\mathbb{R}^{n}$. Given a map $p:B\to B\mathrm{Top}\left(n\right)$ of Kan complexes, the $\infty$-category of $B$-framed $n$-manifolds [AF15, Definition 2.7] is modeled by the fiber product
$\mathcal{M}\mathsf{fld}_{n}^{B}=\mathcal{M}\mathsf{fld}_{n}\times_{\mathcal{S}_{/B\mathrm{Top}\left(n\right)}}\mathcal{S}_{/p}.$
Since the functor $\mathcal{S}_{/p}\to\mathcal{S}_{/B\mathrm{Top}\left(n\right)}$ is a right fibration, so is its pullback $\mathcal{M}\mathsf{fld}_{n}^{B}\to\mathcal{M}\mathsf{fld}_{n}$. Therefore, we have analogs of Theorem 4.4 and 4.5 for $B$-framed manifolds and $B$-framed disks.
5. Boundary Case
The results in Sections 1 through 4 extend to the case of manifolds with boundaries, with very little modification to the arguments. We list the key changes below, leaving the details to the reader. Throughout this section, we fix an integer $n\geq 1$ and an $\left(n-1\right)$-manifold $Z$.
  • We write $\mathcal{M}\mathsf{fld}_{Z}$ for the homotopy coherent nerve of the topological category whose objects are $n$-manifolds $M$ with boundary equipped with a homeomorphism $\partial M\cong Z$, and whose mapping spaces are given by the subspace $\operatorname{Emb}_{\partial}\left(M,N\right)\subset\operatorname{Emb}\left(M,N\right)$ for the subspace consisting of the open embeddings $f:M\to N$ extending the boundary identifications. We also define (the nerve of) an ordinary category $\mathsf{Mfld}_{Z}$ similarly. From section 3 onward, $\mathcal{M}\mathsf{fld}_{n}$ must be replaced by $\mathcal{M}\mathsf{fld}_{Z}$, and whenever a term “manifold” appears in a definition, theorem, proposition, etc, we must replace it by a “manifold with boundary ZZ”.
  • We write $\mathsf{Disk}_{Z}\subset\mathsf{Mfld}_{Z}$ for the full subcategory spanned by the objects isomorphic to $(Z\times\mathbb{R}_{\geq 0})\amalg(\mathbb{R}^{n}\times S)$ for some finite set $S$. We then write $\mathcal{D}\mathsf{isk}_{Z}\subset\mathcal{M}\mathsf{fld}_{Z}$ for the full subcategory spanned by the objects of $\mathsf{Disk}_{Z}$. We define $\infty$-categories $\mathcal{D}\mathsf{isk}_{Z}^{\leq k}$ and $\mathsf{Disk}_{Z}^{\leq k}$ in the obvious manner. From section 3 onward, $\mathcal{D}\mathsf{isk}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ must be replaced by $\mathcal{D}\mathsf{isk}_{Z}$ and $\mathsf{Disk}_{Z}^{\leq k}$.
  • From section 3 onward, $\operatorname{Open}\left(M\right)$ must be replaced by the poset $\operatorname{Open}_{\partial}\left(M\right)$ of open sets of $M$ containing $\partial M$. For each $U\in\operatorname{Open}_{\partial}\left(M\right)$, we write $U_{\partial}\subset U$ for the union of the components that intersect $\partial M$, and set $U_{\mathrm{in}}=U\setminus U_{\partial}$.
  • From section 3 onward, we must replace $\operatorname{Disj}\left(M\right)$ by the full subposet $\operatorname{Disj}_{\partial}\left(M\right)\subset\operatorname{Open}_{\partial}\left(M\right)$ of those $U$ of the form $U_{\partial}\amalg U_{\mathrm{in}}$, where $U_{\partial}$ is a neighborhood of $M$ and $U_{\mathrm{in}}$ is an open set of $M^{\circ}$. A similar remark applies to $\operatorname{Disj}^{\leq k}$. We then define $I_{M}^{=k}$ as the non-full subposet of $\operatorname{Disj}_{\partial}^{\leq k}\left(M\right)$ spanned by the inclusions $U\subset V$ such that $\pi_{0}\left(U\right)\to\pi_{0}\left(V\right)$ is bijective and $\left|\pi_{0}\left(U_{\mathrm{in}}\right)\right|=k$.
  • One of the key ingredient of Section 3 is Ayala–Francis’s theorem (Theorem A.1), and we need this for for $M\in\mathcal{M}\mathsf{fld}_{Z}$ (with various disk categories replaced by the $Z$-counterparts). An inspection reveals that we only need the following generalization of Proposition C.9:
    Proposition 5.1.
    For every finite set $S$ and every morphism $\phi:M\to N$ in $\mathsf{Mfld}_{Z}$, the square
    $\operatorname{Emb}_{\partial}(Z\times[0,\infty))\amalg(\mathbb{R}^{n}\times S),M)$$\operatorname{Emb}_{\partial}((Z\times[0,\infty))\amalg(\mathbb{R}^{n}\times S),N)$$\operatorname{Conf}(S,M^{\circ})$$\operatorname{Conf}(S,N^{\circ})$}}
    is homotopy cartesian.
    Proof.
    For each continuous map $\varepsilon:\partial M\to(0,\infty)$, we define
    $Z\left(\varepsilon\right)=\{\left(p,t\right)\in\partial M\times\mathbb{R}_{\geq 0}\mid 0\leq t<\varepsilon\left(p\right)\}.$
    We then define a simplicial set $\operatorname{Germ}_{\partial}\left(S,M\right)$ as the colimit
    $\operatorname{Germ}_{\partial}\left(S,M\right)=\operatorname{colim}_{r,\varepsilon}\operatorname{Sing}\operatorname{Emb}_{\partial}\left(Z\left(\varepsilon\right)\amalg\left(S\times B^{n}\left(r\right)\right),M\right),$
    where the colimit is indexed over all positive reals $r>0$ and all continuous maps $\varepsilon:Z\to(0,\infty)$. The map $\operatorname{Emb}_{\partial}\left((Z\times[0,\infty))\amalg(\mathbb{R}^{n}\times S),M\right)\to\operatorname{Germ}_{\partial}\left(S,M\right)$ is a homotopy equivalence, so it suffices to show that the square
    $\operatorname{Germ}_{\partial}(S,M)$$\operatorname{Germ}_{\partial}(S,N)$$\operatorname{Sing}\operatorname{Conf}(S,M^{\circ})$$\operatorname{Sing}\operatorname{Conf}(S,N^{\circ})$}}
    is homotopy cartesian. A choice of a collar of $M$ gives us an isomorphism of simplicial sets $\operatorname{Germ}_{\partial}\left(S,M\right)\cong\operatorname{Germ}_{\partial}\left(\emptyset,Z\times[0,1)\right)\times\operatorname{Germ}\left(S,M^{\circ}\right)$, so the claim follows from Proposition C.9. ∎
  • In the definition of $k$-excisive functors (Definitions 1.2 and 4.1), we must assume that the closed sets $A_{i}$ lie in $M^{\circ}$.
  • In Proposition 3.3, Corollary 3.5, Proposition 3.7, and Theorem 3.16, we must replace $B_{k}\left(-\right)$ by $B_{k}\left(\left(-\right)^{\circ}\right)$.
  • Lemma 3.9 will be a claim about $U\in\mathcal{D}\mathsf{isk}_{Z}$ and the map
    $\int^{V\in\operatorname{Disj}_{\partial}^{\leq k-1}\left(U\right)}\operatorname{Sing}B_{k}\left(V^{\circ}\right)\to\operatorname{Sing}\left(B_{k}\left(U^{\circ}\right)\setminus B^{\prime}_{k}\left(U_{\mathrm{in}}\right)\right).$
  • In Proposition 3.7, $B_{k}\left(U\right)$ in the definition of $\widetilde{F}$ should be replaced by $B_{k}\left(U_{\mathrm{in}}\right)$.
  • In the definition of fat diagonal (Definition 3.13), we must adjoin to $\blacktriangle^{k}M$ the images of the points $\left(p_{1},\dots,p_{k}\right)\in M^{k}$ such that some $p_{i}$ lies in $\partial M$.
  • In Definition 3.14, we must consider the functor $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(\left(-\right)^{\circ}\right)\right):\operatorname{Open}_{\partial}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ defined by
    $\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(U^{\circ}\right)\right)=\operatorname{colim}_{Q\in\operatorname{Nbd}\left(\blacktriangle_{k}\left(U\right)\right)^{\mathrm{op}}}\Gamma\left(F;Q\cap B_{k}\left(U^{\circ}\right)\right).$
    We must also consider the functor $\Gamma\left(F;B_{k}\left(\left(-\right)^{\circ}\right)\right)$. We then define $\Gamma_{\blacktriangle}\left(F;B_{k}\left(\left(-\right)^{\circ}\right)\right)$ as the fiber of the map $\Gamma\left(F;B_{k}\left(\left(-\right)^{\circ}\right)\right)\to\Gamma^{\prime}_{\blacktriangle}\left(F;B_{k}\left(\left(-\right)^{\circ}\right)\right)$.
Appendix A (Non)-Localization Theorems
Let $L:\mathcal{C}\to\mathcal{D}$ be a functor of $\infty$-categories, and let $S$ be a set of morphisms of $\mathcal{C}$. Recall ([Lan21, Definition 2.4.2]) that $L$ is said to exhibit $\mathcal{D}$ as a (Dwyer–Kan) localization of $\mathcal{C}$ with respect to $S$ it satisfies the following conditions:
  • The functor $L$ carries every morphism in $S$ to an equivalence.
  • For every $\infty$-category $\mathcal{E}$, the functor
    $\operatorname{Fun}\left(\mathcal{D},\mathcal{E}\right)\to\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{E}\right)$
    is a categorical equivalence, where $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{E}\right)$ denotes the full subcategory of $\operatorname{Fun}\left(\mathcal{C},\mathcal{E}\right)$ spanned by the functors $\mathcal{C}\to\mathcal{E}$ carrying every morphism in $S$ to an equivalence.
A recurring theme in homotopy theory is that many nontrivial $\infty$-categories arise as localizations of ordinary categories. A famous theorem of Ayala and Francis [AF15, Proposition 2.19] is one manifestation of this: It asserts that for each $n$-manifold $M$, the functor
$\mathsf{Disk}_{n/M}\to\mathcal{D}\mathsf{isk}_{n/M}$
exhibits $\mathcal{D}\mathsf{isk}_{n/M}$ as a localization of $\mathsf{Disk}_{n/M}$ with respect to isotopy equivalences. The goal of this subsection is to present a detailed proof of this theorem and some of its variants. [^5] The original proof of Ayala and Francis is missing some important justifications. The author learned the current proof from David Ayala and is grateful to him for his explanation.
A.1. Ayala–Francis’s Theorem on Localizations at Isotopy Equivalences
In this subsection, we give a proof of the following theorem. Point (1) is due to Ayala and Francis:
Theorem A.1.
Let $n\geq 0$, and let $M$ be an $n$-manifold.
  1. (1)
    The functor $\mathsf{Disk}_{n/M}\to\mathcal{D}\mathsf{isk}_{n/M}$ exhibits $\mathcal{D}\mathsf{isk}_{n/M}$ as a localization of $\mathsf{Disk}_{n/M}$ with respect to isotopy equivalences.
  2. (2)
    For every $k\geq 0$, the functor $\mathsf{Disk}_{n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ exhibits $\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$ as a localization of $\mathsf{Disk}_{n/M}^{\leq k}$ with respect to isotopy equivalences.
We will establish Theorem A.1 after a few preliminaries. We begin with a variation of Quillen’s Theorem B.
Definition A.2.
[Cis19, Definition 4.6.3] A map $E\to B$ of simplicial sets is said to be locally constant if for each morphism $X\to B$ of simplicial sets, the square
$X\times_{B}E$}}E{E}X{X}B{B}
is homotopy cartesian in the Kan–Quillen model structure.
Proposition A.3 (A Variation of Quillen’s Theorem B).
Let $\pi:E\to B$ be a cocartesian fibration of simplicial sets. The following conditions are equivalent:
  1. (1)
    The map $\pi$ is locally constant.
  2. (2)
    For each edge $\alpha:b\to b^{\prime}$ in $B$, the induced functor
    $\alpha_{!}:E_{b}=E\times_{B}\{b\}\to E_{b^{\prime}}$
    is a weak homotopy equivalence.
Proof.
[…]
$K\times_{B}E\to L\times_{B}E$
is a weak homotopy equivalence. Let $\mathscr{{M}}$ denote the class of morphisms $S\to T$ of simplicial sets such that, for any map $T\to B$, the induced map $S\times_{B}E\to T\times_{B}E$ is a weak homotopy equivalence. We must show that $\mathscr{{M}}$ contains all weak homotopy equivalences. By [Cis19, Proposition 4.6.1], it suffices to show that $\mathscr{{M}}$ contains all morphisms of the form $\Delta^{0}\to\Delta^{n}$, for any $n\geq 0$.
Let $0\leq i\leq n$ be integers and let $\sigma:\Delta^{n}\to B$ be an $n$-simplex, which we depict as $b_{0}\to\cdots\to b_{n}$. We wish to show that the map $E_{b_{i}}\to\Delta^{n}\times_{B}E$ is a weak homotopy equivalence. Choose a cocartesian natural transformation $h:\Delta^{\{i,n\}}\times E_{b_{i}}\to\Delta^{n}\times_{B}E$ rendering the diagram
$\{i\}\times E_{b_{i}}$$\Delta^{n}\times_{B}E$$\Delta^{\{i,n\}}\times E_{b_{i}}$$\Delta^{n}$}}
commutative. It will suffice to show that the restriction $h|\{n\}\times E_{b_{i}}:\{n\}\times E_{b_{i}}\to\Delta^{n}\times_{B}E$ is a weak homotopy equivalence. We can factor this map as
$\{n\}\times E_{b_{i}}\xrightarrow{h^{\prime}}E_{b_{n}}\xrightarrow{h^{\prime\prime}}\Delta^{n}\times_{B}E,$
where $h^{\prime\prime}$ denotes the inclusion. The map $h^{\prime}$ is a weak homotopy equivalence by hypothesis. Also, since $p$ is a cocartesian fibration and the inclusion $\{n\}\subset\Delta^{n}$ is final, [Lur09, Proposition 4.1.2.15] shows that $h^{\prime\prime}$ is final. In particular, $h^{\prime\prime}$ is a weak homotopy equivalence. It follows that the map $h|\{n\}\times E_{b_{i}}$ is a composite of weak homotopy equivalences, for it is the composite of two weak homotopy equivalences. The claim follows. ∎
We use Proposition A.3 to prove the following recognition result for localizations.
Proposition A.4.
Let $f:\mathcal{C}\to\mathcal{D}$ be a functor of $\infty$-categories. Set $\mathcal{W}=\mathcal{C}\times_{\mathcal{D}}\mathcal{D}^{\simeq}$. Suppose that, for each object $C\in\mathcal{C}$, the map $\mathcal{C}_{/C}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}_{/f\left(C\right)}\right)^{\simeq}$ is a weak homotopy equivalence. The following conditions are equivalent:
  1. (1)
    The functor $f$ exhibits $\mathcal{D}$ as a localization of $\mathcal{C}$ with respect to morphisms in $\mathcal{W}$.
  2. (2)
    The map $\mathcal{W}\to\mathcal{D}^{\simeq}$ is a weak homotopy equivalence.
Proof.
We will use the localization criterion using Rezk’s classification diagram, due to Mazel-Gee. For each simplicial set $K$, let $\operatorname{Fun}^{\mathcal{W}}\left(K,\mathcal{C}\right)\subset\operatorname{Fun}\left(K,\mathcal{C}\right)$ denote the subcategory spanned by the natural transformations whose components are morphisms of $\mathcal{W}$. Recall that the classification diagram $N\left(\mathcal{C},\mathcal{W}\right)$ is the simplicial object in $S\SS$ whose $n$th simplicial set is given by
$N\left(\mathcal{C},\mathcal{W}\right)_{n}=\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{n},\mathcal{C}\right)=\operatorname{Fun}\left(\Delta^{n},\mathcal{C}\right)\times_{\mathcal{C}^{n+1}}\mathcal{W}^{n+1}.$
We will write $N\left(\mathcal{D}\right)=N\left(\mathcal{D},\mathcal{D}^{\simeq}\right)$; it is a complete Segal space [Rez01], called the classifying diagram of $\mathcal{D}$. According to [MG19, Theorem 3.8] (see also [Ara23a, Corollary 4.6] and [AC25]), condition (1) is equivalent to the following condition:
  • (1)
    The map $N\left(\mathcal{C},\mathcal{W}\right)\to N\left(\mathcal{D}\right)$ is a weak equivalence of the complete Segal space model structure.
We will show that condition (1) is equivalent to condition (2).
First we show that the map $d_{0}:N\left(\mathcal{C},\mathcal{W}\right)_{1}\to N\left(\mathcal{C},\mathcal{W}\right)_{0}$ is locally constant. Consider the commutative diagram
$\operatorname{Fun}^{\mathcal{W}}(\Delta^{1},\mathcal{C})$$\operatorname{Fun}(\Delta^{1},\mathcal{D})^{\simeq}$$\operatorname{Fun}^{\mathcal{W}}(\{1\},\mathcal{C})$$\operatorname{Fun}(\{1\},\mathcal{D})^{\simeq}.$$\scriptstyle{\phi}$$\scriptstyle{\pi}$$\scriptstyle{\pi^{\prime}}$
We wish to show that $\pi$ is locally constant. According to [Lur25, Tag 0478], the maps $\pi$ and $\pi^{\prime}$ are cocartesian fibrations and $\phi$ carries $\pi$-cocartesian morphisms to $\pi^{\prime}$-cocartesian morphisms. Therefore, for each morphism $\alpha:C\to C^{\prime}$, the square
$\mathcal{C}^{/C}\times_{\mathcal{C}}\mathcal{W}$$(\mathcal{D}^{/f(C)})^{\simeq}$$\mathcal{C}^{/C^{\prime}}\times_{\mathcal{C}}\mathcal{W}$$(\mathcal{D}^{/f(C^{\prime})})^{\simeq}$$\scriptstyle{\alpha_{!}}$$\scriptstyle{f(\alpha)_{!}}$
consisting of fibers of $\pi$ and $\pi^{\prime}$ commutes up to natural equivalence. By hypothesis (and the equivalence of the upper slices and lower slices [Lur09, Proposition 4.2.1.5]), the horizontal arrows are weak homotopy equivalences. Also, since $f\left(\alpha\right)$ is an equivalence, the functor $f\left(\alpha\right)_{!}$ is a homotopy equivalence. Hence $\alpha_{!}$ is a weak homotopy equivalence. It follows from Proposition A.3 that $\pi$ is locally constant, as desired.
[…]
$X_{k}$$X_{1}$$X_{k-1}$$X_{0}$}}
is homotopy cartesian for all $k\geq 2$, where the maps in the diagrams are induced by the inclusions $\{1\}\to\{0,1\}\hookrightarrow[k]$ and $\{1\}\to\{1,\dots,k\}\hookrightarrow[k]$. Consequently, it suffices to show that for each $k\geq 2$, the square
$\operatorname{Fun}^{\mathcal{W}}(\Delta^{k},\mathcal{C})$$\operatorname{Fun}^{\mathcal{W}}(\Delta^{1},\mathcal{C})$$\operatorname{Fun}^{\mathcal{W}}(\Delta^{\{1,\dotsm,k\}},\mathcal{C})$$\operatorname{Fun}^{\mathcal{W}}(\{1\},\mathcal{C})$$\scriptstyle{\pi}$
is homotopy cartesian in the Kan–Quillen model structure. Since $\pi$ is locally constant, this is equivalent to the assertion that the map
$\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{k},\mathcal{C}\right)\to\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1}\amalg_{\{1\}}\Delta^{\{1,\dots,k\}},\mathcal{C}\right)$
be a weak homotopy equivalence. But this map is a trivial fibration, for the inclusion $\Delta^{1}\amalg_{\{1\}}\Delta^{\{1,\dots,k\}}\subset\Delta^{k}$ is a weak categorical equivalence [Lur09, Lemma 5.4.5.10].
[…]
$N\left(\mathcal{C},\mathcal{W}\right)\xrightarrow{i}\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)\xrightarrow{F}N\left(\mathcal{D}\right).$
Since $i$ is a levelwise weak homotopy equivalence, it suffices to show that the map $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{0}\to N\left(\mathcal{D}\right)_{0}$ is a weak homotopy equivalence. By hypothesis, the map $F$ is a weak equivalence of the complete Segal space model structure. Therefore, it suffices to show that the Segal space $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ is a complete Segal space (for weak equivalences of the complete Segal space model structure between complete Segal spaces are nothing but levelwise homotopy equivalences).
Let $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}\subset\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$ denote the union of the components whose vertices are homotopy equivalences of the Segal space $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$. We must show that the map
$\theta:\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{0}\to\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}$
is a homotopy equivalence. Since $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ is a Segal space, the map $F$ is a Dwyer–Kan equivalence of Segal spaces [Rez01, Theorem 7.1]. Therefore, given a morphism $\alpha$ of $\mathcal{C}$, the morphism $i\left(\alpha\right)$ is a homotopy equivalence of $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)$ if and only if $Fi\left(\alpha\right)$ is a homotopy equivalence of $N\left(\mathcal{D}\right)$. By the definition of $\mathcal{W}$, the latter condition holds if and only if $\alpha$ belongs to $\mathcal{W}$. Therefore, the inverse image of $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}\subset\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$ under the weak homotopy equivalence
$i_{1}:\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1},\mathcal{C}\right)\xrightarrow{\simeq}\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{1}$
is the simplicial subset $\operatorname{Fun}\left(\Delta^{1},\mathcal{W}\right)\subset\operatorname{Fun}^{\mathcal{W}}\left(\Delta^{1},\mathcal{C}\right)$. In particular (since $\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}$ is a union of components) the map $i_{1}$ restricts to a weak homotopy equivalence
$i_{1}^{\prime}:\operatorname{Fun}\left(\Delta^{1},\mathcal{W}\right)\xrightarrow{\simeq}\widetilde{N}\left(\mathcal{C},\mathcal{W}\right)_{{\rm hoeq}}.$
Now consider the commutative diagram
$\mathcal{W}$$\operatorname{Fun}(\Delta^{1},\mathcal{W})$$\operatorname{Fun}^{\mathcal{W}}(\Delta^{1},\mathcal{C})$$\widetilde{N}(\mathcal{C},\mathcal{W})_{0}$$\widetilde{N}(\mathcal{C},\mathcal{W})_{\mathrm{hoeq}}$$\widetilde{N}(\mathcal{C},\mathcal{W})_{1}.$$\scriptstyle{i^{\prime}_{1}}$$\scriptstyle{i_{1}}$$\scriptstyle{\theta}$$\scriptstyle{\theta^{\prime}}$$\scriptstyle{i_{0}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$
The maps $i_{0},i_{1}^{\prime},i_{1}$ are weak homotopy equivalences, and so is the map $\theta^{\prime}$ (for it is a left (and a right) adjoint). Therefore, the map $\theta$ is a weak homotopy equivalence, as required.
[…]
$\operatorname{Fun}^{\mathcal{W}}(\Delta^{1},\mathcal{C})$$\operatorname{Fun}(\Delta^{1},\mathcal{D})^{\simeq}$$\operatorname{Fun}^{\mathcal{W}}(\{1\},\mathcal{C})$$\operatorname{Fun}(\{1\},\mathcal{D})^{\simeq}.$$\scriptstyle{\pi}$$\scriptstyle{\pi^{\prime}}$$\scriptstyle{N(f)_{0}}$$\scriptstyle{N(f)_{1}}$
As we already saw in the second paragraph of the proof, the vertical maps are locally constant cocartesian fibrations. (The map $\pi^{\prime}$ is even a Kan fibration). Also, our hypothesis ensures that, for each $C\in\operatorname{Fun}^{\mathcal{W}}\left(\{1\},\mathcal{C}\right)$, the induced map $\pi^{-1}\left(C\right)\to\left(\pi^{\prime}\right)^{-1}\left(f\left(C\right)\right)$ between the fibers is a weak homotopy equivalence. Hence the square is homotopy cartesian in the Kan–Quillen model structure. Since $N\left(f\right)_{0}$ is a weak homotopy equivalence, so must be $N\left(f\right)_{1}$, and the proof is complete. ∎
Corollary A.5.
Let $\mathcal{C}$ and $\mathcal{D}$ be $\infty$-categories, let $\mathcal{C}^{\prime}\subset\mathcal{C}$ and $\mathcal{D}^{\prime}\subset\mathcal{D}$ be full subcategories, and let $f:\mathcal{C}\to\mathcal{D}$ be a functor which carries $\mathcal{C}^{\prime}$ into $\mathcal{D}^{\prime}$. Set $\mathcal{W}=\mathcal{C}\times_{\mathcal{D}}\mathcal{D}^{\simeq}$. Suppose that, for each $C\in\mathcal{C}$, the functor
$\mathcal{C}^{\prime}_{/C}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}^{\prime}_{/f\left(C\right)}\right)^{\simeq}$
is a weak homotopy equivalence. Then for each $C\in\mathcal{C}$, the functor
$\mathcal{C}^{\prime}_{/C}\to\mathcal{D}^{\prime}_{/f\left(C\right)}$
is a localization with respect to the morphisms whose images in $\mathcal{D}^{\prime}_{f/\left(C\right)}$ are equivalences.
Proof.
We apply Proposition A.4 to the functor $\mathcal{C}^{\prime}_{/C}\to\mathcal{D}^{\prime}_{/f\left(C\right)}$. It suffices to show that, for each object $\left(\alpha:C^{\prime}\to C\right)\in\mathcal{C}^{\prime}_{/C}$, the map
$\theta:\mathcal{C}^{\prime}_{/\alpha}\times_{\mathcal{C}}\mathcal{W}\to\left(\mathcal{D}^{\prime}_{/f\left(\alpha\right)}\right)^{\simeq}$
is a weak homotopy equivalence. Consider the commutative diagram
$\mathcal{C}^{\prime}_{/\alpha}\times_{\mathcal{C}}\mathcal{W}$$(\mathcal{D}^{\prime}_{/f(\alpha)})^{\simeq}$$\mathcal{C}^{\prime}_{/C^{\prime}}\times_{\mathcal{C}}\mathcal{W}$$(\mathcal{D}^{\prime}_{/f(C^{\prime})})^{\simeq}.$$\scriptstyle{\theta}$$\scriptstyle{\theta^{\prime}}$
The vertical maps are trivial fibrations, and $\theta^{\prime}$ is a weak homotopy equivalence by hypothesis. Hence $\theta$ is a weak homotopy equivalence, as desired. ∎
To apply Corollary A.5 to our context, we need a lemma. For each $n,k\geq 0$, write $\mathsf{Disk}_{n}^{=k}$ and $\mathcal{D}\mathsf{isk}_{n}^{=k}$ for the full subcategories of $\mathsf{Disk}_{n}$ and $\mathsf{Disk}_{n}$ spanned by the objects homeomorphic to $\mathbb{R}^{n}\times\{1,\dots,k\}$. Note that, by Kister’s theorem [Kis64], the subcategory $\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\subset\mathsf{Disk}_{n/M}^{=k}$ is spanned by the morphisms which induce bijections between the sets of components.
Lemma A.6.
Let $n,k\geq 0$, and let $M$ be an $n$-manifold. The map
$\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{=k}\right)^{\simeq}$
is a weak homotopy equivalence.
For the proof of Lemma A.6, we introduce a bit of notation.
Notation A.7.
Let $k\geq 0$, and let $M$ be a manifold.
  • We will write $B^{\prime}_{k}\left(M\right)\subset B_{k}\left(M\right)$ for the open set consisting of those subsets $S\subset M$ of cardinality $k$ such that the map $\pi_{0}\left(S\right)\to\pi_{0}\left(M\right)$ is injective.
  • Let $k\geq 0$, and let $M$ be a manifold. We let $I_{M}^{=k}$ denote the (non-full) subcategory of $\operatorname{Disj}\left(M\right)$ spanned by the isotopy equivalences between the elements of $\operatorname{Disj}\left(M\right)$ with exactly $k$ components.
Proof of Lemma A.6.
The assignment $U\mapsto\left(U\hookrightarrow M\right)$ induces a categorical equivalence $I_{M}^{=k}\xrightarrow{\simeq}\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}$, so it suffices to show that the composite
$\theta:I_{M}^{=k}\to\mathsf{Disk}_{n/M}^{=k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{=k}\right)^{\simeq}$
is a weak homotopy equivalence.
Given $n$-manifolds $X,Y$, let $\operatorname{Emb}^{\simeq}\left(X,Y\right)\subset\operatorname{Emb}\left(X,Y\right)$ denote the subspace consisting of the isotopy equivalences from $X$ to $Y$. According to Proposition B.4 and [Lur09, Proposition 4.2.4.1], the map $\theta$ is a weak homotopy equivalence if and only if the following condition holds:
  • ($\ast$)
    For each object $U\in\mathcal{D}\mathsf{isk}_{n}^{=k}$, the map
    $\operatorname{hocolim}_{V\in I_{M}^{=k}}\operatorname{Sing}\operatorname{Emb}^{\simeq}\left(U,V\right)\to\operatorname{Sing}\operatorname{Emb}\left(U,M\right)$
    is a weak homotopy equivalence.
We will prove ($\ast$). The claim is obvious if $k=0$, so assume that $k\geq 1$. Fix a homeomorphism $U\cong\mathbb{R}^{n}\times\{1,\dots,k\}$ and let $V_{1},\dots,V_{k}\subset V$ denote the components of $V$. Evaluation at the origin gives us a commutative diagram
$\operatorname{Emb}^{\simeq}(U,V)$$\operatorname{Emb}(U,V)$$\operatorname{Emb}(U,M)$$\coprod_{\sigma\in\Sigma_{k}}\prod_{i=1}^{k}V_{\sigma(i)}$$\operatorname{Conf}(k,V)$$\operatorname{Conf}(k,M)$$B^{\prime}_{k}(V)$$B_{k}(V)$$B_{k}(M)$$\scriptstyle{(1)}$$\scriptstyle{(3)}$$\scriptstyle{(4)}$$\scriptstyle{(2)}$
of topological spaces. Proposition C.9 shows that square (1) is homotopy cartesian. Squares (2), (3), (4) are homotopy cartesian because they are strictly cartesian and their vertical arrows are Serre fibrations. It follows that the outer square is homotopy cartesian. Since colimits in $\mathcal{S}$ are universal [Lur09, Lemma 6.1.3.14], we are reduced to showing that the map
$\operatorname{hocolim}_{V\in I_{M}^{=k}}\operatorname{Sing}B^{\prime}_{k}\left(V\right)\to\operatorname{Sing}B_{k}\left(M\right)$
is a weak homotopy equivalence. This follows from [Lur17, Theoerm A.3.1]. ∎
We now arrive at the proof of Theorem A.1.
Proof of Theorem A.1.
We will apply Corollary A.5 to the functor $\mathsf{Mfld}_{n}\to\mathcal{M}\mathsf{fld}_{n}$ and the full subcategories $\mathsf{Disk}_{n}\subset\mathsf{Mfld}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}\subset\mathcal{M}\mathsf{fld}_{n}$ for part (1), and the full subcategories $\mathsf{Disk}_{n}^{\leq k}\subset\mathsf{Mfld}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}^{\leq k}\subset\mathcal{M}\mathsf{fld}_{n}$ for part (2). For part (1), we must show that for each $n$-manifold $M$, the functor
$\mathsf{Disk}_{n/M}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}\right)^{\simeq}$
is a weak homotopy equivalence, and for part (2), we must show that the functor
$\mathsf{Disk}_{n/M}^{\leq k}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}\right)^{\simeq}$
is a weak homotopy equivalence. Both of these assertions follow from Lemma A.6. ∎
Remark A.8.
The proof of Theorem A.1 carries over to the smooth case (using Remark C.10). In other words, for each $n,k\geq 0$ and each smooth $n$-manifold $M$, the functors $\mathsf{Disk}_{\mathrm{sm},n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n/M}^{\leq k}$ and $\mathsf{Disk}_{\mathrm{sm},n/M}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n/M}$ are localization with respect to smooth isotopy equivalences.
Remark A.9.
[…]
Variant A.10.
Let $n\geq 0$, and let $M$ be an $n$-manifold. Let $\mathcal{O}$ be a basis of the topology of $M$ whose elements are homeomorphic to $\mathbb{R}^{n}$, and let $\mathsf{Disk}_{n/M}^{\mathcal{O}}\subset\mathsf{Disk}_{n/M}$ denote the full subcategory spanned by the objects whose images are finite disjoint union of elements in $\mathcal{O}$. The functor $\mathsf{Disk}_{n/M}^{\mathcal{O}}\to\mathcal{D}\mathsf{isk}_{n/M}$ is a localization at isotopy equivalences.
Indeed, by applying Corollary A.5 to the full subcategories $\mathsf{Disk}_{n/M}^{\mathcal{O}}\subset\left(\mathsf{Disk}_{n/M}^{\mathcal{O}}\right)^{\triangleright}$ and $\mathcal{D}\mathsf{isk}_{n/M}\subset\left(\mathcal{D}\mathsf{isk}_{n/M}\right)^{\triangleright}$, we are reduced to showing that the maps
$\mathsf{Disk}_{n/N}^{\mathcal{O}}\times_{\mathcal{M}\mathsf{fld}_{n}}\mathcal{M}\mathsf{fld}_{n}^{\simeq}\to\left(\mathcal{D}\mathsf{isk}_{n/N}\right)^{\simeq}$
is a weak homotopy equivalence for all $N\in\mathcal{M}\mathsf{fld}_{n}$, which can be proved exactly as in Lemma A.6.
By a similar argument, we find that the functors
$\mathsf{Disk}_{n/M}^{\mathcal{O}}\cap\mathsf{Disk}_{n/M}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n/M}^{\leq k}$
is a localizations at isotopy equivalences for every $k\geq 0$.
A.2. Global case
Ayala–Francis’s theorem (Theorem A.1), the main result of the previous subsection, is crucial in the development of non-context-free manifold calculus. In this section, we consider a context-free analog of the localization theorem. The results in this section will not be used elsewhere in this paper.
It turns out that the localization theorem is a bit nuanced in the context-free case. (see also [AF20, 2.2.13] for a relevant observation, and [DWW03, Corollary 2.7] for a closely related result.) The validity of the theorem depends on which category (topological or smooth) we work in, as the following theorem shows.
Theorem A.11.
Let $n,k\geq 1$.
  1. (1)
    Both of the functors $\mathsf{Disk}_{n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ and $\mathsf{Disk}_{n}\to\mathcal{D}\mathsf{isk}_{n}$ are localizations with respect to isotopy equivalences.
  2. (2)
    Neither of the functors $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$ and $\mathsf{Disk}_{\mathrm{sm},n}\to\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}$ is a localization with respect to smooth isotopy equivalences.
Proof.
We begin with (1). We will focus on the functor $\mathsf{Disk}_{n}\to\mathcal{D}\mathsf{isk}_{n}$; the proof that the functor $\mathsf{Disk}_{n}^{\leq k}\to\mathcal{D}\mathsf{isk}_{n}^{\leq k}$ is a localization is similar. By Proposition A.4 and Lemma A.6, it suffices to show that the functor
$\mathsf{Disk}_{n}\times_{\mathcal{D}\mathsf{isk}_{n}}\mathcal{D}\mathsf{isk}_{n}^{\simeq}\to\mathcal{D}\mathsf{isk}_{n}^{\simeq}$
is a weak homotopy equivalence. For convenience, in this proof we will replace $\mathsf{Disk}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}$ by their full subcategories spanned by the objects $\{\mathbb{R}^{n}\times\{1,\dots,k\}\}_{k=0,1,\dots}$ and still denote them by $\mathsf{Disk}_{n}$ and $\mathcal{D}\mathsf{isk}_{n}$.
Regard the $k$th symmetric group $\Sigma_{k}$ as a category with one object $\{1,\dots,k\}$ with morphism given by bijections, and let $N\Sigma_{k}$ denote its nerve. We consider the commutative diagram
$\mathsf{Disk}_{n}\times_{\mathcal{D}\mathsf{isk}_{n}}\mathcal{D}\mathsf{isk}_{n}^{\simeq}$$\mathcal{D}\mathsf{isk}_{n}^{\simeq}$$\coprod_{k\geq 0}N\Sigma_{k}$$\scriptstyle{q}$$\scriptstyle{p}$
where $p$ and $q$ are given by $f\mapsto\pi_{0}\left(f\right)$. The functors $p$ and $q$ are cartesian fibrations. Since every morphism of $\coprod_{k\geq 0}N\Sigma_{k}$ is an equivalence, both $p$ and $q$ are locally constant (Definition A.2). Therefore, it suffices to show that for each $k\geq 0$, the map
$\theta_{k}:p^{-1}\left(\{1,\dots,k\}\right)\to q^{-1}\left(\{1,\dots,k\}\right)$
is a weak homotopy equivalence.
We can identify the map $\theta_{k}$ with the map
$\left(N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\right)^{k}\to\left(N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)\right)^{k},$
where $N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$ denotes the homotopy coherent nerve of the topological monoid $\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$ (regarded as a topological category with a single object) and $N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}$ denotes the nerve of the same category with the discrete topology. It thus suffices to show that the map
$N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\to N\operatorname{Emb}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$
is a weak homotopy equivalence. This is a consequence of McDuff’s theorem [McD80, Corollary 2.15]. (For a comparison between the homotopy coherent nerve and McDuff’s model of classifying spaces, see [Ara24a, Corollary 4.2].)
Next, we prove (2). As in the previous paragraph, it suffices to show that the map
$N\operatorname{Emb}_{\mathrm{sm}}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)_{\delta}\to N\operatorname{Emb}_{\mathrm{sm}}\left(\mathbb{R}^{n},\mathbb{R}^{n}\right)$
is not a weak homotopy equivalence. Arguing as in [McD80, Proof of Theorem 1.1], we can reduce this to showing that the map $B\Gamma_{n}^{\infty}\to BGL_{n}\left(\mathbb{R}\right)$ is not a weak homotopy equivalence of topological spaces, where $\Gamma_{n}^{\infty}$ denotes Haefliger’s groupoid [Hae71, p.143] for codimension $n$ smooth foliations. (See also Subsection A.3.) This is a consequence of Bott’s theorem [Bot70] (see [Hae71, p. 143, I.8 (a)]). ∎
A.3. A Remark on Classifying Spaces of Groupoids Internal to $\mathsf{Top}$
A little care is necessary in the final step of Theorem A.11, because in [McD80] and [Hae71], McDuff and Haefliger use different models of classifying spaces of groupoids internal to the category $\mathsf{Top}$ of topological spaces. The equivalence of the two models is probably folklore, but the author is not aware of a convenient reference for this. Thus we record a proof below.
We start by recalling McDuff and Haefliger’s models.
Construction A.12.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. Abusing notation, let $\mathcal{G}$ denote the simplicial topological space corresponding to $\mathcal{G}$. (Thus $\mathcal{G}_{1}=\operatorname{mor}\mathcal{G}$, $\mathcal{G}_{2}=\operatorname{mor}\mathcal{G}\times_{\operatorname{ob}\mathcal{G}}\operatorname{mor}\mathcal{G}$, etc.) McDuff and Haefliger define the classifying space of $\mathcal{G}$ as follows:
  • In [McD80], McDuff defines the classifying space of $\mathcal{G}$ to be the fat realization of the simplicial topological space $\mathcal{G}$. We denote this topological space by $B_{{\rm MD}}\left(\mathcal{G}\right)$. [^6] This model was introduced by Segal [Seg68].
  • Haefliger’s model is slightly more complicated. First, consider the space $E_{{\rm Hae}}\mathcal{G}$ whose points are the equivalence classes of formal expressions of the form
    $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$
    where $g_{0},g_{1},\dots$ are morphisms of $\mathcal{G}$ with a common codomain and $t_{0},t_{1},\dots$ are non-negative real numbers, all but finitely many of which are zero, and $\sum_{i=0}^{\infty}t_{i}=1$. Two expressions $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$ and $\left(t^{\prime}_{0},g^{\prime}_{0},t^{\prime}_{1},g^{\prime}_{1},\dots\right)$ are equivalent if $t_{i}=t_{i}^{\prime}$ for all $i$ and $g_{i}=g^{\prime}_{i}$ whenever $t_{i}>0$. The equivalence class of $\left(t_{0},g_{0},t_{1},g_{1},\dots\right)$ will be denoted by $\bigoplus_{i\geq 0}t_{i}g_{i}=t_{0}g_{0}\oplus t_{1}g_{1}\oplus\cdots$. We will write $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$ for the set map $\tau_{i}\left(\bigoplus_{j\geq 0}t_{j}g_{j}\right)=t_{i}$, and define $p_{i}:\tau_{i}^{-1}\left((0,1]\right)\to\operatorname{mor}\mathcal{G}$ by $p_{i}\left(\bigoplus_{j\geq 0}t_{j}g_{j}\right)=g_{i}$. We topologize $E_{{\rm Hae}}\mathcal{G}$ so that it has the following universal property: If $X$ is a topological space, then a set map $X\to E_{{\rm Hae}}\mathcal{G}$ is continuous if and only if the composites $\tau_{i}f:X\to[0,1]$ and $p_{i}f:\left(\tau_{i}f\right)^{-1}\left((0,1]\right)\to\operatorname{mor}\mathcal{G}$ are continuous.
    Now introduce an equivalence relation on $E_{{\rm Hae}}\mathcal{G}$ as follows: Two points $\bigoplus_{j\geq 0}t_{j}g_{j}$ and $\bigoplus_{j\geq 0}t^{\prime}_{j}g^{\prime}_{j}$ are equivalent if there is some morphism $g$ of $\mathcal{G}$ such that $\bigoplus_{j\geq 0}t_{j}gg_{j}=\bigoplus_{j\geq 0}t_{j}gg^{\prime}_{j}$. The resulting quotient space will be denoted by $B_{{\rm Hae}}\mathcal{G}$. We still denote by $\tau_{i}:B_{{\rm Hae}}\mathcal{G}\to[0,1]$ the map induced by $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$. The space $B_{{\rm Hae}}\mathcal{G}$ is the model of the classifying space of $\mathcal{G}$ used by Haefliger in [Hae71]. [^7] This model was introduced by Milnor in [Mil56].
These models are equivalent in the following sense:
Proposition A.13.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. There is a zig-zag of weak homotopy equivalences between $B_{{\rm MD}}\mathcal{G}$ and $B_{{\rm Hae}}\mathcal{G}$, which is natural in $\mathcal{G}$.
[…]
Lemma A.14.
Let $\mathcal{G}$ be a groupoid internal to $\mathsf{Top}$. There is a bijective continuous map
$\phi:\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\to B_{{\rm Hae}}\mathcal{G},$
which is a homotopy equivalence.
Proof.
We will write $X=\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|$. By definition, a point of $X$ can be represented by a sequence
$\left(\left(t_{0},\dots,t_{n}\right),x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n},k_{0}<\cdots<k_{n}\right),$
where $\left(t_{0},\dots,t_{n}\right)$ is a point of $\left|\Delta^{n}\right|$, $x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n}$ is a point of $\mathcal{G}_{n}$, and $k_{0}<\cdots<k_{n}$ is an element of $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)_{n}$. We declare that $\phi$ carries such a point to the point
$\left(\bigoplus_{0\leq i<k_{0}}0\right)\oplus t_{0}f_{n}\dots f_{1}\oplus\left(\bigoplus_{k_{0}<i<k_{1}}0\right)\oplus t_{1}f_{n}\dots f_{2}\oplus\cdots\oplus t_{n}\operatorname{id}_{x_{n}}\oplus\left(\bigoplus_{i<k_{n}}0\right).$
This defines a bijection $\phi:X\to B_{{\rm Hae}}\mathcal{G}$, which is continuous by the definition of the topology on $B\mathcal{G}$.
To show that $\phi$ is a homotopy equivalence, we first construct a homotopy inverse of $\phi$. Use [tD08, 13.1.7] to find a partition of unity $\left(\psi_{j}\right)_{j=0}^{\infty}$ on $B\mathcal{G}$ subordinate to the cover $\left(\tau_{j}^{-1}(0,1]\right)_{j=0}^{\infty}$. We define a set map $\psi:B_{{\rm Hae}}\mathcal{G}\to X$ as follows: Let $b=\bigoplus_{i=0}^{\infty}t_{i}g_{i}\in B_{{\rm Hae}}\mathcal{G}$ be a point. Let $k_{0}<\dots<k_{n}$ be the enumeration of the integers $i\geq 0$ such that $t_{i}>0$. Then $\psi\left(b\right)$ is represented by the point
$\left(\left(\psi_{k_{0}}\left(b\right)t_{k_{0}},\dots,\psi_{k_{n}}\left(b\right)t_{k_{n}}\right),x_{0}\xrightarrow{g_{k_{1}}^{-1}g_{k_{0}}}\cdots\xrightarrow{g_{k_{n}}^{-1}g_{k_{n-1}}}x_{n},k_{0}<\dots<k_{n}\right),$
where $x_{i}$ denotes the domain of $g_{k_{i}}$. To check the continuity of $\psi$, choose a neighborhood $U$ of $b$ such that $\psi_{j}|U=0$ for all $j\in\mathbb{Z}_{\geq 0}\setminus\{k_{0},\dots,k_{n}\}$. (This is possible because $\operatorname{supp}\psi_{j}\subset\tau_{j}^{-1}\left((0,1]\right)$ and the supports of $\{\psi_{j}\}_{j}$ are locally finite.) Without loss of generality, we may assume that $\tau_{k_{0}},\dots,\tau_{k_{n}}$ are positive on $U$. Then on $U$, the map $\psi$ can be written as a composite
$U\to\left|\Delta^{k}\right|\times\mathcal{G}_{n}\times\{k_{0}<\cdots<k_{n}\}\to X,$
each of which is continuous. This proves that $\psi$ is continuous.
We claim that $\psi$ is a homotopy inverse of $\phi$. Define a map $h:X\times[0,1]\to X$ as follows: Let $x=[\left(t_{0},\dots,t_{n}\right),\sigma,k_{0}<\cdots<k_{n}]$ be a point of $X$ and let $0\leq s\leq 1$. Then $h\left(x,s\right)$ is represented by
$\left(\left(\left(\left(1-s\right)+s\psi_{k_{0}}\left(\phi\left(x\right)\right)\right)t_{0},\dots,\left(\left(1-s\right)+s\psi_{k_{0}}\left(\phi\left(x\right)\right)\right)t_{n}\right),\sigma,k_{0}<\dots<k_{n}\right).$
This defines a homotopy from $\operatorname{id}_{X}$ to $\psi\phi$. Likewise, there is a map $B_{{\rm Hae}}\mathcal{G}\times[0,1]\to B_{{\rm Hae}}\mathcal{G}$ given by
$\left(b,s\right)=\left(\bigoplus_{i\geq 0}t_{i}g_{i},s\right)\mapsto\bigoplus_{i\geq 0}\left(\left(1-s\right)+s\psi_{i}\left(b\right)\right)t_{i}g_{i},$
which is a homotopy from $\operatorname{id}_{B\mathcal{G}}$ to $\phi\psi$. The claim follows. ∎
For the next lemma, we write $\mathbf{\Delta}_{\mathrm{inj}}\subset\mathbf{\Delta}$ for the subcategory spanned by the injective poset maps. We also identify sets with discrete simplicial sets. Note that homotopy colimits over $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$ do not commute with products (up to weak equivalence) in general, as its nerve is not sifted.
Lemma A.15.
Let $\mathcal{C}$ be a category internal to $S\SS$. The map
$\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(\mathcal{C}\times N^{\mathrm{nd}}\left(\mathbb{Z}_{\geq 0}\right)\right)\to\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\mathcal{C}$
is a weak homotopy equivalence.
Remark A.16.
Note that in general homotopy colimits over $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$ do not commute with products even up to weak equivalence, as its nerve is not sifted. In contrast, the nerve of $\mathbf{\Delta}^{\mathrm{op}}$ is sifted [Lur09, Lemma 5.5.8.4].
Proof.
[…]
$\Phi\left(A\right)=\int^{[n]\in\mathbf{\Delta}_{\mathrm{inj}}}A_{n}\times\Delta^{n}.$
To see that $SS\Phi:\SS^{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\to\SS$ is actually a model of homotopy colimits, observe that its right adjoint $S\mapsto S^{\Delta^{\bullet}}$ is right Quillen for the Reedy model structure by inspection, and that it is naturally weakly equivalent to the diagonal functor. Since every semisimplicial simplicial set is Reedy cofibrant, this means that $\Phi$ is a model of homotopy colimits.
[…]
$\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(X_{n}\right)$ $\cong\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\operatorname{hocolim}_{[m]\in\mathbf{\Delta}^{\mathrm{op}}}\left(X_{n,m}\right)$
$\simeq\operatorname{hocolim}_{[m]\in\mathbf{\Delta}^{\mathrm{op}}}\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(X_{n,m}\right),$
which is natural in $X$. This implies that, given a map $f:X\to Y$ of semisimplicial simplicial sets, if the map
$\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}X_{n,m}\to\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}Y_{n,m}$
is a weak homotopy equivalence for every $m\geq 0$, then $\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}f$ is a weak homotopy equivalence. Substituting $\mathcal{C}\times N^{\mathrm{nd}}\left(\mathbb{Z}_{\geq 0}\right)$ for $X$ and $\mathcal{C}$ for $Y$, we are reduced to the case where each $\mathcal{C}_{n}$ is a (discrete simplicial) set, i.e., $\mathcal{C}=N\left(\mathsf{C}\right)$ for some ordinary category $\mathsf{C}$.
Let $\mathsf{C}^{\prime}\subset\mathsf{C}\times\mathbb{Z}_{\geq 0}$ denote the subcategory generated by the morphisms $\left(X,n\right)\to\left(Y,m\right)$ such that $n<m$. We can identify the semisimplicial set $N\left(\mathsf{C}\right)\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)$ with the restriction of the nerve $N\left(\mathsf{C}^{\prime}\right)$ to $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$. Under this identification, our goal is to show that the projection $\pi:\mathsf{C}^{\prime}\to\mathsf{C}$ gives a weak homotopy equivalence
$\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}N\left(\mathsf{C}^{\prime}\right)\to\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}N\left(\mathsf{C}\right).$
Since the inclusion $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}\hookrightarrow\mathbf{\Delta}^{\mathrm{op}}$ is homotopy final (being a right adjoint), we may replace the indexing category of homotopy colimits by $\mathbf{\Delta}^{\mathrm{op}}$. With our model of $\operatorname{hocolim}_{\mathbf{\Delta}^{\mathrm{op}}}$, the relevant map will be just the nerve of $\pi$. Thus, we are reduced to showing that $N\left(\pi\right)$ is a weak homotopy equivalence.
To prove that $N\left(\pi\right)$ is a weak homotopy equivalence, we factor $\pi$ as $\mathsf{C}^{\prime}\xrightarrow{\iota}\mathsf{C}\times\mathbb{Z}_{\geq 0}\xrightarrow{\pi^{\prime}}\mathsf{C}$, where $\iota$ is the inclusion and $\pi^{\prime}$ is the projection. Since $N\left(\pi^{\prime}\right)$ is a weak homotopy equivalence (as $N\left(\mathbb{Z}_{\geq 0}\right)$ is weakly contractible), it will suffice to show that $N\left(\iota\right)$ is a weak homotopy equivalence. We prove this by using Quillen’s Theorem A: We must show that, for each $\left(X,n\right)\in\mathsf{C}\times\mathbb{Z}_{\geq 0}$, the fiber product
$\mathsf{P}=\mathsf{C}^{\prime}\times_{\mathsf{C}\times\mathbb{Z}_{\geq 0}}\left(\mathsf{C}\times\mathbb{Z}_{\geq 0}\right)_{\left(X,n\right)/}$
has weakly contractible nerve. The inclusion $\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)/}\hookrightarrow\mathsf{P}$ is a left adjoint (with right adjoint given by $\left(Y,m\right)\mapsto\left(X,n\right)$ if $m=n$ and $\left(Y,m\right)\mapsto\left(Y,m\right)$ if $m>n$), so it suffices to show that $N\left(\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)/}\right)$ is weakly contractible. This is clear, as $\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)}$ has an initial object. The proof is now complete. ∎
Proof of Proposition A.13.
Consider the maps
$B_{{\rm MD}}\mathcal{G}=\left\|\mathcal{G}\right\|\xleftarrow{\phi^{\prime}}\left\|\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right\|\xrightarrow{\phi}B_{{\rm Hae}}\mathcal{G},$
where $\phi$ is the map of Lemma A.14 and $\phi^{\prime}$ is induced by the projection. The map $\phi$ is a weak homotopy equivalence by Lemma A.14. We will complete the proof by showing that $\phi^{\prime}$ is a weak homotopy equivalence.
[…]
In light of the discussion in the previous paragraph, we can identify $\phi^{\prime}$ with the map
$\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(\mathcal{G}\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)\right)\to\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\mathcal{G}.$
This is a weak homotopy equivalence by the Quillen equivalence between $\mathrm{Top}$ and $S\SS$ and Lemma A.15, and we are done. ∎
Appendix B Some Results on $\infty$-Categories
In this section, we record some general results on $\infty$-categories. Results in the first three subsections (B.1, B.2, and B.3) are well-known, but their references are hard to find. The content of the final subsection (B.4) is new, at least to the author’s knowledge.
B.1. Mapping Spaces of Arrow Categories
In this section, we give a proof of the following (folklore) result, and then use it to prove a few of its corollaries.
Proposition B.1.
Let $\mathcal{C}$ be an $\infty$-category. There is a pullback square
$\operatorname{Fun}(\Delta^{1},\mathcal{C})(f,g)$$\mathcal{C}(X_{0},Y_{0})$$\mathcal{C}(X_{1},Y_{1})$$\mathcal{C}(X_{0},Y_{1})$$\scriptstyle{g_{\ast}}$$\scriptstyle{f^{*}}$
in $\mathcal{S}$, natural in $\left(f:X_{0}\to X_{1},g:Y_{0}\to Y_{1}\right)\in\operatorname{Fun}\left(\Delta^{1},\mathcal{C}\right)^{\mathrm{op}}\times\operatorname{Fun}\left(\Delta^{1},\mathcal{C}\right)$.
Proof.
Recall that for an $\infty$-category $\mathcal{X}$, the twisted arrow construction produces a left fibration $\operatorname{Tw}\left(\mathcal{X}\right)\to\mathcal{X}$ classifying the hom-functor $\mathcal{X}\left(-,-\right):\mathcal{X}^{\mathrm{op}}\times\mathcal{X}\to\mathcal{S}$ [Lur25, Tag 03JF]. According to [AGS23, Theorem 4.4], the square
$\operatorname{Tw}(\operatorname{Fun}(\Delta^{1},\mathcal{C}))$$\operatorname{Fun}(\operatorname{Tw}(\Delta^{1}),\operatorname{Tw}(\mathcal{C}))$$\operatorname{Fun}(\Delta^{1},\mathcal{C})^{\mathrm{op}}\times\operatorname{Fun}(\Delta^{1},\mathcal{C})$$\operatorname{Fun}(\operatorname{Tw}(\Delta^{1}),\mathcal{C}^{\mathrm{op}}\times\mathcal{C})$}}
is homotopy cartesian. We now observe that $\operatorname{Tw}\left(\Delta^{1}\right)$ is the poset $00\to 01\leftarrow 11$, i.e., a walking cospan. Consequently, the claim follows from the following assertion:
  • ($\ast$)
    Let $p:\mathcal{E}\to\mathcal{B}$ be a left fibration classified by a functor $f:\mathcal{B}\to\mathcal{S}$, and let $\mathcal{X}$ be an $\infty$-category. The left fibration $\operatorname{Fun}\left(\mathcal{X},p\right)$ is classified by the composite $\operatorname{Fun}\left(\mathcal{X},\mathcal{B}\right)\xrightarrow{\operatorname{Fun}\left(\mathcal{X},f\right)}\operatorname{Fun}\left(\mathcal{X},\mathcal{S}\right)\xrightarrow{\operatorname{lim}}\mathcal{S}$.
To prove ($\ast$), it suffices to consider the case where $p=p_{\mathrm{univ}}:\mathcal{S}_{\ast}\to\mathcal{S}$ is the universal left fibration. In this case, consider the following diagram
$\operatorname{Fun}(\mathcal{X},\mathcal{S}_{\ast})$$\operatorname{Fun}(\mathcal{X}^{\triangleleft},\mathcal{S}_{\ast})$$\mathcal{S}_{\ast}$$\operatorname{Fun}(\mathcal{X},\mathcal{S})$$\operatorname{Fun}(\mathcal{X}^{\triangleleft},\mathcal{S})$$\mathcal{S}.$$\scriptstyle{\mathrm{Ran}}$$\scriptstyle{\mathrm{ev}_{\infty}}$$\scriptstyle{p_{\mathrm{univ}}}$$\scriptstyle{\mathrm{Ran}}$$\scriptstyle{\mathrm{ev}_{\infty}}$
The left-hand square commutes up to natural equivalence because $p_{\mathrm{univ}}$ preserves limits. Combining this with the fact that the right Kan extension functors are fully faithful [Lur09, Proposition 4.3.2.15], we find that the left-hand square is a pullback square of possibly large $\infty$-categories. Also, since the inclusion $\{\infty\}\hookrightarrow\mathcal{X}^{\triangleleft}$ is left anodyne and $p_{\mathrm{univ}}$ is a left fibration, [Lur09, Corollary 2.1.2.7] implies that the right-hand square is also a pullback. The proof is now complete, as the bottom horizontal composite is nothing but the limit functor. ∎
Corollary B.2.
Let $\mathcal{C}$ be an $\infty$-category. For each object $X\in\mathcal{C}$, there is a pullback square
$\mathcal{C}_{/X}(a,b)$$\mathcal{C}(A,B)$$\ast$$\mathcal{C}(A,X)$$\scriptstyle{b_{\ast}}$$\scriptstyle{a}$
in $\mathcal{S}$, natural in $\left(a:A\to X,b:B\to X\right)\in\left(\mathcal{C}_{/X}\right)^{\mathrm{op}}\times\mathcal{C}_{/X}$.
Proof.
This follows from Proposition B.1 and the equivalence $\mathcal{C}_{/X}\xrightarrow{\simeq}\{X\}\times_{\operatorname{Fun}\left(\{0\},\mathcal{C}\right)}\operatorname{Fun}\left(\Delta^{1},\mathcal{C}\right)$ of [Lur09, Proposition 4.2.1.5]. ∎
We will need the following two results in the main body of the paper:
Proposition B.3.
Let $\mathcal{C}$ be an $\infty$-category, let $\mathcal{C}_{0}\subset\mathcal{C}$ be a full subcategory, let $X\in\mathcal{C}$ be an object, and let $\mathcal{I}$ be a small category. Let $f:\mathcal{I}^{\triangleright}\to\mathcal{C}$ be a functor which carries $\mathcal{I}$ into $\mathcal{C}_{0}$ and the cone point to $X$. The following conditions are equivalent:
  1. (1)
    The functor $\mathcal{I}\to\mathcal{C}_{/X}^{0}$ is final.
  2. (2)
    For each object $C\in\mathcal{C}_{0}$ admitting a morphism $\phi:C\to X$ in $\mathcal{C}$, the map
    $\operatorname{colim}_{I\in\mathcal{I}}\mathcal{C}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
    is an equivalence of $\infty$-groupoids.
Proof.
Condition (1) is equivalent to the condition that, for each object $\left(c:C\to X\right)\in\mathcal{C}_{/X}^{0}$, the $\infty$-category $\left(\mathcal{C}_{/X}^{0}\right)_{c/}\times_{\mathcal{C}^{0}}\mathcal{I}$ is weakly contractible. By [Lur09, Corollary 3.3.4.6] and [Lur09, Theorem 4.2.4.1], the $\infty$-groupoid $\left(\mathcal{C}_{/X}^{0}\right)_{c/}\times_{\mathcal{C}^{0}}\mathcal{I}$ has the homotopy type of $\operatorname{colim}_{I\in\mathcal{I}}\left(\mathcal{C}_{/X}^{0}\right)_{c/}\left(c,f\left(I\right)\right)\in\mathcal{S}$. Since colimits in $\mathcal{S}$ are universal [Lur09, Lemma 6.1.3.14], it follows from Corollary B.2 that $\operatorname{colim}_{I\in\mathcal{I}}\mathcal{C}_{/X}\left(c,f\left(I\right)\right)$ has the weak homotopy type of the fiber of the map
$\operatorname{colim}_{I\in\mathcal{I}}\mathcal{C}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
over $\alpha$. The claim follows. ∎
For later reference, we record a variant of Proposition B.3.
Proposition B.4.
Let $\mathcal{C}$ be an $\infty$-category, let $\mathcal{C}_{0}\subset\mathcal{C}$ be a full subcategory, let $X\in\mathcal{C}$ be an object, and let $\mathcal{I}$ be an $\infty$-category. Let $f:\mathcal{I}^{\triangleright}\to\mathcal{C}$ be a functor which carries $\mathcal{I}$ into $\mathcal{C}_{0}^{\simeq}$ and the cone point to $X$. The following conditions are equivalent:
  1. (1)
    The functor $f^{\prime}:\mathcal{I}\to\left(\left(\mathcal{C}_{0}\right)_{/X}\right)^{\simeq}$ which is adjoint to $f$ is a weak homotopy equivalence.
  2. (2)
    For each object $C\in\mathcal{C}_{0}$ which admits a morphism $C\to X$ in $\mathcal{C}$, the map
    $\operatorname{colim}_{I\in\mathcal{I}}\mathcal{C}^{\simeq}\left(C,f\left(I\right)\right)\to\mathcal{C}\left(C,X\right)$
    is an equivalence of $\infty$-groupoids.
The proof of Proposition B.4 is nearly identical to that of Proposition B.3, noting that for a Kan complex $X$, a map $K\to X$ of simplicial sets is final if and only if it is a weak homotopy equivalence. Details are left to the reader.
B.2. Kan Extension
In this subsection, we summarize a few key facts on Kan extensions. The starting point is the following lemma.
Lemma B.5.
[Lur25, Tag 02ZM] Let $\mathcal{C},\mathcal{D},$ and $\mathcal{E}$ be $\infty$-categories. Let $p:\mathcal{C}\to\mathcal{D}$, $f:\mathcal{C}\to\mathcal{E}$, and $l:\mathcal{D}\to\mathcal{E}$ be functors, and let $\alpha:f\to lp$ be a natural transformation. Suppose that $p$ is a cocartesian fibration. The following conditions are equivalent:
  1. (1)
    The natural transformation $\alpha$ exhibits $l$ as a left Kan extension of $f$ along $p$.
  2. (2)
    For each object $D\in\mathcal{D}$, the map $\alpha$ restricts to a colimit diagram $f|\mathcal{C}_{D}\to\underline{lp\left(D\right)}$, where we set $\mathcal{C}_{D}=\mathcal{C}\times_{\mathcal{D}}\{D\}$.
Remark B.6.
Informally, Lemma B.5 says that a left Kan extension $\operatorname{Lan}_{p}f$ of $f$ along $p$, if it exists, is given by the formula
$\operatorname{Lan}_{p}f\left(D\right)\simeq\operatorname{colim}_{C\in\mathcal{C}_{D}}f\left(C\right).$
This informal notation is quite useful, and we will use it frequently. Thus, for instance, if we say “define a functor $g$ by $g\left(D\right)=\operatorname{colim}_{C\in\mathcal{C}_{D}}f\left(C\right)$,” what we really mean is that “we define $g$ to be a left Kan extension of $f$ along pp.” (In the applications we have in mind, $\mathcal{D}$ will be the nerve of an ordinary category $\mathcal{D}_{0}$, and $p$ will be the relative nerve of an ordinary functor $S\mathcal{D}_{0}\to\SS$.)
Lemma B.5 has the following consequences.
Proposition B.7.
Let
$\mathcal{A}$$\mathcal{B}$$\mathcal{C}$$\scriptstyle{f}$$\scriptstyle{q}$$\scriptstyle{p}$
be a commutative diagram of $\infty$-categories. Suppose that $p$ and $q$ are cocartesian fibrations and that, for each object $C\in\mathcal{C}$, the map
$\mathcal{A}\times_{\mathcal{C}}\{C\}=\mathcal{A}_{C}\to\mathcal{B}_{C}$
induced by $f$ is final. Then $f$ is final.
Proof.
We wish to show that, for each object $B\in\mathcal{B}$, the $\infty$-category $\mathcal{A}_{B/}$ is weakly contractible. According to [Lur09, Corollary 3.3.4.6], the simplicial set $\mathcal{A}_{B/}$ has the weak homotopy type of the colimit of the diagram $\mathcal{A}\xrightarrow{f}\mathcal{B}\xrightarrow{F_{B}}\mathcal{S}$, where $F_{B}$ is the functor corepresented by $B$. By the same token, $\mathcal{B}_{B/}$ has the weak homotopy type of the colimit of $F_{B}$. Therefore, to prove that $\mathcal{A}_{B/}$ is weakly contractible, it suffices to prove that precomposing $f$ does not change the colimit of any diagram $\mathcal{B}\to\mathcal{S}$. In other words, it suffices to show that, for each diagram $g:\mathcal{B}\to\mathcal{S}$, the map
$\mathcal{S}^{g/}\to\mathcal{S}^{gf/}$
is a covariant equivalence over $\mathcal{S}$.
Find a functor $l:\mathcal{C}\to\mathcal{S}$ and a natural transformation $\alpha:g\to lq$ which exhibits $l$ as a left Kan extension of $g$ along $q$, and find also a Kan complex $K$ and a natural transformation $\beta:l\to\underline{K}$ which exhibits $K$ as a colimit of $l$. Then a composite natural transformation
$\theta:g\xrightarrow{\alpha}lq\xrightarrow{\beta q}\underline{K}$
exhibits $K$ as a colimit of $g$ (by the transitivity of Kan extensions[Lur25, Tag 031M]), and we wish to show that the natural transformation $\theta f:gf\to\underline{K}$ exhibits $K$ as a colimit of $gf$. For this, it suffices to show that the natural transformation $\alpha f:gf\to lp$ exhibits $l$ as a left Kan extension of $gf$ along $p$. By Lemma B.5, we must show that, for each $C\in\mathcal{C}$, the natural transformation
$\alpha f|\mathcal{A}_{C}:gf|\mathcal{A}_{C}\to\underline{l\left(C\right)}$
exhibits $l\left(C\right)$ as a colimit of $gf|\mathcal{A}_{C}$. Since $f|\mathcal{A}_{C}:\mathcal{A}_{C}\to\mathcal{B}_{C}$ is final, we are reduced to showing that the natural transformation
$\alpha|\mathcal{B}_{C}:g|\mathcal{B}_{C}\to\underline{l\left(C\right)}$
is a colimit diagram. This follows from Lemma B.5. ∎
Proposition B.8.
Let $\overline{p}:\overline{\mathcal{E}}\to\mathcal{B}^{\triangleright}$ be a cocartesian fibration of $\infty$-categories classifying a colimit diagram in $\mathcal{B}^{\triangleright}\to\mathcal{C}\mathsf{at}_{\infty}$, and let $f:\overline{\mathcal{E}}\to\mathcal{C}$ be a diagram that carries $\overline{p}$-cocartesian edges to equivalences. If $f$ admits a left Kan extension $F:\mathcal{B}^{\triangleright}\to\mathcal{C}$ along $p$, then $F$ is a colimit diagram.
Remark B.9.
In the situation of Proposition B.8, the functor $F$ is given by $F\left(B\right)=\operatorname{colim}_{E\in\mathcal{E}_{B}}f\left(E\right)$ (Lemma B.5). Thus, informally, the proposition says that
$\operatorname{colim}_{E\in\operatorname{colim}_{B\in\mathcal{B}}\mathcal{E}_{B}}f\left(E\right)\simeq\operatorname{colim}_{B\in\mathcal{B}}\operatorname{colim}_{E\in\mathcal{E}_{B}}f\left(E\right).$
Remark B.10.
There is an obvious dual version of Proposition B.8: Let $\overline{p}:\overline{\mathcal{E}}\to\left(\mathcal{B}^{\triangleright}\right)^{\mathrm{op}}$ is a cartesian fibration of $\infty$-categories classifying a colimit diagram $\mathcal{B}^{\triangleright}\to\mathcal{C}\mathsf{at}_{\infty}$, and let $f:\overline{\mathcal{E}}\to\mathcal{C}$ be a functor carrying $\overline{p}$-cartesian edges to equivalences. If $f$ has a left Kan extension $F:\left(\mathcal{B}^{\triangleright}\right)^{\mathrm{op}}\to\mathcal{C}$ along $p$, then $F$ is a limit diagram.
Proof of Proposition B.8.
Consider the following diagram and natural transformations:
$\mathcal{E}$$\overline{\mathcal{E}}$$\mathcal{C}$$\mathcal{B}$$\mathcal{B}^{\triangleright}$$\{\infty\}$$\scriptstyle{j}$$\scriptstyle{p}$$\scriptstyle{f}$$\scriptstyle{\overline{p}}$$\scriptstyle{i}$$\scriptstyle{F}$$\scriptstyle{k}$$\scriptstyle{\alpha}$$\scriptstyle{\beta}$
Here the maps $i,j,k$ are inclusions, $p$ is the pullback of $\overline{p}$ along $i$, and $\alpha$ is the structure natural transformation for a left Kan extension. Our goal is to show that $F\beta$ is a colimit diagram. Since $\alpha j$ exhibits $Fi$ as a left Kan extension of $fj$ along $p$ (by Lemma B.5), we only need to show that the composite $2$-cell
$\theta:fj\xrightarrow{\alpha j}F\overline{p}j=Fip\xrightarrow{F\beta p}\underline{F\left(\infty\right)}$
is a colimit diagram.
Let $j_{\infty}:\overline{\mathcal{E}}_{\infty}\to\overline{\mathcal{E}}$ denote the inclusion, and let $\gamma:j\to j_{\infty}r$ denote the cocartesian natural transformation covering $\beta$ (i.e., $\beta p=\overline{p}\gamma$), where $r:\mathcal{E}\to\overline{\mathcal{E}}_{\infty}$ is the codomain restriction of $\gamma|\mathcal{E}\times\{1\}$. The map $\theta$ is then homotopic to the pasting of the $2$-cells
$\overline{\mathcal{E}_{\infty}}$$\mathcal{E}$$\overline{\mathcal{E}}$$\mathcal{C}$$\mathcal{B}$$\mathcal{B}^{\triangleright}$$\scriptstyle{j_{\infty}}$$\scriptstyle{r}$$\scriptstyle{j}$$\scriptstyle{p}$$\scriptstyle{f}$$\scriptstyle{\overline{p}}$$\scriptstyle{i}$$\scriptstyle{F}$$\scriptstyle{\gamma}$$\scriptstyle{\alpha}$
which can be written as a composite
$fj\xrightarrow{f\gamma}fj_{\infty}r\xrightarrow{\alpha j_{\infty}r}F\overline{p}j_{\infty}r.$
Since $f$ carries cocartesian morphisms to equivalences, the map $f\gamma$ is an equivalence. Therefore, it will suffice to show that $\alpha j_{\infty}r$ is a colimit diagram. Since $\overline{p}$ classifies a colimit diagram, [Lur09, Lemma 3.3.4.1 and Proposition 3.3.4.2] show that the map $r$ is a localization at $p$-cocartesian morphisms (in the sense of Definition A.1). Since localizations are final [Lur25, Tag 02N9], we are reduced to showing that $\alpha j_{\infty}$ is a colimit diagram. This follows from Lemma B.5, and we are done. ∎
B.3. t-Structure and Homotopy Groups
Let $\mathcal{C}$ be a stable $\infty$-category. Recall that a (homological) t-structure on $\mathcal{C}$ is a collection of full subcategories $\{\mathcal{C}_{\geq n}\}_{n\in\mathbb{Z}}$ and $\{\mathcal{C}_{\leq n}\}_{n\in\mathbb{Z}}$ of $\mathcal{C}$ that determine the ordinary t-structure on the triangulated category $\operatorname{ho}\left(\mathcal{C}\right)$ [Lur17, 1.2.1]. If $\mathcal{C}$ is equipped with a t-structure, its heart $\mathcal{C}^{\heartsuit}=\mathcal{C}_{\leq 0}\cap\mathcal{C}_{\geq 0}$ is equivalent to its homotopy category, which is an abelian category. For each $n\geq 0$, there is an $n$th homotopy group functor $\pi_{n}:\mathcal{C}\to\mathcal{C}^{\heartsuit}$ [Lur17, Definition 1.2.1.11], and every fiber sequence $X\to Y\to Z$ in $\mathcal{C}$ gives rise to a long exact sequence
$\cdots\to\pi_{n+1}\left(Z\right)\to\pi_{n}\left(X\right)\to\pi_{n}\left(Y\right)\to\pi_{n}\left(Z\right)\to\pi_{n-1}\left(X\right)\to\cdots$
in $\mathcal{C}^{\heartsuit}$ [GM03, IV.4. Theorem 11]. In this subsection, we record some results on homotopy groups that we will use in the main body of the paper.
The datum of a t-structure give us a measurement of the failure of a cube to be cartesian. The following sequences of results, up to Corollary B.15, give an estimate of this sort using estimates of smaller cubes. These results and arguments are due to Goodwillie, but we record them for completeness.
Definition B.11.
Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, and let $k$ be an integer. An object $X\in\mathcal{C}$ is said to be $k$-connected if it belongs to $\mathcal{C}_{\geq k+1}$, or equivalently, $\pi_{i}\left(X\right)\simeq 0$ for all $i\leq k$. A map $f:X\to Y$ in $\mathcal{C}$ is said to be $k$-connected if its fiber is $\left(k-1\right)$-connected. Given a finite set $S$ and an $S$-cube $X$ in $\mathcal{C}$, we say that $X$ is $k$-cartesian if the map
$X\left(\emptyset\right)\to\operatorname{lim}\left(X|\mathcal{P}_{0}\left(S\right)\right)$
is $k$-connected.
Proposition B.12.
[Goo92, Proposition 1.5] Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $k$ be an integer, and let $X\xrightarrow{f}Y\xrightarrow{g}Z$ be a pair of maps in $\mathcal{C}$.
  1. (1)
    If $f$ and $g$ are $k$-connected, so is $gf$.
  2. (2)
    If $g$ is $\left(k+1\right)$-connected and $gf$ is $k$-connected, then $f$ is $k$-connected.
  3. (3)
    If $gf$ is $k$-connected and $f$ is $\left(k-1\right)$-connected, then $g$ is $k$-connected.
Proof.
This follows from the long exact sequence associated with the fiber sequence $\mathrm{fib}\left(f\right)\to\mathrm{fib}\left(gf\right)\to\mathrm{fib}\left(g\right)$. ∎
Proposition B.13.
[Goo92, Proposition 1.6] Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $S$ be a finite set, and let $F$ be an $\left(S\amalg\{\ast\}\right)$-cube in $\mathcal{C}$. Define $S$-cubes $X,Y$ by $X=F|\mathcal{P}\left(S\right)$ and $Y=F\left(-\cup\{\ast\}\right)$.
  1. (i)
    If $Y$ is $k$-cartesian and $F$ is $k$-cartesian, then $X$ is $k$-cartesian.
  2. (ii)
    If $X$ is $k$-cartesian and $Y$ is $\left(k+1\right)$-cartesian, then $F$ is $k$-cartesian.
Proof.
We will apply Proposition B.12 to the maps
$X\left(\emptyset\right)\xrightarrow{f}\operatorname{lim}\left(F|\mathcal{P}_{0}\left(S\amalg\{\ast\}\right)\right)\xrightarrow{g}\operatorname{lim}\left(X|\mathcal{P}_{0}\left(S\right)\right).$
To obtain the conclusion, it will suffice to show that $g$ is a pullback of the map $Y\left(\emptyset\right)\to\operatorname{lim}\left(Y|\mathcal{P}_{0}\left(S\right)\right)$. This follows from Remark B.10, as the square
$(\mathcal{P}(S)\times[1])\setminus\{(\emptyset,0)\}$$\mathcal{P}_{0}(S)\times[1]$$\mathcal{P}(S)\times\{1\}$$\mathcal{P}_{0}(S)\times\{1\}$}}
in a pushout in $\mathcal{C}\mathsf{at}_{\infty}$ (because the Joyal model structure is left proper). ∎
Proposition B.14.
[Goo92, Proposition 1.20] Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $S$ and $T$ be finite sets, and let $X:\mathcal{P}\left(S\amalg T\right)\to\mathcal{D}$ be an $S\amalg T$-cube. Suppose that:
  1. (1)
    $X$ is $k$-cartesian.
  2. (2)
    For each $U\in\mathcal{P}_{0}\left(S\right)$, the $T$-cube $X\left(U\amalg-\right)$ is $\left(k+\left|U\right|-1\right)$-cartesian.
Then the $T$-cube $X\left(\emptyset\amalg-\right)$ is $k$-cartesian.
Proof.
We prove the claim by induction on $n=\left|S\right|$. If $n=0$, there is nothing to prove. For the inductive step, fix an element $\ast\in S$, and set $S^{\prime}=S\setminus\{\ast\}$. Applying the induction hypothesis to the $S^{\prime}\amalg\left(\{\ast\}\amalg T\right)$-cube $X$, we deduce that the $\left(\{\ast\}\amalg T\right)$-cube $X\left(\emptyset\amalg-\amalg-\right)$ is $k$-cartesian. Also, by condition (2), the $T$-cube $X\left(\emptyset\amalg\{\ast\}\amalg-\right)$ is $k$-cartesian. It follows from part (i) of Proposition B.13 that the $T$-cube $X\left(\emptyset\amalg\emptyset\amalg-\right)$ is $k$-cartesian, as desired. ∎
Corollary B.15.
[Goo92, Proposition 1.22] Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure, let $S$ be a finite set, and let $X:\mathcal{P}_{0}\left(S\right)\times[1]\to\mathcal{D}$ be a diagram. Suppose that, for each $U\in\mathcal{P}_{0}\left(S\right)$, the map
$X\left(U,0\right)\to X\left(U,1\right)$
is $k_{U}$-connected. Then the map
$\operatorname{lim}_{U\in\mathcal{P}_{0}\left(S\right)}X\left(U,0\right)\to\operatorname{lim}_{U\in\mathcal{P}_{0}\left(S\right)}X\left(U,1\right)$
is $\min\{1-\left|U\right|+k_{U}\mid U\in\mathcal{P}_{0}\left(S\right)\}$-connected.
Proof.
Let $\overline{X}:\mathcal{P}\left(S\right)\times[1]\to\mathcal{D}$ denote the right Kan extension of $X$. Since the inclusion $\mathcal{P}\left(S\right)\to\mathcal{P}\left(S\amalg\{\ast\}\right)$ is initial, the $\left(S\amalg\{\ast\}\right)$-cube $\overline{X}$ is cartesian. So the claim follows from applying Proposition B.14 to $\overline{X}$. ∎
We conclude this subsection with two more miscellaneous results.
Lemma B.16 (Milnor exact sequence).
Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure. Suppose that $\mathcal{C}$ has countable products and that the functor $\pi_{0}:\mathcal{C}\to\mathcal{C}^{\heartsuit}$ preserves them. Given a diagram $X_{\bullet}:\mathbb{Z}\to\mathcal{C}$ depicted as
$\cdots\xrightarrow{p_{2}}X_{2}\xrightarrow{p_{1}}X_{1}\xrightarrow{p_{0}}X_{0}\xrightarrow{p_{-1}}\cdots,$
there is an exact sequence
$0\to\operatorname{lim}_{i}^{1}\pi_{n+1}\left(X_{i}\right)\to\pi_{n}\left(\operatorname{lim}_{i}X_{i}\right)\to\operatorname{lim}_{i}\pi_{n}\left(X_{i}\right)\to 0$
in the abelian category $\mathcal{C}^{\heartsuit}\simeq\operatorname{ho}\left(\mathcal{C}^{\heartsuit}\right)$, which is natural in $X_{\bullet}$.
Proof.
Set $X=\operatorname{lim}_{i}X_{i}$. Since $\pi_{0}$ carries fiber sequences to exact sequences [GM03, IV.4., Theorem 11], it suffices to prove the following two assertions:
  1. (1)
    $X$ is naturally equivalent to equalizer of the maps $\operatorname{id},p:\prod_{i}X_{i}\rightrightarrows\prod_{i}X_{i}$, where $p$ is determined by the maps $\{p_{i}\}_{i\geq 0}$.
  2. (2)
    Given a pair of maps $f,g:A\rightrightarrows B$ in $\mathcal{C}$, the equalizer of $f$ and $g$ is equivalent to naturally equivalent to the kernel of fgf-g.
Indeed, assertions (1) and (2) will give us a fiber sequence $X\to\prod_{i}X_{i}\xrightarrow{\operatorname{id}-p}\prod_{i}X_{i}$, and then the associated long exact sequence of homotopy groups gives us the desired conclusion.
To prove (1), let $\mathcal{B}=\{0\mathrel{{\mathop{\rightrightarrows}\limits_{b}^{a}}}1\}$ denote (the nerve of) the free walking parallel arrows, and let $\mathcal{E}$ denote the Grothendieck construction of the functor $\mathcal{B}\to\mathsf{Cat}$ corresponding to the maps $\operatorname{id},-1:\mathbb{Z}\rightrightarrows\mathbb{Z}$. Thus we can depict $\mathcal{E}$ as as
[…]
and there is a projection $\mathcal{E}\to\mathcal{B}$ carrying $a_{i}$ to $a$ and $b_{i}$ to $b$. The assignment $\left(n,i\right)\mapsto n$ determines a functor $\pi:\mathcal{E}\to\mathbb{Z}$. Using the fact that the classifying space of $[1]$ is contractible, we deduce that $\pi$ is a localization at the maps $\{a_{i}\}_{i\geq 0}$. In particular, $\pi$ is initial [Lur25, Tag 02N9], so
$\operatorname{lim}_{i}X_{i}\simeq\operatorname{lim}\left(\mathcal{E}\xrightarrow{\pi}\mathbb{Z}\xrightarrow{X_{\bullet}}\mathcal{C}\right).$
The right-hand limit is equivalent to the limit of the right Kan extension of $X_{\bullet}\circ\pi$ along the projection $\mathcal{E}\to\mathcal{B}$, which, by Remark B.10, is the equalizer of $\operatorname{id}$ and $p$.
Next, for (2), we consider the pullback squares
[…]
The pullback $P$ is nothing but the equalizer of $f$ and $g$ [Lur25, Tag 03HC], and the pullback of the outer rectangle is the kernel of fgf-g. The proof is now complete. ∎
Lemma B.17.
Let $\mathcal{C}$ be a stable $\infty$-category with a t-structure. Suppose that $\mathcal{C}$ has small limits and that 0-connected objects are stable under small products. Suppose we are given a pushout diagram
[…]
in $\mathcal{S}$, where $I$ is a set and $d\geq 0$. For every diagram $\phi:Y\to\mathcal{C}$ carrying each object to a 0-connected object, the map
$\theta:\operatorname{lim}\phi\to\operatorname{lim}\left(\phi|X\right)$
is $\left(-d+1\right)$-connected.
Proof.
By Remark B.10, the map $\theta$ is a pullback of the map
$\prod_{i\in I}\operatorname{lim}\left(\phi|\left(D^{d}\times\{i\}\right)\right)\to\prod_{i\in I}\operatorname{lim}\left(\phi|\left(S^{d-1}\times\{i\}\right)\right).$
Therefore, it suffices to show that the limit $\operatorname{lim}\left(\phi|\left(S^{d-1}\times\{i\}\right)\right)$ is $\left(-d+1\right)$-connected for each $i\in I$. When $d=0$, the limit is a zero object, so there is nothing to prove. If $d\geq 1$, let $C\in\mathcal{C}$ denote an image of a vertex of $S^{d-1}$. Since $D^{d}$ is contractible, the diagram $\phi|S^{d-1}\times\{i\}$ is equivalent to the constant diagram at $C$. Thus
$\operatorname{lim}\phi|\left(S^{d-1}\times\{i\}\right)\simeq C^{S^{d-1}}\simeq C^{\Sigma^{d-1}S^{0}}\simeq\Omega^{d-1}\left(C^{S^{0}}\right)\simeq\Omega^{d-1}\left(C\times C\right).$
The right-hand side is $\left(-d+1\right)$-connected, so we are done. ∎
B.4. Extension by Zero Objects
Consider the following problem: Let $\mathcal{C}$ be an $\infty$-category and let $\mathcal{C}_{0},\mathcal{C}_{1}\subset\mathcal{C}$ be full subcategories such that every object of $\mathcal{C}$ belongs to exactly one of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$. Let $\mathcal{D}$ be a pointed $\infty$-category and suppose we are given a functor $f:\mathcal{C}\to\mathcal{D}$ carrying each object of $\mathcal{C}_{0}$ to a zero object. One would expect that $f$ is equivalent to a smaller piece of data. In other words, we would expect that there is a categorical equivalence
$\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\simeq\{\text{something smaller}\}$
where $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\subset\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)$ denotes the full subcategory spanned by the objects carrying each object in $\mathcal{C}_{0}$ to a zero object. We want to figure out what the smaller thing is.
A naïve guess of the “smaller thing” is $\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$, but this is too naïve. Indeed, there could be a morphism in $\mathcal{C}_{1}$, say $\alpha$, which factors through a morphism of $\mathcal{C}_{0}$, and there is no reason why an arbitrary functor $\mathcal{C}_{1}\to\mathcal{D}$ carries $\alpha$ to a zero map. So the restriction $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$ is, in general, not essentially surjective.
The next guess is that the essential image of the restriction functor $\operatorname{Fun}^{\prime}\left(\mathcal{C},\mathcal{D}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1},\mathcal{D}\right)$ will suffice. If $\mathcal{C}$ and $\mathcal{D}$ are nerves of ordinary categories, this is true. But for arbitrary $\infty$-categories, this is false; the problem is that there could be a nontrivial homotopy between zero maps, as one can see from the particular case where $\mathcal{C}=\Delta^{2}$ and $\mathcal{C}_{1}=\{1\}$.
In this subsection, we will give one answer to this question in the case where $\mathcal{C}_{1}$ is “isolated” in $\mathcal{C}$.
To state our main result, we need a bit of terminology.
Definition B.18.
Let $\mathcal{C}$ be an $\infty$-category, and let $\mathcal{C}^{\prime}\subset\mathcal{C}$ be a full subcategory. A morphism $X\to Y$ of $\mathcal{C}^{\prime}$ is said to be isolated in $\mathcal{C}$ if it does not factor through any object of $\mathcal{C}$ which lies outside of $\mathcal{C}^{\prime}$. We say that $\mathcal{C}^{\prime}$ is isolated in $\mathcal{C}$ if the morphisms of $\mathcal{C}^{\prime}$ that are isolated in $\mathcal{C}$ are closed under compositions.
Here is the main result of this subsection.
Proposition B.19.
Let $\mathcal{C}$ be an $\infty$-category, and let $\mathcal{C}_{0},\mathcal{C}_{1}\subset\mathcal{C}$ be full subcategories such that every object of $\mathcal{C}$ belongs to exactly one of $\mathcal{C}_{0}$ or $\mathcal{C}_{1}$. Let $\mathcal{D}$ be a pointed $\infty$-category, and let $\mathcal{Z}\subset\mathcal{D}$ denote the full subcategory spanned by the zero objects. Suppose that $\mathcal{C}_{1}$ is isolated in $\mathcal{C}$, and let $\mathcal{C}_{1}^{{\rm isltd}}\subset\mathcal{C}_{1}$ denote the subcategory spanned by the morphisms isolated in $\mathcal{C}$. The functor
$\theta:\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{Z}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1}^{{\rm isltd}},\mathcal{D}\right)$
is a trivial fibration.
Proof.
Let us say that a morphism $f$ of $\mathcal{C}$ is basic if either $f$ is a morphism of $\mathcal{C}_{1}^{{\rm isltd}}$ or $f$ is not a morphism of $\mathcal{C}^{1}$. We will say that a simplex $\sigma:\Delta^{d}\to\mathcal{C}$ is basic if for each $0\leq i<d$, the restriction $\sigma|\Delta^{\{i,i+1\}}$ is basic. We let $\{\sigma_{\alpha}:\Delta^{d_{\alpha}}\to\mathcal{C}\}_{\alpha\in A}$ denote the set of all basic simplices of $\mathcal{C}$. Note that the map $\coprod_{\alpha\in A}\Delta^{d_{\alpha}}\to\mathcal{C}$ is an epimorphism of simplicial sets because every simplex of $\mathcal{C}$ is a face of a basic simplex.
For each $\alpha\in A$, set $P_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{0}$, $Q_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{1}$, and $Q^{\prime}_{\alpha}=\Delta^{d_{\alpha}}\times_{\mathcal{C}}\mathcal{C}_{1}^{{\rm isltd}}$. Since $\theta$ is a pullback of the map
$\operatorname{Fun}\left(\coprod_{\alpha\in A}\Delta^{d_{\alpha}},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\coprod_{\alpha\in A}P_{\alpha},\mathcal{D}\right)}\operatorname{Fun}\left(\coprod_{\alpha\in A}P_{\alpha},\mathcal{Z}\right)\to\operatorname{Fun}\left(\coprod_{\alpha\in A}Q_{\alpha}^{\prime},\mathcal{D}\right),$
it suffices to show that for each $\alpha\in A$, the map
$\theta_{\alpha}:\operatorname{Fun}\left(\Delta^{d_{\alpha}},\mathcal{D}\right)\times_{\operatorname{Fun}\left(P_{\alpha},\mathcal{D}\right)}\operatorname{Fun}\left(P_{\alpha},\mathcal{Z}\right)\to\operatorname{Fun}\left(Q^{\prime}_{\alpha},\mathcal{D}\right)$
is a trivial fibration. Now since $\sigma_{\alpha}$ is isolated, the subcategory $Q^{\prime}_{\alpha}\subset Q_{\alpha}$ is spanned by the morphisms of $Q_{\alpha}$ that are isolated in $\Delta^{d_{\alpha}}$. Therefore, we are reduced to the case where $\mathcal{C}=\Delta^{d}$ for some $d\geq 0$. In this case, we will prove the assertion by induction on $d$.
If $d=0$, the claim is obvious. For the inductive step, suppose we have proved the assertion up to d1d-1. If $\mathcal{C}_{0}$ is empty, the claim is trivial, so assume that $\mathcal{C}_{0}$ is nonempty. Let $m$ be the maximal integer which belongs to $\mathcal{C}_{0}$. There are two cases to consider.
Suppose first that $m=d$. In this case, [Lur09, Proposition 4.3.2.15] shows that the functor
$\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{Z}\right)\to\operatorname{Fun}\left(\Delta^{d-1},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{Z}\right)$
is a trivial fibration. By the induction hypothesis, the functor
$\operatorname{Fun}\left(\mathcal{C}\cap\Delta^{d-1},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0}\cap\Delta^{d-1},\mathcal{Z}\right)\to\operatorname{Fun}\left(\mathcal{C}_{1}^{{\rm isltd}},\mathcal{D}\right)$
is a trivial fibration. It follows that $\theta$ is a composition of trivial fibrations and hence is itself a trivial fibration.
Suppose next that $m<d$. Set $X=\Delta^{\{0,\dots,m\}}\amalg_{\{m\}}\Delta^{\{m\dots,d\}}$. The map $\theta$ factors as
$\operatorname{Fun}\left(\mathcal{C},\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C},\mathcal{Z}\right)$ $\xrightarrow{f}\operatorname{Fun}\left(X,\mathcal{D}\right)\times_{\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{D}\right)}\operatorname{Fun}\left(\mathcal{C}_{0},\mathcal{Z}\right)$
$\xrightarrow{g}\operatorname{Fun}\left(\mathcal{C}_{1}^{{\rm isltd}},\mathcal{D}\right).$
The map $f$ is a trivial fibration because the inclusion $X\to\mathcal{C}$ is a weak categorical equivalence [Lur09, Lemma 5.4.5.10]. It will therefore suffice to show that $g$ is a trivial fibration. Since $g$ is a categorical fibration, it suffices to show that it is a categorical equivalence. We may identify $g$ with the map between the pullbacks of the rows of the following commutative diagram:
$\operatorname{Fun}(\Delta^{m},\mathcal{D})\times_{\operatorname{Fun}(\mathcal{C}_{0}\cap\Delta^{m},\mathcal{D})}\operatorname{Fun}(\mathcal{C}_{0}\cap\Delta^{m},\mathcal{Z})$$\operatorname{Fun}(\{m\},\mathcal{Z})$$\operatorname{Fun}(\Delta^{\{m,\dots,d\}},\mathcal{D})\times_{\operatorname{Fun}(\{m\},\mathcal{D})}\operatorname{Fun}(\{m\},\mathcal{Z})$$\operatorname{Fun}(\Delta^{m}\cap\mathcal{C}_{1}^{\mathrm{isltd}},\mathcal{D})$$\Delta^{0}$$\operatorname{Fun}(\Delta^{\{m,\dots,d\}}\cap\mathcal{C}_{1}^{\mathrm{isltd}},\mathcal{D}).$}}
By the induction hypothesis, the vertical arrows of the above diagrams are trivial fibrations. Moreover, the horizontal arrows are categorical fibrations of $\infty$-categories. Hence $g$ is also a categorical equivalence, and the proof is complete. ∎
Appendix C Miscellany
In this section, we briefly recall some key definitions and results that did not quite fit into the main body of the paper, but are nonetheless used in it.
C.1. Germs
[…]
Definition C.1.
[Lur17, Definition 5.4.1.6] Let $n\geq 0$. For each positive number $r>0$, let $B^{n}\left(r\right)\subset\mathbb{R}^{n}$ denote the open ball of radius $r$ centered at the origin. Given a finite set $S$ and an $n$-manifold $M$, we define the simplicial set $\operatorname{Germ}\left(S,M\right)$ of $S$-germs of $M$ as the colimit
$\operatorname{Germ}\left(S,M\right)=\operatorname{colim}_{r\in\left(\mathbb{R}_{>0}\right)^{\mathrm{op}}}\operatorname{Sing}\operatorname{Emb}\left(S\times B^{n}\left(r\right),M\right).$
If $S$ is a singleton, we simply write $\operatorname{Germ}\left(S,M\right)=\operatorname{Germ}\left(M\right)$. The evaluation at the origin determines a map $\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$. For each $p\in M$, we set $\operatorname{Germ}_{p}\left(M\right)=\operatorname{Germ}\left(M\right)\times_{\operatorname{Sing}M}\{p\}$.
Remark C.2.
Let $n\geq 0$. The simplicial set $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$ has the structure of a simplicial monoid. To see this, note that its $k$-simplex is represented by a map $\sigma:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to\mathbb{R}^{n}$, where $r>0$ is a positive number and for each point $x\in\left|\Delta^{k}\right|$, the map $\sigma\left(x,-\right):B^{n}\left(r\right)\to\mathbb{R}^{n}$ is an embedding which fixes the origin. Given another such map $\tau:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to\mathbb{R}^{n}$, the product $\tau]\circ[\sigma]$ is represented by the map
$\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime\prime}\right)\xrightarrow{\left(\operatorname{id},\sigma\right)}\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\xrightarrow{\tau}\mathbb{R}^{n},$
where $0<r^{\prime\prime}<r$ is chosen so that $\sigma|\Delta^{k}\times B^{n}\left(r^{\prime\prime}\right)$ takes values in $B^{n}\left(r^{\prime}\right)$. (Such a number $r^{\prime\prime}$ exists because $\left|\Delta^{k}\right|$ is compact.)
More generally, if $M$ is an $n$-manifold and $S$ is a finite set, then $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$ acts on $\operatorname{Germ}\left(S,M\right)$ (from the right) as follows: A $k$-simplex of $\operatorname{Germ}\left(S,M\right)\times\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$ is represented by maps $\tau:\left|\Delta^{k}\right|\times\coprod_{s\in S}B^{n}\left(r\right)\to M$ and $\left(\sigma_{s}:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to\mathbb{R}^{n}\right)_{s\in S}$ where $r,r^{\prime}>0$. Its image in $\operatorname{Germ}\left(S,M\right)$ is represented by the composite
$\left|\Delta^{k}\right|\times\coprod_{s\in S}B^{n}\left(r^{\prime\prime}\right)\xrightarrow{\left(\operatorname{id},\coprod_{s}\sigma_{s}\right)}\left|\Delta^{k}\right|\times\coprod_{s\in S}B^{n}\left(r^{\prime}\right)\xrightarrow{\tau}M,$
where $r^{\prime\prime}>0$ is chosen so that the above composite is defined.
Remark C.3.
Let $n\geq 0$, let $M$ be an $n$-manifold, and let $S$ be a finite set. As a filtered colimit of Kan complexes, the simplicial set $\operatorname{Germ}\left(S,M\right)$ is a Kan complex. Moreover, the map
$\operatorname{Sing}\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\to\operatorname{Germ}\left(S,M\right)$
is a homotopy equivalence of Kan complexes [Lur17, Proposition 5.4.1.8].
The importance of the germ construction lies in the following propositions:
Proposition C.4.
Let $n\geq 0$. The simplicial monoid $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$ is a simplicial group.
Proposition C.5.
Let $n\geq 0$, let $M$ be an $n$-manifold, and let $S$ be a finite set. The map
$\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$
is a principal $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)^{S}$-fibration.
The proofs of Propositions C.4 and C.5 require a few preliminaries.
Lemma C.6.
Let $n\geq 1$, let $X$ be a locally path-connected topological space, let $M$ be an $n$-manifold, and let
$f,g:X\times D^{n}\to M$
be continuous maps. Suppose that, for each $x\in X$, the map $f_{x}=f\left(x,-\right):D^{n}\to M$ is injective. If there is a point $x_{0}\in X$ such that $g_{x_{0}}\left(D^{n}\right)\subset f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$, then there is a neighborhood $U$ of $x_{0}$ such that $g_{x}\left(D^{n}\right)\subset f_{x}\left(\operatorname{Int}D^{n}\right)$ for all $x\in U$.
Proof.
Since $f_{x_{0}}|\operatorname{Int}D^{n}$ is an injective continuous map between $n$-manifolds, it is open. In particular, the set $V=f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$ is open. Thus, replacing $X$ by a neighborhood of $x_{0}$ if necessary, we may assume that $g$ takes values in $f_{x_{0}}\left(\operatorname{Int}D^{n}\right)$. The subset $f_{x_{0}}^{-1}\left(g_{x_{0}}\left(D^{n}\right)\right)\subset D^{n}$ is compact and lies in the interior of $D^{n}$, so we can find some number $0<r<1$ such that $f_{x_{0}}^{-1}\left(g_{x_{0}}\left(D^{n}\right)\right)\subset B^{n}\left(r\right)$. Replacing $X$ by a neighborhood of $x_{0}$ if necessary, we may assume that $f\left(X\times\overline{B^{n}\left(r\right)}\right)\subset V$.
[…]
$\sup_{x\in U}\operatorname{dist}\left(f_{x}\left(\partial\overline{B^{n}\left(r\right)}\right),f_{x_{0}}\left(\partial\overline{B^{n}\left(r\right)}\right)\right)$ $<\varepsilon/2,$
$\sup_{x\in U}\operatorname{dist}\left(g_{x}\left(D^{n}\right),g_{x_{0}}\left(D^{n}\right)\right)$ $<\varepsilon/2.$
We claim that $U$ has the desired properties. In fact, we will show that $g_{x}\left(D^{n}\right)\subset f_{x}\left(B^{n}\left(r\right)\right)$ for all $x\in U$.
Let $x_{1}\in U$. We will derive a contradiction by assuming that there is a point $p\in D^{n}$ such that $q=g_{x_{1}}\left(p\right)\not\in f_{x_{1}}\left(B^{n}\left(r\right)\right)$. By construction, for each $x\in U$, the point $q$ does not belong to $f_{x}\left(\partial\overline{B^{n}\left(r\right)}\right)$. Since $U$ is path-connected, the maps $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ and $f_{x_{1}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ are homotopic. The latter map is null homotopic because it extends to all of $\overline{B^{n}\left(r\right)}$. Hence the map $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ is also nullhomotopic. On the other hand, the distance from $q$ to $g_{x_{0}}\left(D^{n}\right)$ is less than $\varepsilon/2$, so $q$ belongs to $f_{x_{0}}\left(B_{n}\left(r\right)\right)$. It follows that the map $f_{x_{0}}:\partial\overline{B^{n}\left(r\right)}\to V\setminus\{q\}$ is a homotopy equivalence. Thus $\partial\overline{B^{n}\left(r\right)}$ is contractible, which is a contradiction. ∎
Proposition C.7.
Let $n\geq 0$. For any $n$-manifold $M$ and any finite set $S$, the evaluation at the origin determines a Serre fibration
$\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\to\operatorname{Conf}\left(S,M\right).$
The following proof of Proposition C.7 is due to Alexander Kupers [sku].
Proof.
Since Serre fibrations can be recognized locally [tD08, 6.3.3], it suffices to prove the following:
  • (A)
    For any collection $\{U_{s}\}_{s\in S}$ of pairwise disjoint open sets of $M$ such that each $U_{s}$ is homeomorphic to $\mathbb{R}^{n}$, the map
    $\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\times_{\operatorname{Conf}\left(S,M\right)}\prod_{s\in S}U_{s}\to\prod_{s\in S}U_{s}$
    is a Serre fibration.
We begin with a preliminary assertion. Let $\operatorname{Homeo}\left(M\right)\subset\operatorname{Emb}\left(M,M\right)$ denote the subspace consisting of the self-homeomorphisms of $M$. We prove the following:
  • (B)
    Let $U\subset M$ be an open set homeomorphic to $\mathbb{R}^{n}$, and let $K\subset U$ be a compact subset. There is a continuous map
    $\chi:K\times K\to\operatorname{Homeo}\left(M\right)$
    with the following properties:
    1. $\bullet$
      For each $\left(x,y\right)\in K\times K$, the support of $\chi\left(x,y\right)$ is compact and lies in $U$ and, and moreover $\chi\left(x,y\right)\left(x\right)=y$.
    2. $\bullet$
      For each $x\in K$, we have $\chi\left(x,x\right)=\operatorname{id}_{\mathbb{R}^{n}}$.
To prove (B), it suffices to consider the case where $U=M=\mathbb{R}^{n}$ and $K=D^{n}$. Choose a compactly supported smooth function $\phi:\mathbb{R}\to\mathbb{R}$ such that $\phi\left(t\right)=1$ if $\left|t\right|\leq 1$ and $\sup_{t\in\mathbb{R}}\left|\phi^{\prime}\left(t\right)\right|<1/2$. We define $\chi:D^{n}\times D^{n}\to\operatorname{Homeo}\left(\mathbb{R}^{n}\right)$ by
$\chi\left(x,y\right)\left(z\right)=\left(\left(1-\phi\left(z_{i}\right)\right)z_{i}+\phi\left(z_{i}\right)\left(z_{i}+y_{i}-x_{i}\right)\right)_{i=1}^{n}.$
Note that this is well-defined. To see this, it suffices to show that, for any real number $r\in[-2,2]$, the function
$z\mapsto\left(1-\phi\left(z\right)\right)z+\phi\left(z\right)\left(z+r\right)=z+\phi\left(z\right)r$
is a self-homeomorphism of $\mathbb{R}$. This is clear, as the right-hand side is an increasing function in $z$ (because $\sup_{t\in\mathbb{R}}\left|\phi^{\prime}\left(t\right)\right|<1/2$). The function $\chi$ has the desired properties, so we have proved (B).
Now we prove assertion (A). Let $k\geq 0$ and consider a solid commutative diagram
$I^{k}\times\{0\}$$\operatorname{Emb}(\mathbb{R}^{n}\times S,M)\times_{\operatorname{Conf}(S,M)}\prod_{s\in S}U_{s}$$I^{k}\times I$$\prod_{s\in S}U_{s}$$\scriptstyle{F}$$\scriptstyle{G}$
We wish to find a map $I^{k}\times I\to\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right)\times_{\operatorname{Conf}\left(S,M\right)}\prod_{s\in S}U_{s}$ rendering the diagram commutative. Using assertion (B), we can construct a map
$\Phi:I^{k}\times I\to\operatorname{Homeo}\left(M\right)$
such that $\Phi\left(x,0\right)=\operatorname{id}_{M}$ and $\Phi\left(x,t\right)\left(G_{s}\left(x,0\right)\right)=G_{s}\left(x,t\right)$ for all $s\in S$ and $\left(x,t\right)\in I^{k}\times I$; here $G_{s}$ denotes the $s$th component of $G$. The map
$I^{k}\times I\to\operatorname{Emb}\left(\mathbb{R}^{n}\times S,M\right),\,\left(x,t\right)\mapsto\Phi\left(x,t\right)\circ F\left(x,0\right)$
gives the desired filler. ∎
Corollary C.8.
Let $n\geq 0$. For any $n$-manifold $M$ and any finite set $S$, the map
$\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\operatorname{Conf}\left(S,M\right)$
is a Kan fibration.
Proof.
This follows from Proposition C.7, for Kan fibrations are stable under filtered colimits. ∎
We can now prove Propositions C.4 and C.5.
Proof of Proposition C.4.
The claim is trivial if $n=0$, so we will assume that $n\geq 1$. Let $\sigma:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to\mathbb{R}^{n}$ be a representative of a $k$-simplex $\sigma]\in\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)$. We wish to construct an inverse of $\sigma]$.
For each $x\in\left|\Delta^{k}\right|$, let $\sigma_{x}=\sigma\left(x,-\right)$ denote the restriction of $\sigma$. We will first show that there is some $s>0$ which satisfies $\overline{B^{n}\left(s\right)}\subset\sigma_{x}\left(\overline{B^{n}\left(r/2\right)}\right)$ for every $x\in\left|\Delta^{k}\right|$. Since $\left|\Delta^{k}\right|$ is compact, it will suffice to prove the following:
  • ($\ast$)
    For each $x_{0}\in\left|\Delta^{k}\right|$, there are neighborhoods $U_{x_{0}}$ of $x_{0}$ and some number $s\left(x_{0}\right)>0$ such that $\overline{B^{n}\left(s\left(x_{0}\right)\right)}\subset\sigma_{x}\left(\overline{B^{n}\left(r/2\right)}\right)$ for every $x\in U_{x_{0}}$.
Assertion ($\ast$) follows from Lemma C.6.
Now we construct the inverse of $\sigma]$. Choose $s>0$ as in the previous paragraph. The map
$F:\sigma^{-1}\left(\overline{B^{n}\left(s\right)}\right)\to\left|\Delta^{k}\right|\times\overline{B^{n}\left(s\right)},\,\left(x,p\right)\mapsto\left(x,\sigma_{x}\left(p\right)\right)$
is a continuous bijection of compact Hausdorff spaces, so it is a homeomorphism. Consider the composite
$\tau:\left|\Delta^{k}\right|\times B^{n}\left(s\right)\xrightarrow{F^{-1}}\sigma^{-1}\left(\overline{B^{n}\left(s\right)}\right)\xrightarrow{\operatorname{pr}}\mathbb{R}^{n}.$
We have $\tau][\sigma]=1=[\sigma][\tau]$ in $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}$, so $\tau]$ is the desired inverse of $\sigma]$. ∎
Proof of Proposition C.5.
Since the map $\operatorname{Germ}\left(S,M\right)\to\operatorname{Sing}\left(\operatorname{Conf}\left(S,M\right)\right)$ is a pullback of the map $\operatorname{Germ}\left(M\right)^{S}\to\operatorname{Sing}M^{S}$, it suffices to consider the case where $S$ is a singleton. Let $k\geq 0$. The action of $\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}$ on $\operatorname{Germ}\left(M\right)_{k}$ is clearly free. It will therefore suffice to show that the induced map
$\theta:\operatorname{Germ}\left(M\right)_{k}/\operatorname{Germ}_{0}\left(\mathbb{R}^{n}\right)_{k}\to\operatorname{Sing}\left(M\right)_{k}$
is bijective. Surjectivity of $\theta$ is immediate from Corollary C.8. To prove that $\theta$ is injective, we must prove the following:
  • ($\ast$)
    Let $r>0$ and let $\sigma,\tau:\left|\Delta^{k}\right|\times B^{n}\left(r\right)\to M$ be continuous maps such that, for each $x\in\left|\Delta^{k}\right|$, the maps $\sigma_{x}=\sigma\left(x,0\right):B^{n}\left(r\right)\to M$ and $\tau_{x}=\tau\left(x,0\right):B^{n}\left(r\right)\to M$ are embeddings satisfying $\sigma_{x}\left(0\right)=\tau_{x}\left(0\right)$. Then there are some number $r^{\prime}>0$ and a map $g:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to B^{n}\left(r\right)$ such that
    $\sigma\left(x,g\left(x,p\right)\right)=\tau\left(x,p\right)$
    for every $\left(x,p\right)\in\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)$.
Using Lemma C.6 and the compactness of $\left|\Delta^{k}\right|$, we can find some $0<r^{\prime}<r$ which satisfies $\tau_{x}\left(\overline{B^{n}\left(r^{\prime}\right)}\right)\subset\sigma_{x}\left(B^{n}\left(r/2\right)\right)$ for every $x\in\left|\Delta^{k}\right|$. The map
$F:\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}$ $\to\left|\Delta^{k}\right|\times M$
$\left(x,p\right)$ $\mapsto\left(x,\sigma_{x}\left(p\right)\right)$
is a continuous injection from a compact space to a Hausdorff space, so it is a homeomorphism onto its image. We define $g:\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\to B^{n}\left(r\right)$ as the composite
$\left|\Delta^{k}\right|\times B^{n}\left(r^{\prime}\right)\xrightarrow{\tau^{\prime}}F\left(\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}\right)\xrightarrow{F^{-1}}\left|\Delta^{k}\right|\times\overline{B^{n}\left(r/2\right)}\xrightarrow{\operatorname{pr}}B^{n}\left(r\right),$
where $\tau^{\prime}\left(x,p\right)=\left(x,\tau_{x}\left(p\right)\right)$. The map $g$ has the desired properties. ∎
We conclude this subsection with another consequence of Corollary C.8.
Proposition C.9.
Let $n\geq 0$, let $S$ be a finite set, let $M$ and $N$ be $n$-manifolds, and let $\phi:M\to N$ be an embedding. The square
$\operatorname{Emb}(\mathbb{R}^{n}\times S,M)$$\operatorname{Emb}(\mathbb{R}^{n}\times S,N)$$\operatorname{Conf}(S,M)$$\operatorname{Conf}(S,N),$}}
determined by the evaluation at the origin, is homotopy cartesian (with respect to the Quillen model structure for topological spaces).
Proof.
It will suffice to show that the square becomes homotopy cartesian after applying the singular complex functor. Using Remark C.3, we are reduced to showing that the square
$\operatorname{Germ}(S,M)$$\operatorname{Germ}(S,N)$$\operatorname{Sing}\operatorname{Conf}(S,M)$$\operatorname{Sing}\operatorname{Conf}(S,N)$}}
is homotopy cartesian. By Corollary C.8, the vertical arrows of this diagram are Kan fibrations. It will therefore suffice to show that the square is strictly cartesian, which is clear. ∎
Remark C.10.
Proposition C.9 remains valid if $\phi$ is a smooth embedding of smooth manifolds of the same dimension and $\operatorname{Emb}$ is replaced by $\operatorname{Emb}_{\mathrm{sm}}$. This is because the smooth version of Proposition C.7 is also true, with the same proof.
C.2. Simplicial Complexes
The term simplicial complex can refer to several closely related notions depending on context. Since definitions are not entirely standardized, we briefly recall the conventions and basic properties relevant to this paper.
Definition C.11.
A simplicial complex $K$ is a locally finite collection of simplices in a Euclidean space, which is closed under finite intersection and the operation of taking faces. A subcomplex of $K$ is a subset $L$ of $K$ which is itself a simplicial complex.
The union of the simplices in $K$ is called its underlying polyhedron and is denoted by $\left|K\right|$. (In the main body of the paper, we will often blur the distinction between simplicial complexes and their underlying polyhedra.) A subdivision of $K$ is a simplicial complex $K^{\prime}$ with the same underlying polyhedron as $K^{\prime}$, such that every simplex of $K^{\prime}$ lies in $K$. In this situation, we write $K^{\prime}\triangleleft K$.
A triangulation of a topological space $X$ is a homeomorphism $t:\left|K\right|\to X$ with $K$ a simplicial complex.
Example C.12.
Let $K$ be a simplicial complex, and let $k\geq 0$. There is a triangulation of $\operatorname{SP}^{k}\left(K\right)=K^{k}/\Sigma_{k}$ containing $\blacktriangle_{k}\left(K\right)$ as its subcomplex. To see this, choose a well-ordering of the vertices of $K$. The products $K^{k}$ is a union of cells of the form $\sigma_{1}\times\cdots\times\sigma_{k}$, where each $\sigma_{i}$ is a simplex of $M$. We subdivide each such cell by the simplices spanned by the vertices of form $\{\left(a_{1}^{\left(i\right)},\dots,a_{k}^{\left(i\right)}\right)\}_{0\leq i\leq d}$, where $\left(a_{1}^{\left(0\right)},\dots,a_{k}^{\left(0\right)}\right)<\cdots<\left(a_{1}^{\left(d\right)},\dots,a_{k}^{\left(d\right)}\right)$ are vertices of $\sigma_{1}\times\cdots\times\sigma_{k}$. (The ordering on the product is given by the categorical product of posets.) This gives rise to a triangulation on $\operatorname{SP}^{k}\left(K\right)$ with the desired properties.
Definition C.13.
Let $K$ be a simplicial complex, and let $L\subset K$ be a subcomplex. The simplicial neighborhood of $L$ in $K$, denoted by $N\left(L,K\right)$, is the subcomplex of $K$ generated by the simplices intersecting $\left|L\right|$ nontrivially. The simplicial complement of $L$ in $K$, denoted by $C\left(L,K\right)$, is the subcomplex of $K$ consisting of the simplices not intersecting $\left|L\right|$.
A subdivision $K^{\prime}\triangleleft K$ is said to be obtained from $K$ by deriving it near $L$ if it is obtained by replacing the simplices of $K\setminus\left(L\cup C\left(L,K\right)\right)$ by the following procedure: For each $\sigma\in K\setminus\left(L\cup C\left(L,K\right)\right)$, we choose a point $a_{\sigma}$ in the interior of $\sigma$. We then adjoin $K\setminus\left(L\cup C\left(L,K\right)\right)$ to the faces simplices of the form $\sigma\star a_{\tau_{1}}\star\cdots\star a_{\tau_{k}}$, where $k\geq 1$, $\sigma\in L\cup C\left(L,K\right)$, and $\tau_{1}\subset\cdots\subset\tau_{k}$ are simplices in $K\setminus\left(L\cup C\left(L,K\right)\right)$ with $\sigma\subset\tau_{1}$. (Here $\star$ denotes the join operation.) We call $N\left(L,K^{\prime}\right)$ a derived neighborhood of $L$ in $K$.
Remark C.14.
Let $K$ be a simplicial complex, and let $L\subset K$ be a subcomplex. Then $\left|N\left(L,K\right)\right|$ is a neighborhood of $\left|L\right|$ in $\left|K\right|$. Moreover, $N\left(L,K\right)$ is the smallest subcomplex of $K$ having this property.
Remark C.15.
Let $K$ be a simplicial complex, and let $L\subset K$ be a subcomplex. Then $\left|C\left(L,K\right)\right|$ is a deformation retract of $\left|K\right|\setminus\left|L\right|$. Indeed, every point of $\left|K\right|\setminus\left(\left|L\right|\cup\left|C\left(L,K\right)\right|\right)$ can be written uniquely as $tx+\left(1-t\right)y$, where $x\in\left|L\right|$, $y\in\left|C\left(L,K\right)\right|$, and $t\in\left(0,1\right)$, and an inverse homotopy equivalence is given by mapping such a point to $y$.
Remark C.16.
Simplicial neighborhoods are compatible with the operation of taking intersections with subcomplexes, in the following sense: Let $K$ be a simplicial complex, and let $K_{0},L\subset K$ be subcomplexes. Set $L_{0}=K_{0}\cap L$. We have
$N\left(K,L\right)\cap K_{0}=N\left(K_{0},L_{0}\right).$
Remark C.17.
Derived neighborhoods are compatible with the operation of taking intersections with subcomplexes, in the following sense: Let $K$ be a simplicial complex, and let $K_{0},L\subset K$ be subcomplexes. Suppose we are given a subdivision $K^{\prime}\triangleleft K$ is obtained by deriving it near $L$, and let $K^{\prime}_{0}\triangleleft K_{0}$ be the subdivision consisting of the simplices of $K^{\prime}$ lying in $\left|K_{0}\right|$. Then $K^{\prime}_{0}$ is obtained by deriving $K_{0}$ near $L_{0}$, and $L_{0}=K^{\prime}_{0}\cap L$. So by Remark C.16, we have
$N\left(K^{\prime},L\right)\cap K^{\prime}_{0}=N\left(K^{\prime}_{0},L_{0}\right).$
In particular, we have $\left|N\left(K^{\prime},L\right)\right|\cap\left|K_{0}\right|=\left|N\left(K^{\prime}_{0},L_{0}\right)\right|$.
Lemma C.18.
Let $K$ be a simplicial complex, let $L\subset K$ be a subcomplex, and let $U\subset\left|K\right|$ be a neighborhood of $\left|L\right|$. There is a derived neighborhood of $L$ in $K$ whose underlying polyhedron lies in $U$.
Proof.
Define $d_{L}:\left|K\right|\to[0,1]$ by mapping each vertex in $L$ to 0 and the remaining vertices to $1$, and then extending linearly on each simplex. Since simplices are compact, for each simplex $\sigma\in K\setminus\left(L\cup C\left(L,K\right)\right)$, there is some $\varepsilon_{\sigma}\in\left(0,1\right)$ such that $d_{L}^{-1}\left([0,\varepsilon_{\sigma}]\right)\cap\sigma\subset U$. By choosing $\varepsilon_{\sigma}$ in the order of decreasing dimension, we may assume that if $\sigma\subset\tau$, then $\varepsilon_{\sigma}\leq\varepsilon_{\tau}$. We then choose a point $a_{\sigma}\in\sigma^{\circ}\cap d_{L}^{-1}\left(\varepsilon_{\sigma}\right)$, and form a derived subdivision $K^{\prime}$ of $K$ by using the points $a_{\sigma}$. We claim that $\left|N\left(L,K^{\prime}\right)\right|\subset U$.
The simplicial complex $N\left(L,K^{\prime}\right)$ is generated by the simplices of the form $\sigma\star a_{\sigma_{1}\star\tau_{1}}\star\cdots\star a_{\sigma_{k}\star\tau_{k}}$, where $\sigma\subset\sigma_{1}\subset\cdots\subset\sigma_{k}$ are simplices of $L$ and $\tau_{1}\subset\cdots\subset\tau_{k}$ are simplices of $C\left(K,L\right)$. Such a cell is contained in $d_{L}^{-1}\left([0,\varepsilon_{\sigma_{k}\star\tau_{k}}]\right)\cap\left(\sigma_{k}\star\tau_{k}\right)$ and hence in $U$. Hence $\left|N\left(L,K^{\prime}\right)\right|\subset U$, as required. ∎